### v1

# FedFDP: Federated Learning with Fairness and Differential Privacy

Abstract
Federated learning (FL) is a new machine learning paradigm to overcome the challenge of data silos and has garnered significant attention. However, through our observations, a globally effective trained model may performance disparities in different clients. This implies that the jointly trained models by clients may lead to unfair outcomes. On the other hand, relevant studies indicate that the transmission of gradients or models in federated learning can also give rise to privacy leakage issues, such as membership inference attacks.
To address the first issue mentioned above, we propose a federated algorithm with fairness, termed FedFair. Building upon FedFair, we introduce privacy protection to form the FedFDP algorithm to address the second issue mentioned above. In FedFDP, we devise a fairness-aware clipping strategy to achieve differential privacy while adjusting fairness. Additionally, for the extra uploaded loss values, we present an adaptive clipping approach to maximize utility. Furthermore, we theoretically prove that our algorithm converges and ensures differential privacy. Lastly, Extensive experimental results demonstrate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. The code is accessible at https://anonymous.4open.science/r/FedFDP-E754.
1 Introduction
Federated Learning (FL) McMahan et al. (2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without revealing their individual datasets. After training their local models, each client transmits only the model parameters, not the raw training data, to a central server. The server aggregates these parameters to update the global model, which is then sent back to clients for further training. Because its ability to overcome data silos by jointly improving model performance without the need to upload local data, FL has gained significant attention in recent year Zhu et al. (2021)Lee et al. (2022). In particular, many studies have made great breakthroughs in model performance Wang et al. (2020); Li et al. (2020a, 2021b) and communication costs Wang et al. (2019); Ling et al. (2023) in FL.
However, recent research indicates that trained parameters can leak data privacy, such as the recovery of original data Zhu et al. (2019) or the membership inference attack Song et al. (2017). Moreover, Melis et al. Melis et al. (2022) demonstrated that clients’ training data can be leaked through shared models. In order to enhance data privacy in the FL, numerous scholars have begun integrating techniques such as differential privacy (DP) Dwork et al. (2014), homomorphic encryption Paillier (1999), and secure multi-party computation Yao (1982) into the federated learning framework. Among these, DP has emerged as the most widely acknowledged approach for privacy protection in federated learning, owing to its minimal computational overhead and rigorous mathematical framework. Many works Abadi et al. (2016); Carlini et al. (2019); Feldman (2020) have demonstrated that by adding proper DP noise in the training phase, the resulted machine learning models can prevent unintentional leakage of private training data, such as membership inference attacks.
Figure 1: The test accuracy of 50 clients utilizing a globally trained model in FL.
Another easily overlooked critical issue in federated learning is fairness. The globally trained model collaboratively by clients may exhibit significant performance disparities across datasets held by different clients. As shown in Fig.1, where the red dots indicate the test accuracy of the global model across various clients after the completion of federated learning using the FedAvg McMahan et al. (2017) algorithm. And we can see the globally trained model applied to individual clients, the performance significant performance disparities across differential clients. More and more scholars have begun to notice this unfair phenomenon and initiated research on it  Wang et al. (2020); Li et al. (2020b).
While numerous articles separately discuss fairness and differential privacy protection in federated learning, there is a lack of research integrating both aspects into federated learning. Balancing model performance, fairness, and differential privacy protection in federated learning presents a challenging problem. This paper aims to investigate fairness and differential privacy protection mechanisms in federated learning. By analyzing the relationship between them, we propose a federated learning framework that ensures both fairness and differential privacy protection, effectively guaranteeing model performance.
Our main contributions are listed below:
  1. 1.
    Firstly, we propose a fairness-aware federated algorithm, denoted as algorithm FedFair by designing a new local loss function that is constructed to simultaneously optimize the two objectives of fairness and model performance (Section 4). FedFair algorithm achieves the excellent fairness, as shown by the green dots in Fig.1.
  2. 2.
    Building upon FedFair, we propose algorithm FedFDP to achieve differential privacy protection (Section 5) . Particularly, we device a fairness-aware gradient clipping strategy to ensures both differential privacy and allows for the adjustment of fairness in federated learning (Section 5.1). Additionally, we propose an adaptive clipping approach method for the additional loss values uploaded by each client to achieve optimal utility (Section 5.2).
  3. 3.
    Furthermore, we demonstrate there exist an optimal fairness parameter $\lambda$ value to balance the model performance and fairness through the convergence analysis (Section 6).
  4. 4.
    And then, we perform privacy loss analysis of our algorithm FedFDP by the notion of Rényi differential privacy (RDP) to prove that it satisfies the differential privacy (Section 7).
  5. 5.
    Lastly, through a comprehensive empirical evaluation on three public datasets, we confirm that FedFair and FedFDP outperforms the state-of-the-art solutions in terms of model performance and fairness (Section 8).
2 Related Work
2.1 Fairness in Federated Learning
In the field of machine learning, a multitude of fairness concepts have been advanced (Mehrabi et al. Mehrabi et al. (2021), Yuxin et al. Shi et al. (2021)), each highlighting specific facets and addressing the concerns of diverse stakeholder groups. For instance, Mohri et al. Mohri et al. (2019) advocate for Good-Intent Fairness, while Zhou et al. Zhou et al. (2021) embrace Selection Fairness. Cong and Lyu et al. Cong et al. (2020)Lyu et al. (2020) employ principles from game theory Rabin (1993) to steer the allocation of rewards.
Currently, the predominant fairness criterion in Federated Learning is “Performance Parity”, as proposed by Li et al. Li et al. (2020b). This criterion aims to ensure fairness by evaluating the consistency of performance across different client devices in federated learning. This definition is widely adopted. Zhao et al. Zhao and Joshi (2022) suggested replacing the loss amplification mechanism in Li et al. (2020b) with a straightforward weight reallocation mechanism. Li et al. Li et al. (2021b) focus on optimizing the fairness of local models in personalized federated learning.
2.2 Federated Learning with Differential Privacy
Federated learning with differential privacy has become a common privacy protection method in federated learning. Enayat et al. Ullah et al. (2023) dynamically adjust the compression rate to minimize communication rounds in federated learning by responding to changes in training loss. Zhou et al. Zhou et al. (2023) investigated the effects of local iterations when utilizing the Laplace mechanism for perturbation. Fu et al. Fu et al. (2022) introduced adaptive noise adding, customizing it for different clients and communication rounds based on observed fluctuations in loss values. Ling et al. Ling et al. (2023) determined the optimal number of local epochs through convergence analysis, striking a balance between communication resources and privacy budget. Raouf et al. Kerkouche et al. (2021) improved model accuracy by updating only a fixed subset of model weights while keeping others unchanged.
The aforementioned works each independently address either fairness or privacy protection in federated learning. To the best of our knowledge, we are the first to concurrently consider both fairness and privacy in the context of federated learning.
3 Preliminaries
3.1 Federated Learning
We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model. The objective of Federated Learning is to minimize $F(\mathbf{w})$, in other words, to seek:
$\mathbf{w}^{*}\triangleq\arg\min F(\mathbf{w}).$ (1)
The function $F(\cdot)$ is defined as follows:
$F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (2)
Here, $F_{i}(\cdot)$ represents the local loss function of client $i$. The weights are defined as $p_{i}=\frac{|D_{i}|}{|D|}$, $\sum_{i=1}^{N}p_{i}=1$, $D_{i}$ represents the dataset of client $i$, and $D$ represents the union of all $D_{i}$.
3.2 DPSGD
In gradient-based methods, DPSGDAbadi et al. (2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
  • Firstly, computing the gradient of loss function:
    $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}).$ (3)
  • Secondly, clipping the gradient of per-sample to get the sensitivity:
    $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right).$ (4)
  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients:
    $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$ (5)
The final step is to proceed with gradient descent as usual. Here, $C$ represents the clipping bound, and $\sigma$ is the noise multiplier. If each client in FL uses DPSGD as the gradient descent method to protect their private data, FL will transform into Differentially Private Federated Learning (DPFL).
3.3 Fairness in Federated Learning
Definition 1.
(Fairness of performance distributionLi et al. (2020b)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally say that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
In this study, we employ the loss of model denote the performance. Equation (6) represents the specific form of fairness in federated learning in this paper, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients. Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients.
$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (6)
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, fairness hyperparameter $\lambda$, batch sample ratio $q$, local dataset $D_{i}$.
Output: the final trained model $\mathbf{w}_{T}$
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
2SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
5                   $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
14       $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
15       for $e\in[1,\cdots,E]$ do
16             for batch $b\in\mathcal{B}_{i}$ do
17                   $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i},b)-F(\mathbf{w}_{t})$;
18                   $\eta_{i}=\eta\cdot\left(1+\lambda\cdot\Delta_{i}\right)$;
19                   $\mathbf{w}_{t}^{i}=\mathbf{w}_{t}^{i}-\eta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},b)$;
20
21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ and compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
22       return $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})$;
23
Algorithm 1 FedFair
4 Proposed Algorithm: FedFair
In the previous section, we identified the performance objective function as shown in Equation (2) and the fairness objective function as shown in Equation (6). Both of these objective functions are composed of individual client objectives. In this section, we will focus on the performance and fairness of each client and combine these two objectives to construct a new objective function, based on which we propose a fair federated learning algorithm.
We simultaneously optimize performance and fairness, and can get the optimization objective for client $i$ as:
$\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}_{t})\right)^{2},$ (7)
where $F(\mathbf{w}_{t})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t-1})$ [^1] The reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training.and $\lambda$ is a hyperparameter used to adjust the degree of fairness, $\lambda\geq 0$. When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness.
The gradient of Equation (7) as follows:
$\nabla H_{i}(\mathbf{w}_{t}^{i})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})$
$=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (8)
where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t})$.
The stochastic gradient descent for Equation (7) as show:
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})$
$=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (9)
where $\eta_{i}=\eta(1+\lambda\Delta_{i})$.
Therefore, our approach can be viewed as establishing an adaptive learning rate that allows the model update to achieve both performance and fairness objectives simultaneously. Based on these principles, we introduce the FedFair algorithm, as depicted in Algorithm 1.
  • Server (lines 1-1): The server executes $T$ rounds of communication, each round sending the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
  • Local (lines 1-1): For $E$ epochs and each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
Furthermore, as the iterations advance, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$. In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods Zinkevich (2003).
Figure 2: FedFDP framework
5 Proposed Algorithm: FedFDP
In this section, we propose the FedFDP by integrating differential privacy protection base on FedFair. The flow of FedFDP algorithm as shown in Fig. 2. The FedFDP algorithm adds the following two additional processes to FedFair. Firstly, in Section 5.1, we present a fair-clipping strategy that not only achieves sensitivity but also allows for adjustments to the level of fairness. Furthermore, in Section 5.2, we expand the use of differential privacy protection to the additional loss values that local clients need to upload, suggesting an adaptive clipping method to maximize utility. The overall FedFDP algorithm is described in Section 5.3.
5.1 Fair-clipping Strategy for Gradient
In the general DPSGD algorithm Fu et al. (2023); Abadi et al. (2016); Bu et al. (2022); Yang et al. (2022), per-sample clipping is an essential process for obtaining sensitivity. In order to be compatible with per-sample clipping, we expand our objective function of client $i$ in FedFair, Equation (7), into Equation (10):
$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})\right)^{2},$ (10)
where $\xi_{j}\in\mathcal{B}_{t}^{i}$, $j=\{1,2,\cdots,|\mathcal{B}_{i}|\}$, and $|\mathcal{B}_{i}|$ is the batch size of client $i$.
Next, we compute the gradient of Equation (10):
$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\lambda\cdot\Delta_{loss}^{j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$
$=\left(1+\lambda\cdot\Delta_{loss}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$ (11)
where $\Delta_{loss}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$.
To achieve differential privacy protection, we replace loss function gradient by Equation (5.1) based on Equations (3)-(5), and reorganize the gradient descent equation for client $i$:
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta^{j}_{loss}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta^{j}_{loss}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (12)
where:
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta^{j}_{loss},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$ (13)
Hence, our strategy can be regarded as an fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model. When $\lambda=0$, Equation (13) corresponds to the traditional DP clipping method, and FedFDP no longer ensures fairness. For large $C$, Equation (13) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness. Conversely, for small $C$, Equation (13) simplifies to the second term, resulting in the gradient being clipped to the bound $C$. Our conclusions can be verified from the experiments 5.
5.2 Adaptive Clipping Method for Loss
Since the calculation of $\Delta_{loss}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy. As shown in lines 2-2 of Algorithm 2, We need to clip $F_{i}(\mathbf{w}_{t})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $C_{l}^{i,t}\cdot\sigma_{l}$ to ensure differential privacy. However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t})$ is challenging. A large $C_{l}^{i,t}$ helps avoid loss of fidelity to the original values, but it also leads to large injected noise.
Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t})$. For the $i$-th client, the adaptive clipping method uses the $differentially$ $private$ $mean$ $of$ $the$ $previous$ $round$ as the clipping bound for the current round. The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
$C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$ (14)
where
$\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))),$ (15)
To ensure the achievement of differential privacy during the clipping process, we use the noise-added loss as the clipping bound $C_{l}^{i,t}$. Each round’s clipping bound adheres to the post-processing nature of differential privacy when $t=\{1,\dots,T\}$.
5.3 Overall Algorithm
We utilize the equation form presented in (12) to enhance Algorithm 2. Compared with the Algorithm 1, FedFDP solely necessitates the application of differential privacy to gradient and loss.
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, noise multiplier for training $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
Output: the final trained model $\mathbf{w}_{T}$
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0}),C_{l}^{0}=\text{Initial()}$();
2SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
5                   $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
14       Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
15       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
16             $\Delta_{loss}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
17             Compute $C^{i,j}_{t}$ by Equation (13);
18             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
19
20      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\mathcal{N}(0,\mathbf{I})\right)$;
21       Compute clipping bound of loss $C_{l}^{i,t}$ by Equation (14) and (15) ;
22       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
23             Compute $F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})$;
24             $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
25
26      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})\leftarrow\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\mathcal{N}(0,\mathbf{I})\right)$;
27       return $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
28
Algorithm 2 FedFDP
  • DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$. Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 2 and execute the gradient decent.
  • DP for local loss (lines 2-2): As described in 5.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
6 Utility Analysis
In this section, we examine the convergence of the FedFDP algorithm to identify an optimal $\lambda$ that minimizes the convergence bound.
6.1 Convergence Analysis
We analyze the convergence of Algorithm 2 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound. Prior to proving this, we need to establish several assumptions:
Assumption 1.
$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 2.
$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 3.
Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random. The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=1,\cdots,N$ , $t=1,\cdots,T-1$ and $j=1,\cdots,|\mathcal{B}_{i}|$
Theorem 1.
Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 2 as follows:
$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (16)
where:
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
The meaning of $\Gamma$ is consistent with Li et al. (2020c), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive (16). For the detailed proof, please refer to Appendix A.1.
6.2 Optimal $\lambda$
When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{loss}^{j}$. We assume $Q_{0}\leq\Delta_{loss}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in (16) as a function of $\lambda$,
$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{2Q_{0}\lambda+1},$ (17)
where $a_{1}=G^{2}Q_{1}^{3}$, $a_{2}=6G^{2}Q_{1}^{2}$, $a_{3}=9Q_{1}G^{2}+2L\Gamma Q_{1}+2Q_{0}\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{4}=4G^{2}+2L\Gamma+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$.
We sketched a rough graph of $P(\lambda)$, where $\lambda\in[0,+\infty)$. Upon observing Fig. 3, it becomes evident that when $a_{1}$,$a_{2}$,$a_{3}$,$a_{4}$ greater than 00, there exists an optimal value of $\lambda$ that minimizes the right-hand side of Equation (16), aligning with the experimental results presented in Fig. 4.
Figure 3: An example of $P(\lambda)$
7 Privacy Analysis
Because each client’s locally trained pair of data samples has the same sampling rate and noise multiplier in each round, each client has the same privacy loss in each round. As a result, we only assess each client’s privacy loss from the perspective of a single client. Following that, we examine the privacy loss of FedPDF for the $i$-th client as follows.
Theorem 2.
(Privacy loss of FedFDP). The privacy loss in $i$-th client of FedFDP satisfies:
[…]
where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 3 in Appendix A.2, and $R_{loss}^{i}(\alpha$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 4 in Appendix A.2.
The details of the proof please refer to Appendix A.2.
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 98.67$\pm$0.09 7.3e-3$\pm$5.4e-3 87.05$\pm$1.02 3.6e-1$\pm$2.2e-1 62.03$\pm$1.00 1.1e0$\pm$8.9e-1
SCAFFLOD 98.45$\pm$0.03 3.0e-2$\pm$2.6e-2 86.77$\pm$0.25 6.1e0$\pm$4.0e0 62.21$\pm$1.37 9.9e0$\pm$1.1e1
FedProx 98.70$\pm$0.13 7.5e-3$\pm$5.1e-3 87.30$\pm$0.51 3.5e-1$\pm$2.1e-1 62.13$\pm$0.20 1.2e0$\pm$3.1e-1
FedDyn 98.40$\pm$0.23 8.0e-3$\pm$2.1e-3 87.48$\pm$0.31 6.2e-1$\pm$3.5e-1 62.33$\pm$1.15 3.2e0$\pm$8.8e-1
q-FFL 98.72$\pm$0.03 4.7e-3$\pm$1.0e-3 87.35$\pm$0.13 5.1e-1$\pm$5.2e-2 63.33$\pm$0.21 9.4e-1$\pm$3.3e-2
FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
Table 1: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 93.40$\pm$0.47 1.1e11$\pm$0.5e10 84.15$\pm$1.97 3.0e8$\pm$1.4e8 52.61$\pm$0.17 6.1e9$\pm$5.2e9
SCAFFLOD 93.95$\pm$1.39 7.0e10$\pm$4.9e9 79.41$\pm$4.87 3.3e9$\pm$4.7e9 53.85$\pm$1.16 6.2e9$\pm$1.3e9
FedProx 90.50$\pm$3.95 5.3e10$\pm$7.3e9 83.85$\pm$0.98 7.9e8$\pm$4.5e8 51.34$\pm$0.91 4.4e9$\pm$2.1e9
FedDyn 91.42$\pm$2.25 5.6e10$\pm$9.8e9 84.04$\pm$1.21 8.8e8$\pm$3.2e8 52.14$\pm$1.39 5.5e9$\pm$1.2e9
ALI-DPFL 90.89$\pm$1.65 3.3e10$\pm$8.3e9 83.65$\pm$1.68 6.6e8$\pm$3.5e8 52.05$\pm$1.26 3.6e9$\pm$1.6e9
q-FFL 93.74$\pm$1.41 7.8e10$\pm$1.1e11 83.13$\pm$2.74 4.2e9$\pm$2.6e9 48.46$\pm$1.00 4.7e9$\pm$1.9e9
FedFDP 95.13$\pm$0.83 2.3e10$\pm$1.0e10 85.99$\pm$0.76 2.8e8$\pm$0.8e8 54.21$\pm$0.98 2.6e9$\pm$1.6e9
Table 2: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
8 Experiments
In this section, we conduct experiments to demonstrate the performance and fairness of FedFair and FedFDP using three real datasets. Additionally, we investigate the impact of various factors on FedFDP, such as fairness hyperparameter $\lambda$, original clipping bound $C$, heterogeneity and scalability. The impact of heterogeneity and scalability please refer to Appendix A.3).
8.1 Default Settings
We conducted experiments on three computer vision (CV) datasets: MNIST LeCun et al. (1998), FashionMNIST Xiao et al. (2017), and CIFAR10 Krizhevsky (2009). And we comparised our algorithms to six baselines, which include FedAvgMcMahan et al. (2017), SCAFFLODKarimireddy et al. (2020), FedProxLi et al. (2020a), FedDynAcar et al. (2021), ALI-DPFLLing et al. (2023) and q-FFLLi et al. (2020b). For the details of parameters settings and the descriptions of baselines please refer to Appendix A.3.1.
8.2 Results
Performance and Fairness. The experimental results of the FedFair algorithm and baseline methods are shown in Table 1. FedFair notably enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.5%. And Table 2 shows the experimental results of the FedFDP algorithm and baseline methods. Under the same privacy budget, FedFDP demonstrates improvements in model accuracy of 1.2%, 1.8%, and 0.4% compared to the best baseline on the three datasets, while also enhancing fairness by 30.3%, 6.7%, and 27.8% respectively compared to the best baseline.
Figure 4: Impact of different $\lambda$ in FedFDF algorithm
Impact of different $\lambda$. We investigated the correlation between $\lambda$ and fairness $\Psi$ in FedFDF algorithm. As depicted in Figure 4, our findings align with the theoretical analysis in Section 6.1: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase.
Figure 5: Impact of different $C$ in FedFDF algorithm
Impact of different $C$. In Section 5.1, we emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously. As illustrated in Figure 5, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values of [0.01, 0.1, 1, 5, 10]. Although $\Psi$ reaches its minimum when $C=1$, at this point, due to excessive noise addition, the model does not converge. As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$. When $C=0.1$, both accuracy and $\Psi$ achieve relatively good values.
9 Conclusion
In this paper, firstly, we propose a FedFair algorithm effectively addresses fairness concerns by optimizing a novel local loss function. Building upon FedFair, we propose the FedFDP, incorporates a fairness-aware gradient clipping strategy and an adaptive clipping method for additional loss values to achieve both fairness and differential privacy protection. And then, through convergence analysis, we determine the existence of an optimal fairness parameter, $\lambda$, striking a balance between model performance and fairness. We conduct a comprehensive privacy analysis of our approach using RDP and validate our algorithms through extensive experiments. The results indicate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. We believe our work contributes a valuable FL framework for addressing fairness and privacy challenges.
Appendix A Appendix
In the beginning, Table 3 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
Table 3: Summary of main notations
$F(\mathbf{w}$)
Global loss function
$F_{i}(\mathbf{w}$)
Local loss function for client $i$
$T$
Communication rounds
$t$
Communication round index
$D_{i}$
Dataset of client $i$
$D$
The union of $D_{i}$
$\mathbf{w}_{t}^{i}$
Local model parameter at node $i$ at iteration $t$
$\mathbf{w}^{*}$
Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
$\eta$
Gradient descent step size
$\lambda$
Fairness parameter
$N$
Total number of client
$q$
batch sample ratio
$|\mathcal{B}_{i}|$
Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
$d$
The dimension of model
$p_{i}$
Weight of client $i$, equals to $|D_{i}|/|D|$
$\xi_{j}$
Single data sample in $D_{i}$, indexed by $j$
$\sigma$
Noise multiplier for gradient
$\sigma_{l}$
Noise multiplier for loss
$C$
Original clipping bound for gradient
$C_{l}$
Original Clipping bound for loss
|||||$\cdot$||| | ⋅ | |
$L_{2}$-norm
A.1 Proof of Theorem 1
For the purpose of validation, we introduce an additional variable $\mathbf{v}_{t}^{i}$ to represent the immediate result of a single-step DPSGD update from $\mathbf{w}_{t}^{i}$. We interpret $\mathbf{w}_{t+1}^{i}$ as the parameter obtained after a single communication step. Consequently, the fair-clipping DPSGD in client $i$ at iteration $t$ transitions from Equation (12) to:
$\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (19)
where:
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta^{j}_{loss},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by Stich [2018]. Therefore,
$\mathbf{v}_{t+1}$ $=\mathbf{w}_{t}$
$\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (20)
A.1.1 Key Lemma
Lemma 1.
(Results of one iteration.)Assume Assumption 1-3 hold, we have:
$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq\left(1-\mu\eta C_{t}\right)\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
where:
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Let $\Delta_{t}\triangleq\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. It is evident that we always have $\mathbf{w}_{t+1}=\mathbf{v}_{t+1}$. According to Lemma 1, this implies:
$\Delta_{t+1}\leq\left(1-\mu\eta C_{t}\right)\Delta_{t}+\eta^{2}A$
We use mathematical induction to obtain $\Delta_{t}\leq\frac{v}{t}$ where $v=\max\{\frac{\beta^{2}A}{\mu\beta C_{t}-1},\Delta_{1}\}$, $\eta=\frac{\beta}{t}$ for some $\beta>\frac{1}{\mu}$.
STEP 1. When $t=1$, the equation $\Delta_{1}\leq v$ holds obviously.
STEP 2. We assume $\Delta_{t}\leq\frac{v}{t}$ holds.
STEP 3.
$\Delta_{t+1}$ $\leq\left(1-\mu\frac{\beta}{t}C_{t}\right)\frac{v}{t}+\frac{\beta^{2}A}{t^{2}}$
$=\frac{t-1}{t^{2}}v+\left(\frac{\beta^{2}A}{t^{2}}-\frac{\mu\beta C_{t}-1}{t^{2}}v\right)$
$\leq\frac{t-1}{t^{2}}v$
$\leq\frac{v}{t+1}$
Therefore, $\Delta_{t+1}\leq\frac{v}{t+1}$ holds, completing the proof by mathematical induction. Hence, $\Delta_{t}\leq\frac{v}{t}$ holds.
Then by the $L$-smoothness of $F(\cdot)$, let $\beta=\frac{2}{\mu}$, we get
$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2}\Delta_{t}\leq\frac{L}{2t}v\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\Delta_{1}\right)$
A.1.2 Proof of Lemma 1
By the Equation (A.1), we get
$\left\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\right\|^{2}$
$*\displaystyle=\|\mathbf{w}_{t}-\mathbf{w}^{*}$
$-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\|^{2}$
$=\left\|\mathbf{w}_{t}-\mathbf{w}^{*}\right\|^{2}$
$\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\mathcal{N}(0,\mathbf{I})]}_{\mathcal{A}_{1}}\rangle$
$+\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}}_{\mathcal{A}_{2}}$
Firstly, we process $\mathcal{A}_{2}$:
$\mathcal{A}_{2}\leq\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2}}_{\mathcal{B}_{1}}$
$\quad+\underbrace{\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\left\langle\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right\rangle}_{\mathcal{B}_{0}}$
$\quad+\underbrace{\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\cdot\mathcal{N}(0,\mathbf{I})\|^{2}}_{\mathcal{B}_{2}}$
Since $\mathbb{E}\mathcal{B}_{0}=0$, we focus on $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$:
$\mathbb{E}\mathcal{B}_{2}\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
where $\frac{1}{\hat{B}^{2}}\triangleq\max_{i}\frac{1}{|\mathcal{B}_{i}|}$.
By the convexity of $\|\cdot\|^{2}$,
$\mathcal{B}_{1}\leq\eta^{2}\sum_{i=1}^{N}p_{i}\left\|\frac{1}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2},$
taking exception and according to assumption 3:
$\mathbb{E}\mathcal{B}_{1}\leq\eta^{2}\mathbb{E}\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\leq\eta^{2}C_{t}^{2}G^{2},$
where $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i}$, $C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Now, we obtain the bound for the expectation of $\mathcal{A}_{2}$:
$\mathbb{E}\mathcal{A}_{2}$ $\leq\eta^{2}\left(\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}+G^{2}C_{t}^{2}\right)$
The process of $\mathcal{A}_{1}$ show as below:
$\mathcal{A}_{1}$ $=-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
$\quad\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\mathcal{N}(0,\mathbf{I})\rangle}_{\mathcal{C}_{0}}$
$=\mathcal{C}_{0}$
$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{1}}$
$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{2}}$
It’s obvious that $\mathbb{E}\mathcal{C}_{0}=0$.
By Cauchy-Schwarz inequality and AM-GM inequality, we have
$\mathcal{C}_{1}$ $=-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
$\leq\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|^{2}$
$=\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\mathcal{B}_{1}$
So we get
$\mathbb{E}\mathcal{C}_{1}\leq\sum_{i=1}^{N}p_{i}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}$
According to Assumption 2, we know that
$-\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\nabla F_{i}(\mathbf{w}_{t}^{i})\rangle\leq-\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
So we get
$\mathcal{C}_{2}$ $\leq 2\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}\cdot$
$\quad\left[-\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}\right]$
$\leq-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
$=-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}+F^{*}-F_{i}(\mathbf{w}^{*})\right)$
$\quad-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
$=-2\eta C_{t}\underbrace{\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}\right)}_{\mathcal{D}_{1}}-2\eta C_{t}\Gamma-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2},$
where $\Gamma\triangleq\sum_{i=1}^{N}p_{i}(F^{*}-F_{i}^{*})=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$.
Next, we proceed to handle $\mathcal{D}_{1}$.
$\mathcal{D}_{1}$ $=\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}_{t})\right)+\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t})-F^{*}\right)$
$\geq\sum_{i=1}^{N}p_{i}\langle\nabla F_{i}(\mathbf{w}_{t}),\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\rangle+(F(\mathbf{w}_{t})-F^{*})$
(from the Assumption 2)
$\geq-\frac{1}{2}\sum_{i=1}^{N}p_{i}\left[\eta\|\nabla F_{i}(\mathbf{w}_{t})\|^{2}+\frac{1}{\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
$\quad+(F(\mathbf{w}_{t})-F^{*})$
(from the AM-GM inequality)
$\geq-\sum_{i=1}^{N}p_{i}\left[\eta L(F_{i}(\mathbf{w}_{t})-F_{i}^{*})+\frac{1}{2\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
$\quad+(F(\mathbf{w}_{t})-F^{*})$
(from the L-smooth inference)
$\geq-(\eta L+1)\Gamma-\frac{1}{2\eta}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2},$
where L-smooth inference as show:
$\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$ (21)
Thus, we get
$\mathcal{C}_{2}\leq 2\eta^{2}C_{t}L\Gamma+C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
To sum up,
$\mathbb{E}\mathcal{A}_{1}$ $=\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}+2\eta^{2}C_{t}L\Gamma$
$\quad+C_{t}\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$=(1+C_{t})\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$\quad+\eta^{2}\left(G^{2}C_{t}^{2}+2L\Gamma C_{t}\right),$
and
$\quad\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}$
$=\mathbb{E}\sum_{i=1}^{N}p_{i}\|(\mathbf{w}_{t}-\mathbf{v}_{t+1}^{i})-(\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i})\|^{2}$
$\leq\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i}\|^{2}$
$\leq\mathbb{E}\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}$
$=\mathbb{E}\mathcal{A}_{2}=\eta^{2}C_{t}^{2}G^{2}+\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
So we get,
$\mathbb{E}\mathcal{A}_{1}$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}$
$\quad+\eta^{2}\left(2G^{2}C_{t}^{2}+2L\Gamma C_{t}+2\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}\right)$
All in all, we get
$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
where
$A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
FedAvg 98.67 |||| 7.3e-3 98.37 |||| 1.2e-3 89.41 |||| 3.6e-2 62.03 |||| 1.1e0 60.89 |||| 1.5e0 59.78 |||| 1.5e0
SCAFFLOD 98.45 |||| 5.0e-3 98.51 |||| 2.1e-3 86.39 |||| 3.4e-2 62.21 |||| 9.9e0 59.48 |||| 8.7e0 58.64 |||| 2.3e0
FedProx 98.70 |||| 7.5-3 98.35 |||| 1.2e-3 88.33 |||| 3.5e-2 62.13 |||| 1.2e0 60.33 |||| 2.2e0 61.31 |||| 1.6e0
FedDyn 98.40 |||| 8.0e-3 98.36 |||| 1.8e-3 88.26 |||| 3.5e-2 62.33 |||| 9.4e-1 61.35 |||| 2.6e0 62.52 |||| 1.5e0
q-FFL 98.72 |||| 4.7e-3 98.61 |||| 1.1e-3 89.03 |||| 3.6e-2 63.33 |||| 9.4e-1 60.63 |||| 1.3e0 59.65 |||| 1.5e0
FedFair 98.75 |||| 3.9e-3 98.72 |||| 1.2e-3 89.60 |||| 1.7e-2 62.38 |||| 8.5e-1 61.58 |||| 1.1e0 62.90 |||| 1.5e0
Table 4: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
FedAvg 93.40 |||| 1.1e11 94.67 |||| 4.1e9 75.60 |||| 3.7e9 52.61 |||| 6.1e9 51.58 |||| 2.0e9 55.90 |||| 1.5e9
SCAFFLOD 93.95 |||| 7.0e10 94.06 |||| 1.9e9 72.56 |||| 3.6e9 53.85 |||| 6.2e9 54.34 |||| 2.9e8 53.98 |||| 2.2e9
FedProx 90.50 |||| 5.3e10 94.23 |||| 3.1e9 71.38 |||| 5.6e9 51.34 |||| 4.4e9 53.67 |||| 4.5e8 54.14 |||| 2.2e9
FedDyn 91.42 |||| 5.6e10 93.08 |||| 3.8e9 73.21 |||| 9.8e10 52.14 |||| 5.5e9 50.21 |||| 1.2e10 51.86 |||| 2.0e10
ALI-DPFL 90.89 |||| 3.3e10 93.98 |||| 2.6e9 77.66 |||| 2.5e10 52.05 |||| 3.6e9 38.35 |||| 2.4e10 33.46 |||| 1.6e10
q-FFL 93.74 |||| 7.8e10 93.86 |||| 2.5e8 82.20 |||| 7.4e8 48.46 |||| 4.7e9 51.12 |||| 2.7e8 49.83 |||| 1.1e7
FedFDP 95.13 |||| 2.3e10 94.56 |||| 3.4e9 83.77 |||| 6.0e8 54.21 |||| 2.6e9 54.52 |||| 2.2e8 56.49 |||| 5.3e8
Table 5: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
A.2 Proof of Theorem 2
As Algorithm 2 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$. $\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of local client, so they are both to perform differential privacy preservation, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively. We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
A.2.1 Privacy Loss of $\mathbf{w}_{t+1}^{i}$
Theorem 3.
After $T$ rounds local updates, the RDP of the $\mathbf{w}_{T}^{i}$ in $i$-th client satisfies:
$R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$ (22)
where $\sigma$ is noise multiplier of the $w_{t+1}^{i}$, and $\alpha>1$ is the order.
$proof.$ We will prove this in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
Definition 2.
(RDP privacy budget of SGMMironov et al. [2019]). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$ (23)
where
$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$ (24)
with $\vartheta_{0}\triangleq\mathcal{N}\left(0,\sigma^{2}\right),\vartheta_{1}\triangleq\mathcal{N}\left(1,\sigma^{2}\right)\mbox{ and }\vartheta\triangleq(1-q)\vartheta_{0}+q\vartheta_{1}$
Further, it holds $\forall(q,\sigma)\in(0,1],\mathbb{R}^{+*},A_{\alpha}(q,\sigma)\geq B_{\alpha}(q,\sigma)$. Thus, $SG_{q,\sigma}\text{ satisfies }\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
Finally, the existing work Mironov et al. [2019] describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$ (25)
Definition 3.
(Composition of RDPMironov [2017]). For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
From Definition 2 and Definition 3, the Theorem 3 is obtained.
A.2.2 Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$
Theorem 4.
After $T$ rounds local updates, the RDP of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ in $i$-th client satisfies:
$R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$ (26)
where $\sigma_{l}$ is noise multiplier of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$, and $\alpha>1$ is the order.
The proof is similar to that of the $w_{t+1}^{i}$, so we omit it.
A.2.3 Proof Loss of FedFDP
Since both $w_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 3, and then use Lemma 2 to convert it to $(\epsilon,\delta)$-DP. Therefore, the Theorem 2 is hold.
Lemma 2.
(Conversion from RDP to DP Balle et al. [2020]). if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
A.3 Experiment Settings
A.3.1 Parameters Settings
For all datasets, we adopted a widely used heterogeneous settingsLin et al. [2020]Li et al. [2021a] to 10 clients which controlled by a Dirichlet distribution denoted as $Dir(\beta)$, where a smaller $\beta$ indicates a higher degree of heterogeneity. We set the default value of $\beta=0.1$ Lin et al. [2020]Wang et al. [2020].
In all datasets, we use a 4-layer CNN architecture McMahan et al. [2017] which consists of two convolutional layers and two fully connected layers. For FedFair and its baseline algorithms, we set learning rate $\eta=0.1$. For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$. In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
In all experiments, we use 25% of the data will form the test dataset, and the remaining 75% of the data will be used for training. All tasks will be repeated five times using different random seeds, and the mean and standard deviation will be reported.
A.3.2 Baselines introduction
The baseline methods included in the comparison are:
  • FedAvgMcMahan et al. [2017]: FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFDP into the FedAvg algorithm.
  • SCAFFLODKarimireddy et al. [2020]: SCAFFOLD uses control variates (variance reduction) to correct for the ’client-drift’ in its local updates. It requires significantly fewer communication rounds and is not affected by data heterogeneity or client sampling.
  • FedProxLi et al. [2020a]: FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
  • FedDynAcar et al. [2021]: FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
  • ALI-DPFLLing et al. [2023]: ALI-DPFL dynamically obtains the optimal number of epochs for each round of communication by minimizing the convergence upper bound, thereby addressing the imbalance between local privacy budgets and communication resources.
  • q-FFLLi et al. [2020b]: q-FFL amplifies the loss values of clients, equivalent to giving larger update weights to poorly performing clients, in order to cater to the underperforming clients, uniforming model performance across multiple devices.
A.3.3 Herterogeneity and Scalability
As shown in Table 4 and 5, we conducted additional experiments on the MNIST, FashionMNIST, and CIFAR10 datasets to compare the performance of FedFDP with that of six other algorithms in terms of Heterogeneity and Scalability. The Heterogeneity experiments were mainly conducted in three heterogeneous scenarios, extending from the default $Dir(0.1)$ to include $Dir(0.5)$ and $Dir(1)$. Scalability was expanded from the default 10 clients to include 20 and 50 clients.
Heterogeneities. Tables 4 and 5 show that when confronted with diverse data distributions exhibiting varying levels of heterogeneity, both FedFair and FedFDP outperform baseline methods in accuracy while maintaining significant fairness. With the exception of the $Dir(0.5)$ scenario, where they fall slightly short of q-FFL, they respectively demonstrated fairness improvements of 17.0% and 50% for FedFair and 30.3% and 18.9% for FedFDP compared to the best baselines in $Dir(0.1)$ and $Dir(1.0)$.
Scalability. Table 4 demonstrates that, with accuracy mostly superior to baselines, FedFair exhibited fairness improvements of 9.5%, 15.4%, and 0.01% across three client counts. Table 5 indicates that as the number of clients increases, FedFDP outperformed the best baselines by 0.36%, 0.18%, and 0.59% in accuracy, except in the 50 clients scenario, where it demonstrated fairness improvements of 27.8% and 24.1%.

### v2

@@
-FedFDP: Federated Learning with Fairness and Differential Privacy
-Abstract
+FedFDP: Fairness-Aware Federated Learning with Differential Privacy
+Abstract.
 Federated learning (FL) is a new machine learning paradigm to overcome the challenge of data silos and has garnered significant attention.
@@
 On the other hand, relevant studies indicate that the transmission of gradients or models in federated learning can also give rise to privacy leakage issues, such as membership inference attacks.
-To address the first issue mentioned above, we propose a federated algorithm with fairness, termed FedFair.
+To address the first issue mentioned above, we propose a fairness-aware federated learning algorithm, termed FedFair.
 Building upon FedFair, we introduce privacy protection to form the FedFDP algorithm to address the second issue mentioned above.
@@
 Lastly, extensive experimental results demonstrate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness.
-The code is accessible at https://anonymous.4open.science/r/FedFDP-E754.
-1 Introduction
-Federated Learning (FL) McMahan et al.
-(2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without revealing their individual datasets.
+Code and data is accessible at https://anonymous.4open.science/r/FedFDP-5607.
+Federated Learning, Differential Privacy, Fairness
+copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
+1. Introduction
+Federated Learning (FL) (McMahan et al., 2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without revealing their individual datasets.
 After training their local models, each client transmits only the model parameters, not the raw training data, to a central server.
 The server aggregates these parameters to update the global model, which is then sent back to clients for further training.
-Because its ability to overcome data silos by jointly improving model performance without the need to upload local data, FL has gained significant attention in recent year Zhu et al.
-(2021)Lee et al.
-(2022).
-In particular, many studies have made great breakthroughs in model performance Wang et al.
-(2020); Li et al.
-(2020a, 2021b) and communication costs Wang et al.
-(2019); Ling et al.
-(2023) in FL.
-However, recent research indicates that trained parameters can leak data privacy, such as the recovery of original data Zhu et al.
-(2019) or the membership inference attack Song et al.
-(2017).
+Thanks to its ability to overcome data silos by jointly improving model performance without the need to upload raw data, FL has gained significant attention in recent years (Zhu et al., 2021)(Lee et al., 2022).
+In particular, many studies have made great breakthroughs in model performance (Wang et al., 2020; Li et al., 2020b, 2021b) and communication costs (Wang et al., 2019; Ling et al., 2023) in FL.
+Figure 1. The test accuracies of 50 clients utilizing a globally trained model in FL.
+[…]
+To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary (Li et al., 2021c).
+Current machine learning algorithms that protect fairness, such as  (Dwork et al., 2012; Cotter et al., 2019), have not adequately addressed the issue of performance disparities among clients in federated learning.
+As shown in Figure 1, the red dots represent the test accuracies of the global model on different clients after completing federated learning with the FedAvg algorithm (McMahan et al., 2017).
+The global model trained collaboratively by clients exhibits significant unbalanced performance across different datasets held by the various clients.
+An increasing number of scholars are beginning to pay attention to this performance disparities phenomenon and are starting to work on it  (Li et al., 2020c, 2021b; Zhao and Joshi, 2022).
+On the other hand, recent research indicates that trained parameters can leak data privacy, such as the recovery of original data (Zhu et al., 2019) or the membership inference attack (Song et al., 2017).
 Moreover, Melis et al.
-Melis et al.
-(2022) demonstrated that clients’ training data can be leaked through shared models.
-In order to enhance data privacy in the FL, numerous scholars have begun integrating techniques such as differential privacy (DP) Dwork et al.
-(2014), homomorphic encryption Paillier (1999), and secure multi-party computation Yao (1982) into the federated learning framework.
-Among these, DP has emerged as the most widely acknowledged approach for privacy protection in federated learning, owing to its minimal computational overhead and rigorous mathematical framework.
-Many works Abadi et al.
-(2016); Carlini et al.
-(2019); Feldman (2020) have demonstrated that by adding proper DP noise in the training phase, the resulted machine learning models can prevent unintentional leakage of private training data, such as membership inference attacks.
-Figure 1: The test accuracy of 50 clients utilizing a globally trained model in FL.
-Another easily overlooked critical issue in federated learning is fairness.
-The globally trained model collaboratively by clients may exhibit significant performance disparities across datasets held by different clients.
-As shown in Fig.1, where the red dots indicate the test accuracy of the global model across various clients after the completion of federated learning using the FedAvg McMahan et al.
-(2017) algorithm.
-And we can see the globally trained model applied to individual clients, the performance significant performance disparities across differential clients.
-More and more scholars have begun to notice this unfair phenomenon and initiated research on it  Wang et al.
-(2020); Li et al.
-(2020b).
-While numerous articles separately discuss fairness and differential privacy protection in federated learning, there is a lack of research integrating both aspects into federated learning.
-Balancing model performance, fairness, and differential privacy protection in federated learning presents a challenging problem.
+(Melis et al., 2022) demonstrated that clients’ training data can be leaked through shared models.
+In order to enhance data privacy in FL, numerous scholars have begun integrating techniques such as differential privacy (DP) (Dwork et al., 2014), homomorphic encryption (Paillier, 1999), and secure multi-party computation (Yao, 1982) into the federated learning framework.
+Among these, DP stands out as the premier approach to privacy preservation in federated learning, attributed to its modest computational overhead and robust mathematical underpinnings.
+This is particularly relevant in defending membership inference attacks, as several studies (Abadi et al., 2016; Carlini et al., 2019; Feldman, 2020) have shown that incorporating appropriate DP during the training phase can effectively prevent inadvertent disclosure of sensitive training information.
+Although numerous articles have addressed fairness and differential privacy individually within the context of federated learning, there is a scarcity of research that integrates these two aspects.
+Achieving a balance between model performance, fairness, and differential privacy protection within federated learning frameworks poses a significant challenge.
 This paper aims to investigate fairness and differential privacy protection mechanisms in federated learning.
-By analyzing the relationship between them, we propose a federated learning framework that ensures both fairness and differential privacy protection, effectively guaranteeing model performance.
-Our main contributions are listed below:
-  1. 1.
-    Firstly, we propose a fairness-aware federated algorithm, denoted as algorithm FedFair by designing a new local loss function that is constructed to simultaneously optimize the two objectives of fairness and model performance (Section 4).
-FedFair algorithm achieves the excellent fairness, as shown by the green dots in Fig.1.
-  2. 2.
-    Building upon FedFair, we propose algorithm FedFDP to achieve differential privacy protection (Section 5) .
+By analyzing the relationship between them, we propose a fairness-aware federated learning with differential privacy, effectively guaranteeing model performance when achieving fairness and privacy protection.
+Our main contributions are listed as follows:
+  1. (1)
+    Firstly, a fairness-aware federated algorithm is proposed, denoted as FedFair.
+It incorporates a novel loss function, designed to optimize the objectives of fairness and model performance simultaneously (Section 4).
+FedFair algorithm achieves the excellent fairness, as shown by the green dots in Figure1.
+  2. (2)
+    Building upon FedFair, an algorithm referred to as FedFDP is proposed to achieve differential privacy protection (Section 5).
 Particularly, we device a fairness-aware gradient clipping strategy to ensures both differential privacy and allows for the adjustment of fairness in federated learning (Section 5.1).
 Additionally, we propose an adaptive clipping approach method for the additional loss values uploaded by each client to achieve optimal utility (Section 5.2).
-  3. 3.
-    Furthermore, we demonstrate there exist an optimal fairness parameter $\lambda$ value to balance the model performance and fairness through the convergence analysis (Section 6).
-  4. 4.
-    And then, we perform privacy loss analysis of our algorithm FedFDP by the notion of Rényi differential privacy (RDP) to prove that it satisfies the differential privacy (Section 7).
-  5. 5.
-    Lastly, through a comprehensive empirical evaluation on three public datasets, we confirm that FedFair and FedFDP outperforms the state-of-the-art solutions in terms of model performance and fairness (Section 8).
-2 Related Work
-2.1 Fairness in Federated Learning
-In the field of machine learning, a multitude of fairness concepts have been advanced (Mehrabi et al.
-Mehrabi et al.
-(2021), Yuxin et al.
-Shi et al.
-(2021)), each highlighting specific facets and addressing the concerns of diverse stakeholder groups.
-For instance, Mohri et al.
-Mohri et al.
-(2019) advocate for Good-Intent Fairness, while Zhou et al.
-Zhou et al.
-(2021) embrace Selection Fairness.
-Cong and Lyu et al.
-Cong et al.
-(2020)Lyu et al.
-(2020) employ principles from game theory Rabin (1993) to steer the allocation of rewards.
-Currently, the predominant fairness criterion in Federated Learning is “Performance Parity”, as proposed by Li et al.
-Li et al.
-(2020b).
-This criterion aims to ensure fairness by evaluating the consistency of performance across different client devices in federated learning.
-This definition is widely adopted.
-Zhao et al.
-Zhao and Joshi (2022) suggested replacing the loss amplification mechanism in Li et al.
-(2020b) with a straightforward weight reallocation mechanism.
-Li et al.
-Li et al.
-(2021b) focus on optimizing the fairness of local models in personalized federated learning.
-2.2 Federated Learning with Differential Privacy
-Federated learning with differential privacy has become a common privacy protection method in federated learning.
-Enayat et al.
-Ullah et al.
-(2023) dynamically adjust the compression rate to minimize communication rounds in federated learning by responding to changes in training loss.
-Zhou et al.
-Zhou et al.
-(2023) investigated the effects of local iterations when utilizing the Laplace mechanism for perturbation.
-Fu et al.
-Fu et al.
-(2022) introduced adaptive noise adding, customizing it for different clients and communication rounds based on observed fluctuations in loss values.
-Ling et al.
-Ling et al.
-(2023) determined the optimal number of local epochs through convergence analysis, striking a balance between communication resources and privacy budget.
-Raouf et al.
-Kerkouche et al.
-(2021) improved model accuracy by updating only a fixed subset of model weights while keeping others unchanged.
-The aforementioned works each independently address either fairness or privacy protection in federated learning.
-To the best of our knowledge, we are the first to concurrently consider both fairness and privacy in the context of federated learning.
-3 Preliminaries
-3.1 Federated Learning
+  3. (3)
+    Furthermore, it is demonstrated that an optimal fairness parameter $\lambda$ is found, which, when determined, enables the fastest convergence of FedFDP in balancing model performance and fairness through the convergence analysis (Section 6).
+  4. (4)
+    And then, a privacy loss analysis is conducted on algorithm FedFDP using the notion of Rényi differential privacy (RDP) to verify its satisfaction of differential privacy requirements (Section 7).
+  5. (5)
+    Lastly, through a comprehensive empirical evaluation on three public datasets, it is confirmed that FedFair and FedFDP outperform the state-of-the-art solutions in terms of model performance and fairness (Section 8).
+2. Related Work
+Differentially private federated learning (DPFL) is widely used for privacy protection in federated learning.
+Many studies have focused on the trade-off between model performance, communication efficiency, and privacy budget in DPFL (Ullah et al., 2023; Zhou et al., 2023; Fu et al., 2022; Ling et al., 2023; Kerkouche et al., 2021; Fu et al., 2024).
+In addition that, various fairness concepts in machine learning have been proposed, each emphasizing different aspects and addressing the needs of different stakeholders.
+For example, concepts like Good-Intent Fairness (Mohri et al., 2019) and Selection Fairness (Zhou et al., 2021) have been introduced.
+Further discussions on these concepts can be found in additional literature (Cong et al., 2020; Lyu et al., 2020; Mehrabi et al., 2021; Shi et al., 2021).
+At present, the primary criterion for fairness in FL is group fairness, which aims to mitigate biases associated with sensitive attributes like gender and race.
+Relevant research has shown that in federated learning with differential privacy, adjusting parameters such as the clipping threshold and noise multiplier can help enhance the algorithm’s group fairness (Gu et al., 2022).
+There is a growing body of literature that addresses the concept of group fairness specifically within the framework of DPFL (Gálvez et al., 2021; Padala et al., 2021; Gu et al., 2022, 2020; Pentyala et al., 2022).
+The more predominant fairness criterion in FL is balanced performance fairness, as introduced by Li et al.
+(Li et al., 2020c).
+This approach ensures fairness by assessing the uniformity of model performance across participating clients.
+The concept has gained broad acceptance.
+Building on this, Zhao et al.
+(Zhao and Joshi, 2022) proposed an alternative to the loss amplification mechanism, advocating for a direct weight reallocation strategy.
+Concurrently, Li et al.
+(Li et al., 2021b) concentrated on enhancing the fairness of local models within personalized FL.
+Our study pioneers the investigation of balanced performance fairness within the context of DPFL.
+In comparison to definitions such as group fairness, balanced performance fairness is more meaningful for data providers such as enterprises and institutions, moreover, it can effectively motivate clients participate in the federated system.
+However, achieving balanced performance fairness typically requires adjusting the direction of model updates.
+It is a challenge to refine the model for balanced performance fairness while integrating DP techniques, such as clipping and noise addition, into the model’s update process.
+To the best of our knowledge, we are the first to consider the work of balanced performance fairness in DPFL.
+3. Preliminaries
+In the beginning, Table 1 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
+Table 1. Summary of main notations
+$F(\mathbf{w}$) Global loss function
+$F_{i}(\mathbf{w}$) Local loss function for client $i$
+$T$ Communication rounds
+$t$ Communication round index
+$D_{i}$ Dataset of client $i$
+$D$ The union of $D_{i}$
+$\mathbf{w}_{t}^{i}$ Local model parameter at node $i$ at round $t$
+$\mathbf{w}^{*}$ Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
+$\eta$ Gradient descent step size
+$\lambda$ Fairness parameter
+$N$ Total number of clients
+$q$ batch sample ratio
+$|\mathcal{B}_{i}|$ Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
+$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
+$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
+$\sigma$ Noise multiplier for gradient
+$\sigma_{l}$ Noise multiplier for loss
+$C$ Original clipping bound for gradient
+$C_{l}$ Original Clipping bound for loss
+|||||$\cdot$||| | ⋅ | | $L_{2}$-norm
+3.1. Federated Learning and Fairness
 We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model.
-The objective of Federated Learning is to minimize $F(\mathbf{w})$, in other words, to seek:
-$\mathbf{w}^{*}\triangleq\arg\min F(\mathbf{w}).$ (1)
-The function $F(\cdot)$ is defined as follows:
-$F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (2)
+The objective of FL is to minimize $F(\mathbf{\cdot})$, in other words, to seek:
+(1) $\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$
 Here, $F_{i}(\cdot)$ represents the local loss function of client $i$.
 The weights are defined as $p_{i}=\frac{|D_{i}|}{|D|}$, $\sum_{i=1}^{N}p_{i}=1$, $D_{i}$ represents the dataset of client $i$, and $D$ represents the union of all $D_{i}$.
-3.2 DPSGD
-In gradient-based methods, DPSGDAbadi et al. (2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
+We have defined balanced performance fairness in FL as follows.
+Specifically, this type of fairness is defined based on the performance variance of the model across clients: the smaller the variance, the fairer the model.
+Definition 0.
+(Balanced performance fairness (Li et al., 2020c)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
+3.2. Differential Privacy
+Differential privacy is a formal mathematical model that quantifies privacy leakage in data analysis algorithms.
+It stipulates that alterations to a single record within the training data should not induce significant statistical variations in the algorithm’s results.
+Definition 0.
+(Differential Privacy (Dwork et al., 2014)).
+$(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
+(2) $\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$
+By adding random noise, we can achieve differential privacy for a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ according to Definition 2.
+The $l_{k}$-sensitivity determines how much noise is needed and is defined as follow.
+We use $l_{2}$ sensitivity as in the previous work.
+Definition 0.
+($\mathbf{l_{k}}$-Sensitivity(Dwork et al., 2006)) For a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, we define its $l_{k}$ norm sensitivity (denoted as $\Delta_{k}f$) over all neighboring datasets $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ differing in a single sample as
+(3) $\text{sup}_{D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}}||f(D_{i})-f(D_{i}^{\prime})||_{k}\leq\Delta_{k}f.$
+DPSGD. In gradient-based methods, DPSGD(Abadi et al., 2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
+  1. (1)
     Firstly, computing the gradient of loss function:
-    $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}).$ (3)
+    (4) $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}).$
+  2. (2)
     Secondly, clipping the gradient of per-sample to get the sensitivity:
-    $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right).$ (4)
-  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients:
-    $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$ (5)
+    (5) $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right),$
+    where $C$ represents the clipping bound.
+  3. (3)
+    Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier:
+    (6) $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$
 The final step is to proceed with gradient descent as usual.
-Here, $C$ represents the clipping bound, and $\sigma$ is the noise multiplier.
-If each client in FL uses DPSGD as the gradient descent method to protect their private data, FL will transform into Differentially Private Federated Learning (DPFL).
-3.3 Fairness in Federated Learning
-Definition 1.
-(Fairness of performance distributionLi et al.
-(2020b)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally say that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
-In this study, we employ the loss of model denote the performance.
-Equation (6) represents the specific form of fairness in federated learning in this paper, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients.
-Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients.
-$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (6)
+When each client in FL utilizes DPSGD as the optimization algorithm to safeguard their private data, the framework transitions from FL to DPFL.
 Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, fairness hyperparameter $\lambda$, batch sample ratio $q$, local dataset $D_{i}$.
@@
 1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
-2SERVER
+2 SERVER
 3       for $t\in[0,1,\cdots,T-1]$  do
@@
 13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
+14
 15      $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
-15       for $e\in[1,\cdots,E]$ do
 16       for batch $b\in\mathcal{B}_{i}$ do
@@
 20
-21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ and compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
+21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ , compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
 22       return $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})$;
@@
 Algorithm 1 FedFair
-4 Proposed Algorithm: FedFair
-In the previous section, we identified the performance objective function as shown in Equation (2) and the fairness objective function as shown in Equation (6).
-Both of these objective functions are composed of individual client objectives.
-In this section, we will focus on the performance and fairness of each client and combine these two objectives to construct a new objective function, based on which we propose a fair federated learning algorithm.
-We simultaneously optimize performance and fairness, and can get the optimization objective for client $i$ as:
-$\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}_{t})\right)^{2},$ (7)
-where $F(\mathbf{w}_{t})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t-1})$ [^1] The reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training.and $\lambda$ is a hyperparameter used to adjust the degree of fairness, $\lambda\geq 0$.
-When $\lambda=0$, the algorithm becomes traditional federated learning.
+4. Proposed Algorithm: FedFair
+We previously defined the performance objective in Equation (1) and balanced performance fairness in Definition 1.
+Here, we use these to develop a new objective function that informs our proposed fairness-aware federated learning algorithm.
+In this study, we employ the loss of model denote the performance.
+Equation (7) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients.
+Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
+(7) $\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$
+Upon incorporating Equation (7) into Equation (1), we introduce a comprehensive objective function that integrates considerations of fairness:
+(8) $\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$
+For client $i$ during round $t$, the refined objective is articulated as follows:
+(9) $\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$
+where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training.
+from round $t-1$, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning.
 When $\lambda\rightarrow\infty$, the algorithm only focus on fairness.
-The gradient of Equation (7) as follows:
-$\nabla H_{i}(\mathbf{w}_{t}^{i})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})$
-$=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (8)
-where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t})$.
-The stochastic gradient descent for Equation (7) as show:
-$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})$
-$=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (9)
+However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging.
+Fortunately, we theoretically found an optimal value for $\lambda$ in section 6.2.
+In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$..
+Consequently, within the context of Equation (9), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed.
+So, the gradient of Equation (9) as follows:
+(10) $\nabla H_{i}(\mathbf{w}_{t}^{i})=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$
+where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t}^{i})$.
+The stochastic gradient descent by using Equation (10) as show:
+(11) $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$
 where $\eta_{i}=\eta(1+\lambda\Delta_{i})$.
+Figure 2. FedFDP framework.
+[…]
 Therefore, our approach can be viewed as establishing an adaptive learning rate that allows the model update to achieve both performance and fairness objectives simultaneously.
@@
   • Server (lines 1-1): The server executes $T$ rounds of communication, each round sending the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
-  • Local (lines 1-1): For $E$ epochs and each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
-Furthermore, as the iterations advance, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$.
-In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods Zinkevich (2003).
-Figure 2: FedFDP framework
-5 Proposed Algorithm: FedFDP
+  • Local (lines 1-1): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
+Furthermore, as the training process advances, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$.
+In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods (Zinkevich, 2003).
+5. Proposed Algorithm: FedFDP
 In this section, we propose the FedFDP by integrating differential privacy protection base on FedFair.
-The flow of FedFDP algorithm as shown in Fig.
-2.
+The flow of FedFDP algorithm as shown in Figure 2.
 The FedFDP algorithm adds the following two additional processes to FedFair.
-Firstly, in Section 5.1, we present a fair-clipping strategy that not only achieves sensitivity but also allows for adjustments to the level of fairness.
+Firstly, in Section 5.1, we present a fair-clipping strategy that not only achieves sensitivity but also allows adjustments to the level of fairness.
 Furthermore, in Section 5.2, we expand the use of differential privacy protection to the additional loss values that local clients need to upload, suggesting an adaptive clipping method to maximize utility.
 The overall FedFDP algorithm is described in Section 5.3.
-5.1 Fair-clipping Strategy for Gradient
-In the general DPSGD algorithm Fu et al.
-(2023); Abadi et al.
-(2016); Bu et al.
-(2022); Yang et al.
-(2022), per-sample clipping is an essential process for obtaining sensitivity.
-In order to be compatible with per-sample clipping, we expand our objective function of client $i$ in FedFair, Equation (7), into Equation (10):
-$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})\right)^{2},$ (10)
-where $\xi_{j}\in\mathcal{B}_{t}^{i}$, $j=\{1,2,\cdots,|\mathcal{B}_{i}|\}$, and $|\mathcal{B}_{i}|$ is the batch size of client $i$.
-Next, we compute the gradient of Equation (10):
-$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\lambda\cdot\Delta_{loss}^{j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$
-$=\left(1+\lambda\cdot\Delta_{loss}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$ (11)
-where $\Delta_{loss}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$.
-To achieve differential privacy protection, we replace loss function gradient by Equation (5.1) based on Equations (3)-(5), and reorganize the gradient descent equation for client $i$:
-$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta^{j}_{loss}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta^{j}_{loss}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
-$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
-$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (12)
-where:
-$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta^{j}_{loss},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$ (13)
-Hence, our strategy can be regarded as an fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model.
-When $\lambda=0$, Equation (13) corresponds to the traditional DP clipping method, and FedFDP no longer ensures fairness.
-For large $C$, Equation (13) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness.
-Conversely, for small $C$, Equation (13) simplifies to the second term, resulting in the gradient being clipped to the bound $C$.
-Our conclusions can be verified from the experiments 5.
-5.2 Adaptive Clipping Method for Loss
-Since the calculation of $\Delta_{loss}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy.
-As shown in lines 2-2 of Algorithm 2, We need to clip $F_{i}(\mathbf{w}_{t})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $C_{l}^{i,t}\cdot\sigma_{l}$ to ensure differential privacy.
-However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t})$ is challenging.
-A large $C_{l}^{i,t}$ helps avoid loss of fidelity to the original values, but it also leads to large injected noise.
-Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t})$.
-For the $i$-th client, the adaptive clipping method uses the $differentially$ $private$ $mean$ $of$ $the$ $previous$ $round$ as the clipping bound for the current round.
-The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
-$C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$ (14)
-where
-$\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))),$ (15)
-To ensure the achievement of differential privacy during the clipping process, we use the noise-added loss as the clipping bound $C_{l}^{i,t}$.
-Each round’s clipping bound adheres to the post-processing nature of differential privacy when $t=\{1,\dots,T\}$.
-5.3 Overall Algorithm
-We utilize the equation form presented in (12) to enhance Algorithm 2. Compared with the Algorithm 1, FedFDP solely necessitates the application of differential privacy to gradient and loss.
 Input: loss function $F(\mathbf{w})$.
-Parameters: learning rate $\eta$, noise multiplier for training $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
+Parameters: learning rate $\eta$, noise multiplier for gradient $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
 Output: the final trained model $\mathbf{w}_{T}$
-1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0}),C_{l}^{0}=\text{Initial()}$();
-2SERVER
+1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0}),C_{l}^{0}$;
+2 SERVER
 3       for $t\in[0,1,\cdots,T-1]$  do
@@
 15       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
-16             $\Delta_{loss}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
+16             $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
 17             Compute $C^{i,j}_{t}$ by Equation (15);
@@
 19
-20      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\mathcal{N}(0,\mathbf{I})\right)$;
+20      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right)$;
+21
 22      Compute clipping bound of loss $C_{l}^{i,t}$ by Equation (16) and (17) ;
@@
 24             Compute $F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})$;
+25
 26            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
-25
-26      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})\leftarrow\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\mathcal{N}(0,\mathbf{I})\right)$;
+27
+28      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})\leftarrow\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\cdot\mathcal{N}(0,\mathbf{I})\right)$;
 29       return $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
-28
+30
 Algorithm 2 FedFDP
+5.1. Fair-clipping Strategy for Gradient
+In the general DPSGD algorithm (Abadi et al., 2016; Fu et al., 2023; Bu et al., 2022; Yang et al., 2022), per-sample clipping is an essential process for obtaining sensitivity.
+In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (9), into Equation (12):
+(12) $\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$
+where $\xi_{j}\in\mathcal{B}_{t}^{i}$, $j=\{1,2,\cdots,|\mathcal{B}_{i}|\}$, and $|\mathcal{B}_{i}|$ is the batch size of client $i$.
+Next, we compute the gradient of Equation (12):
+$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\lambda\cdot\Delta_{i}^{j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$
+(13) $=\left(1+\lambda\cdot\Delta_{i}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$
+where $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})$.
+To achieve differential privacy protection, we replace the gradient of loss function by Equation (5.1) based on Equations (4)-(6), and reorganize the gradient descent equation for client $i$:
+$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
+$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
+(14) $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$
+where:
+(15) $C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
+Hence, our strategy can be regarded as an fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model.
+When $\lambda=0$, Equation (15) corresponds to the traditional DP clipping method, and FedFDP no longer ensures fairness.
+For large $C$, Equation (15) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness.
+Conversely, for small $C$, Equation (15) simplifies to the second term, resulting in the gradient being clipped to the bound $C$.
+Our conclusions can be verified from the experiments (Figure 4).
+5.2. Adaptive Clipping Method for Loss
+Since the calculation of $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy.
+As shown in lines 2-2 of Algorithm 2, We need to clip $F_{i}(\mathbf{w}_{t}^{i})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy.
+However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging.
+A large $C_{l}^{i,t}$ avoids loss of fidelity to the original values, but it also leads to large injected noise.
+Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t}^{i})$.
+For the $i$-th client, the adaptive clipping method uses the $differentially$ $private$ $mean$ $of$ $the$ $previous$ $round$ as the clipping bound for the current round.
+The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1}^{i})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
+(16) $C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$
+where
+(17) $\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))).$
+To ensure the achievement of differential privacy during the clipping process, we use the noise-added loss as the clipping bound $C_{l}^{i,t}$.
+Each round’s clipping bound adheres to the post-processing nature of differential privacy when $t=\{1,\dots,T\}$.
+5.3. Overall Algorithm
+We utilize the equation form Equation (14) to enhance Algorithm 2. Compared with the Algorithm 1, FedFDP solely necessitates the application of differential privacy to gradient and loss.
   • DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$.
@@
   • DP for local loss (lines 2-2): As described in 5.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
-6 Utility Analysis
-In this section, we examine the convergence of the FedFDP algorithm to identify an optimal $\lambda$ that minimizes the convergence bound.
-6.1 Convergence Analysis
+6. Optimal Fairness Parameter
+In this section, we examine the convergence of the FedFDP algorithm to identify an optimal fairness parameter $\lambda$ that minimizes the convergence bound.
+6.1. Convergence Analysis
 We analyze the convergence of Algorithm 2 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound.
@@
 Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random.
-The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=1,\cdots,N$ , $t=1,\cdots,T-1$ and $j=1,\cdots,|\mathcal{B}_{i}|$
+The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=[1,\cdots,N]$ , $t=[1,\cdots,T-1]$ and $j=[1,\cdots,|\mathcal{B}_{i}|]$
 Theorem 1.
 Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 2 as follows:
-$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (16)
+(18) $\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$
 where:
@@
   • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
-The meaning of $\Gamma$ is consistent with Li et al. (2020c), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
+The meaning of $\Gamma$ is consistent with (Li et al., 2020a), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
 proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$.
 Finally, by utilizing Assumption 1, we derive (18).
-For the detailed proof, please refer to Appendix A.1.
-6.2 Optimal $\lambda$
-When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{loss}^{j}$.
-We assume $Q_{0}\leq\Delta_{loss}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$.
-We rearrange the convergence upper bound in (16) as a function of $\lambda$,
-$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{2Q_{0}\lambda+1},$ (17)
-where $a_{1}=G^{2}Q_{1}^{3}$, $a_{2}=6G^{2}Q_{1}^{2}$, $a_{3}=9Q_{1}G^{2}+2L\Gamma Q_{1}+2Q_{0}\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{4}=4G^{2}+2L\Gamma+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$.
-We sketched a rough graph of $P(\lambda)$, where $\lambda\in[0,+\infty)$.
-Upon observing Fig.
-3, it becomes evident that when $a_{1}$,$a_{2}$,$a_{3}$,$a_{4}$ greater than 00, there exists an optimal value of $\lambda$ that minimizes the right-hand side of Equation (16), aligning with the experimental results presented in Fig.
-4.
-Figure 3: An example of $P(\lambda)$
-7 Privacy Analysis
-Because each client’s locally trained pair of data samples has the same sampling rate and noise multiplier in each round, each client has the same privacy loss in each round.
-As a result, we only assess each client’s privacy loss from the perspective of a single client.
-Following that, we examine the privacy loss of FedPDF for the $i$-th client as follows.
+For the detailed proof, please refer to Appendix A.1333Due to the lengthy proof, please refer to https://anonymous.4open.science/r/FedFDP-5607/APPENDIX/FedFDP_appendix.pdf for the appendix..
+6.2. Optimal $\lambda$
+Theorem 2.
+There exists a computable optimal $\lambda$, denoted as $\lambda^{*}$, which ensures the fastest convergence of the Algorithm 2.
+Proof.
+When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},x_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{i}^{j}$.
+Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$.
+We rearrange the convergence upper bound in Equation (18) to obtain a function in terms of $\lambda$,
+(19) $P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$
+where $a_{1}=G^{2}Q_{1}^{3}$, $a_{2}=6G^{2}Q_{1}^{2}$, $a_{3}=9Q_{1}G^{2}+2L\Gamma Q_{1}+2Q_{0}\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{4}=4G^{2}+2L\Gamma+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{5}=2Q_{0}$.
+Let $\mathcal{F}(\lambda)=\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1}$, then $P(\lambda)=\frac{L}{2\mu T}\mathcal{F}(\lambda)$, where $\frac{L}{2\mu T}$ can be considered a constant term.
+Obviously, when $\mathcal{F}(\lambda)$ is minimized, $P(\lambda)$ is also minimized.
+The derivative of $\mathcal{F}(\lambda)$ is as follows:
+$\mathcal{F}^{\prime}(\lambda)$ $=\frac{(3a_{1}\lambda^{2}+2a_{2}\lambda+a_{3})(a_{5}\lambda+1)}{(a_{5}\lambda+1)^{2}}-\frac{a_{5}(a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4})}{(a_{5}\lambda+1)^{2}}$
+(20) $=\frac{2a_{1}a_{5}\lambda^{3}+(a_{2}a_{5}+3a_{1})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}}{(a_{5}\lambda+1)^{2}}.$
+Let the denominator of $\mathcal{F}^{\prime}(\lambda)$ be a new function:
+(21) $\mathcal{G}(\lambda)=2a_{1}a_{5}\lambda^{3}+(3a_{1}+a_{2}a_{5})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}.$
+Since $(a_{5}\lambda+1)^{2}>0$, the sign of $\mathcal{F}^{\prime}(\lambda)$ is consistent with that of $\mathcal{G}(\lambda)$.
+Using the discriminant of a cubic function $\Delta=b^{2}-3ac$ (a general cubic function is expressed as $a\lambda^{3}+b\lambda^{2}+c\lambda+d$),
+$\Delta$ $=(3a_{1}+a_{2}a_{5})^{2}-3(2a_{1}a_{5}\cdot 2a_{2})$
+$=9a_{1}^{2}+a_{2}^{2}a_{5}^{2}+6a_{1}a_{2}a_{5}-12a_{1}a_{2}a_{5}$
+(22) $=(3a_{1}-a_{2}a_{5})^{2}\geq 0.$
+The derivative of $\mathcal{G}(\lambda)$ is given by
+(23) $\mathcal{G}^{\prime}(\lambda)=6a_{1}a_{5}\lambda^{2}+2(a_{2}a_{5}+3a_{1})\lambda+2a_{2}.$
+It is evident that when $\lambda>0$, $\mathcal{G}^{\prime}(\lambda)>0$.
+If $\mathcal{F}(\lambda)$ has an extremum, then $\mathcal{F}^{\prime}(\lambda)$ has a root, meaning $\mathcal{G}(\lambda)$ has a root.
+Substituting $\lambda=0$, $\mathcal{G}(0)=a_{3}-a_{4}a_{5}$, evidently when $a_{3}<a_{4}a_{5}$, $\mathcal{G}(\lambda)$ has a root.
+Let $\lambda=x-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, replacing $\lambda$ with $x$ in $\mathcal{G}(\lambda)$ yields
+(24) $\mathcal{G}(x)=x^{3}+a_{6}x+a_{7},$
+where
+  • $a_{6}=\frac{-(3a_{1}-a_{2}a_{5})^{2}}{12a_{1}^{2}a_{5}^{2}}$,
+  • $a_{7}=\frac{108a_{1}^{2}a_{5}^{2}(a_{3}-a_{4}a_{5})-36(a_{1}a_{2}a_{5})(3a_{1}+a_{2}a_{5})+2(3a_{1}+a_{2}a_{5})^{3}}{216a_{1}^{3}a_{5}^{3}}$.
+Using the root-finding formula to calculate $\mathcal{G}(x)=0$ gives the solutions $x_{1},x_{2},x_{3}$.
+[…]
+[…]
+[…]
+where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{1,2,3}=x_{1,2,3}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$.
+Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root.
+If the equation has $k$ real roots, the optimal solution for $\lambda$ is $\lambda^{*}=\max_{k}(\lambda_{k})$.
+In conclusion, $\mathcal{F}^{\prime}(\lambda)\geq 0\iff G(\lambda)\geq 0\iff\lambda\geq\lambda^{*}$, therefore $\mathcal{F}(\lambda)$ has a minimum value $\mathcal{F}(\lambda^{*})$, and $P(\lambda)$ has a minimum value $P(\lambda^{*})$.
+7. Privacy Analysis
+As Algorithm 2 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$.
+$\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of local client, so they are both to perform differential privacy preservation, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively.
+We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
+7.1. Privacy Loss of $\mathbf{w}_{t+1}^{i}$
+Theorem 1.
+After $T$ rounds local updates, the RDP of the $\mathbf{w}_{T}^{i}$ in $i$-th client satisfies:
+(25) $R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$
+where $\sigma$ is noise multiplier of the $\mathbf{w}_{t+1}^{i}$, and $\alpha>1$ is the order.
+The detail of proof please refer Appendix A.2.
+7.2. Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$
+Theorem 2.
+After $T$ rounds local updates, the RDP of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ in $i$-th client satisfies:
+(26) $R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$
+where $\sigma_{l}$ is noise multiplier of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$, and $\alpha>1$ is the order.
+The proof is similar to that of the $\mathbf{w}_{t+1}^{i}$, so we omit it.
+7.3. Privacy Loss of FedFDP
+Since both $\mathbf{w}_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 3, and then use Lemma 4 to convert it to $(\epsilon,\delta)$-DP.
+Lastly, we can get the Privacy loss of FedFDP as follows.
 Theorem 3.
@@
 […]
-where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 3 in Appendix A.2, and $R_{loss}^{i}(\alpha$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 4 in Appendix A.2.
-The details of the proof please refer to Appendix A.2.
+where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 1, and $R_{loss}^{i}(\alpha)$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 2.
+Lemma 0.
+(Conversion from RDP to DP (Balle et al., 2020)).
+if a randomized mechanism $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ satisfies $(\alpha,R)$-RDP , then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
+8. Experiments
+Table 2. FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
 Datasets MNIST FashionMNIST CIFAR10
@@
 FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
-Table 1: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
+Table 3. FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
 Datasets MNIST FashionMNIST CIFAR10
@@
 FedAvg 93.40$\pm$0.47 1.1e11$\pm$0.5e10 84.15$\pm$1.97 3.0e8$\pm$1.4e8 52.61$\pm$0.17 6.1e9$\pm$5.2e9
-SCAFFLOD 93.95$\pm$1.39 7.0e10$\pm$4.9e9 79.41$\pm$4.87 3.3e9$\pm$4.7e9 53.85$\pm$1.16 6.2e9$\pm$1.3e9
+SCAFFLOD 93.95$\pm$1.39 7.0e10$\pm$4.9e9 83.41$\pm$4.87 3.3e9$\pm$4.7e9 53.85$\pm$1.16 6.2e9$\pm$1.3e9
 FedProx 90.50$\pm$3.95 5.3e10$\pm$7.3e9 83.85$\pm$0.98 7.9e8$\pm$4.5e8 51.34$\pm$0.91 4.4e9$\pm$2.1e9
@@
 FedFDP 95.13$\pm$0.83 2.3e10$\pm$1.0e10 85.99$\pm$0.76 2.8e8$\pm$0.8e8 54.21$\pm$0.98 2.6e9$\pm$1.6e9
-Table 2: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
-8 Experiments
-In this section, we conduct experiments to demonstrate the performance and fairness of FedFair and FedFDP using three real datasets.
-Additionally, we investigate the impact of various factors on FedFDP, such as fairness hyperparameter $\lambda$, original clipping bound $C$, heterogeneity and scalability.
-The impact of heterogeneity and scalability please refer to Appendix A.3).
-8.1 Default Settings
-We conducted experiments on three computer vision (CV) datasets: MNIST LeCun et al.
-(1998), FashionMNIST Xiao et al.
-(2017), and CIFAR10 Krizhevsky (2009).
-And we comparised our algorithms to six baselines, which include FedAvgMcMahan et al.
-(2017), SCAFFLODKarimireddy et al.
-(2020), FedProxLi et al.
-(2020a), FedDynAcar et al.
-(2021), ALI-DPFLLing et al.
-(2023) and q-FFLLi et al.
-(2020b).
-For the details of parameters settings and the descriptions of baselines please refer to Appendix A.3.1.
-8.2 Results
+Table 4. FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
+Herterogeneity Scalability
+Datasets MNIST FashionMNIST CIFAR10
+Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
+FedAvg 98.67 |||| 7.3e-3 98.37 |||| 1.2e-3 89.41 |||| 3.6e-2 62.03 |||| 1.1e0 60.89 |||| 1.5e0 59.78 |||| 1.5e0
+SCAFFLOD 98.45 |||| 5.0e-3 98.51 |||| 2.1e-3 86.39 |||| 3.4e-2 62.21 |||| 9.9e0 59.48 |||| 8.7e0 58.64 |||| 2.3e0
+FedProx 98.70 |||| 7.5-3 98.35 |||| 1.2e-3 88.33 |||| 3.5e-2 62.13 |||| 1.2e0 60.33 |||| 2.2e0 61.31 |||| 1.6e0
+FedDyn 98.40 |||| 8.0e-3 98.36 |||| 1.8e-3 88.26 |||| 3.5e-2 62.33 |||| 9.4e-1 61.35 |||| 2.6e0 62.52 |||| 1.5e0
+q-FFL 98.72 |||| 4.7e-3 98.61 |||| 1.1e-3 89.03 |||| 3.6e-2 63.33 |||| 9.4e-1 60.63 |||| 1.3e0 59.65 |||| 1.5e0
+FedFair 98.75 |||| 3.9e-3 98.72 |||| 1.2e-3 89.60 |||| 1.7e-2 62.38 |||| 8.5e-1 61.58 |||| 1.1e0 62.90 |||| 1.5e0
+Table 5. FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
+Herterogeneity Scalability
+Datasets MNIST FashionMNIST CIFAR10
+Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
+FedAvg 93.40 |||| 1.1e11 94.17 |||| 4.1e9 85.60 |||| 3.7e9 52.61 |||| 6.1e9 51.58 |||| 2.0e9 55.90 |||| 1.5e9
+SCAFFLOD 93.95 |||| 7.0e10 94.06 |||| 1.9e9 84.56 |||| 3.6e9 53.85 |||| 6.2e9 54.34 |||| 2.9e8 53.98 |||| 2.2e9
+FedProx 90.50 |||| 5.3e10 94.23 |||| 3.1e9 84.38 |||| 5.6e9 51.34 |||| 4.4e9 53.67 |||| 4.5e8 54.14 |||| 2.2e9
+FedDyn 91.42 |||| 5.6e10 93.08 |||| 3.8e9 85.21 |||| 9.8e10 52.14 |||| 5.5e9 50.21 |||| 1.2e10 51.86 |||| 2.0e10
+ALI-DPFL 90.89 |||| 3.3e10 93.98 |||| 2.6e9 83.66 |||| 2.5e10 52.05 |||| 3.6e9 38.35 |||| 2.4e10 33.46 |||| 1.6e10
+q-FFL 93.74 |||| 7.8e10 93.86 |||| 2.5e8 85.20 |||| 7.4e8 48.46 |||| 4.7e9 51.12 |||| 2.7e8 49.83 |||| 1.1e7
+FedFDP 94.13 |||| 2.3e10 94.56 |||| 3.4e9 86.77 |||| 6.0e8 54.21 |||| 2.6e9 54.52 |||| 2.2e8 56.49 |||| 5.3e8
+[…]
+$\epsilon$ 1.0 2.0 3.0 4.0
+Accuracy Fairness Accuracy Fairness Accuracy Fairness Accuracy Fairness
+FedAvg 61.68 1.1e6 82.69 2.3e7 84.07 3.8e8 86.83 8.0e8
+SCAFFLOD 62.12 1.2e6 82.36 2.3e7 83.39 2.6e9 85.41 5.6e8
+FedProx 61.25 1.1e6 81.62 2.3e7 85.32 3.8e8 85.68 2.4e9
+FedDyn 62.38 1.5e6 82.43 3.1e7 83.29 2.9e8 84.51 5.2e9
+ALI-DPFL 61.78 1.1e6 83.01 2.3e7 84.12 3.9e8 85.58 4.5e9
+q-FFL 58.99 3.6e6 80.24 9.6e7 82.92 5.5e8 85.01 3.8e9
+FedFDP 63.36 1.0e6 83.15 2.1e7 85.59 9.5e7 86.97 3.8e8
+This section outlines our experimental framework, beginning with the presentation of baselines and default parameter configurations.
+We then detail the experimental assessment of FedFair and FedFDP across dimensions of performance, fairness, heterogeneity, and scalability.
+Ablation studies focusing on the hyperparameters $\lambda$, $C$ and $\epsilon$ of FedFDP are also included.
+All tasks will be repeated five times using different random seeds, and the mean and standard deviation will be reported.
+8.1. Default Settings
+We conducted experiments on three computer vision (CV) datasets: MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), and CIFAR10 (Krizhevsky, 2009).
+And we comparised our algorithms to six baselines:
+  1. (1)
+    FedAvg(McMahan et al., 2017): FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
+  2. (2)
+    SCAFFLOD(Karimireddy et al., 2020): SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
+  3. (3)
+    FedProx(Li et al., 2020b): FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
+  4. (4)
+    FedDyn(Acar et al., 2021): FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
+  5. (5)
+    ALI-DPFL(Ling et al., 2023): ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
+  6. (6)
+    q-FFL(Li et al., 2020c): q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
+For all datasets, we adopted a widely used heterogeneous settings (Lin et al., 2020; Li et al., 2021a) to 10 clients which controlled by a Dirichlet distribution denoted as $Dir(\beta)$, and the default value of $\beta=0.1$ (Lin et al., 2020; Wang et al., 2020).
+We use a 4-layer CNN architecture (McMahan et al., 2017) which consists of two convolutional layers and two fully connected layers as model architecture.
+In hyper-parameters, for FedFair and its baseline algorithms, we set learning rate $\eta=0.1$.
+For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$.
+In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
+8.2. Results
+In this subsection, we systematically evaluate the performance and fairness of the FedFair and FedFDP under varying heterogeneity levels, expanding from $Dir(0.1)$ to both $Dir(0.5)$ and $Dir(1)$.
+We also assess their scalability by scaling the number of clients from the standard 10 to encompass 20 and 50 clients.
+Furthermore, we examine the influence of three pivotal hyperparameters within the FedFDP algorithm, namely, the adjustments to $\lambda$, $C$ and $\epsilon$.
+8.2.1. FedFair
 Performance and Fairness.
-The experimental results of the FedFair algorithm and baseline methods are shown in Table 1.
-FedFair notably enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.5%.
-And Table 2 shows the experimental results of the FedFDP algorithm and baseline methods.
-Under the same privacy budget, FedFDP demonstrates improvements in model accuracy of 1.2%, 1.8%, and 0.4% compared to the best baseline on the three datasets, while also enhancing fairness by 30.3%, 6.7%, and 27.8% respectively compared to the best baseline.
-Figure 4: Impact of different $\lambda$ in FedFDF algorithm
+Table 2 presents the experimental results of the FedFair algorithm compared to baseline methods.
+FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
+Heterogeneity.
+Table 4 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity.
+Except for the $Dir(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $Dir(0.1)$ and $Dir(1.0)$, respectively.
+Scalability. Table 4 indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
+8.2.2. FedFDP
+Performance and Fairness.
+Table 3 displays the experimental results of the FedFDP algorithm compared to baseline methods.
+Under a consistent privacy budget of $\epsilon=3.52$, when accuracy matches the best baseline, our method significantly enhances fairness, with improvements of 30.3%, 6.7%, and 27.8% observed across three separate datasets.
+Heterogeneity.
+Table 5 shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity.
+Except for the $Dir(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $Dir(0.1)$ and $Dir(1.0)$, respectively.
+Scalability.
+Table 5 demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline.
+Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
+Figure 3. Impact of different $\lambda$ in FedFDP algorithm
+[…]
+Figure 4. Impact of different $C$ in FedFDP algorithm
+[…]
 Impact of Different $\lambda$.
-We investigated the correlation between $\lambda$ and fairness $\Psi$ in FedFDF algorithm.
-As depicted in Figure 4, our findings align with the theoretical analysis in Section 6.1: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase.
-Figure 5: Impact of different $C$ in FedFDF algorithm
+We investigated the correlation between $\lambda$ and fairness $\Psi$ in the FedFDP algorithm.
+As depicted in Figure 3, our findings align with the theoretical analysis in Section 6.1: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
 Impact of Different $C$.
 In Section 5.1, we emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously.
-As illustrated in Figure 5, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values of [0.01, 0.1, 1, 5, 10].
-Although $\Psi$ reaches its minimum when $C=1$, at this point, due to excessive noise addition, the model does not converge.
+As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10].
+Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition.
 As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$.
-When $C=0.1$, both accuracy and $\Psi$ achieve relatively good values.
-9 Conclusion
-In this paper, firstly, we propose a FedFair algorithm effectively addresses fairness concerns by optimizing a novel local loss function.
-Building upon FedFair, we propose the FedFDP, incorporates a fairness-aware gradient clipping strategy and an adaptive clipping method for additional loss values to achieve both fairness and differential privacy protection.
-And then, through convergence analysis, we determine the existence of an optimal fairness parameter, $\lambda$, striking a balance between model performance and fairness.
-We conduct a comprehensive privacy analysis of our approach using RDP and validate our algorithms through extensive experiments.
-The results indicate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness.
-We believe our work contributes a valuable FL framework for addressing fairness and privacy challenges.
+When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
+Different $\epsilon$.
+We conducted experiments with $\epsilon=[1,2,3,4]$ on the FashionMNIST dataset, as detailed in Table 6.
+For $\epsilon=[1,2,3]$, we allowed larger values of $\epsilon$ to support more iterations.
+At $\epsilon=4$, we supported a smaller noise multiplier $\sigma=1.65$ with a similar number of iterations as at $\epsilon=3$.
+The experimental results demonstrate that FedFDP consistently achieves performance on par with the best baseline in terms of accuracy across various privacy budgets.
+In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
+9. Conclusion
+In this paper, a FedFair algorithm is initially proposed to effectively address fairness concerns by optimizing a novel local loss function.
+The FedFDP, building upon FedFair, is introduced to incorporate a fairness-aware gradient clipping strategy and an adaptive clipping method for additional loss values, thereby achieving both fairness and differential privacy protection.
+Then, we find an optimal fairness parameter $\lambda^{*}$ through convergence analysis and numerical analysis methods, striking a balance between model performance and fairness.
+Subsequently, a comprehensive privacy analysis of the approach is conducted using RDP.
+Through extensive experiments, the results indicate that FedFair and FedFDP significantly outperform state-of-the-art solutions in terms of model performance and fairness.
+It is believed that our work contributes a valuable FL framework for addressing fairness and privacy challenges.
 Appendix A Appendix
-In the beginning, Table 3 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
-Table 3: Summary of main notations
-$F(\mathbf{w}$)
-Global loss function
-$F_{i}(\mathbf{w}$)
-Local loss function for client $i$
-$T$
-Communication rounds
-$t$
-Communication round index
-$D_{i}$
-Dataset of client $i$
-$D$
-The union of $D_{i}$
-$\mathbf{w}_{t}^{i}$
-Local model parameter at node $i$ at iteration $t$
-$\mathbf{w}^{*}$
-Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
-$\eta$
-Gradient descent step size
-$\lambda$
-Fairness parameter
-$N$
-Total number of client
-$q$
-batch sample ratio
-$|\mathcal{B}_{i}|$
-Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
-$d$
-The dimension of model
-$p_{i}$
-Weight of client $i$, equals to $|D_{i}|/|D|$
-$\xi_{j}$
-Single data sample in $D_{i}$, indexed by $j$
-$\sigma$
-Noise multiplier for gradient
-$\sigma_{l}$
-Noise multiplier for loss
-$C$
-Original clipping bound for gradient
-$C_{l}$
-Original Clipping bound for loss
-|||||$\cdot$||| | ⋅ | |
-$L_{2}$-norm
-A.1 Proof of Theorem 1
+A.1. Proof of Theorem 1
 For the purpose of validation, we introduce an additional variable $\mathbf{v}_{t}^{i}$ to represent the immediate result of a single-step DPSGD update from $\mathbf{w}_{t}^{i}$.
@@
 Consequently, the fair-clipping DPSGD in client $i$ at iteration $t$ transitions from Equation (14) to:
-$\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (19)
+(28) $\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$
 where:
-$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta^{j}_{loss},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
-In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by Stich [2018].
+$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
+In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by (Stich, 2018).
 Therefore,
 $\mathbf{v}_{t+1}$ $=\mathbf{w}_{t}$
-$\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (20)
-A.1.1 Key Lemma
+(29) $\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
+A.1.1. Key Lemma
 Lemma 0.
@@
 $\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2}\Delta_{t}\leq\frac{L}{2t}v\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\Delta_{1}\right)$
-A.1.2 Proof of Lemma 1
+A.1.2. Proof of Lemma 1
 By the Equation (A.1), we get
 $\left\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\right\|^{2}$
-$*\displaystyle=\|\mathbf{w}_{t}-\mathbf{w}^{*}$
+$=\|\mathbf{w}_{t}-\mathbf{w}^{*}$
 $-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\|^{2}$
@@
 $\quad+\underbrace{\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\cdot\mathcal{N}(0,\mathbf{I})\|^{2}}_{\mathcal{B}_{2}}$
-Since $\mathbb{E}\mathcal{B}_{0}=0$, we focus on $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$:
-$\mathbb{E}\mathcal{B}_{2}\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
+Since $\mathbb{E}[\mathcal{B}_{0}]=0$, we focus on $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$:
+$\mathbb{E}[\mathcal{B}_{2}]\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
 where $\frac{1}{\hat{B}^{2}}\triangleq\max_{i}\frac{1}{|\mathcal{B}_{i}|}$.
@@
 taking exception and according to assumption 3:
-$\mathbb{E}\mathcal{B}_{1}\leq\eta^{2}\mathbb{E}\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\leq\eta^{2}C_{t}^{2}G^{2},$
+$\mathbb{E}[\mathcal{B}_{1}]\leq\eta^{2}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\right]\leq\eta^{2}C_{t}^{2}G^{2},$
 where $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i}$, $C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
 Now, we obtain the bound for the expectation of $\mathcal{A}_{2}$:
-$\mathbb{E}\mathcal{A}_{2}$ $\leq\eta^{2}\left(\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}+G^{2}C_{t}^{2}\right)$
+$\mathbb{E}[\mathcal{A}_{2}]$ $\leq\eta^{2}\left(\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}+G^{2}C_{t}^{2}\right)$
 The process of $\mathcal{A}_{1}$ show as below:
@@
 $\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{2}}$
-It’s obvious that $\mathbb{E}\mathcal{C}_{0}=0$.
+It’s obvious that $\mathbb{E}[\mathcal{C}_{0}]=0$.
 By Cauchy-Schwarz inequality and AM-GM inequality, we have
@@
 So we get
-$\mathbb{E}\mathcal{C}_{1}\leq\sum_{i=1}^{N}p_{i}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}$
+$\mathbb{E}[\mathcal{C}_{1}]\leq\sum_{i=1}^{N}p_{i}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}$
 According to Assumption 2, we know that
@@
 where L-smooth inference as show:
-$\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$ (21)
+(30) $\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$
 Thus, we get
@@
 To sum up,
-$\mathbb{E}\mathcal{A}_{1}$ $=\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}+2\eta^{2}C_{t}L\Gamma$
-$\quad+C_{t}\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
-$=(1+C_{t})\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
+$\mathbb{E}[\mathcal{A}_{1}]$ $=\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}+2\eta^{2}C_{t}L\Gamma$
+$\quad+C_{t}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
+$=(1+C_{t})\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
 $\quad+\eta^{2}\left(G^{2}C_{t}^{2}+2L\Gamma C_{t}\right),$
 and
-$\quad\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}$
-$=\mathbb{E}\sum_{i=1}^{N}p_{i}\|(\mathbf{w}_{t}-\mathbf{v}_{t+1}^{i})-(\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i})\|^{2}$
-$\leq\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i}\|^{2}$
+$\quad\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]$
+$=\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|(\mathbf{w}_{t}-\mathbf{v}_{t+1}^{i})-(\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i})\|^{2}\right]$
+$\leq\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i}\|^{2}\right]$
 $\leq\mathbb{E}\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}$
-$=\mathbb{E}\mathcal{A}_{2}=\eta^{2}C_{t}^{2}G^{2}+\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
+$=\mathbb{E}\left[\mathcal{A}_{2}\right]=\eta^{2}C_{t}^{2}G^{2}+\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
 So we get,
-$\mathbb{E}\mathcal{A}_{1}$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}$
+$\mathbb{E}\left[\mathcal{A}_{1}\right]$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}$
 $\quad+\eta^{2}\left(2G^{2}C_{t}^{2}+2L\Gamma C_{t}+2\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}\right)$
@@
 $A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
-Herterogeneity Scalability
-Datasets MNIST FashionMNIST CIFAR10
-Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
-FedAvg 98.67 |||| 7.3e-3 98.37 |||| 1.2e-3 89.41 |||| 3.6e-2 62.03 |||| 1.1e0 60.89 |||| 1.5e0 59.78 |||| 1.5e0
-SCAFFLOD 98.45 |||| 5.0e-3 98.51 |||| 2.1e-3 86.39 |||| 3.4e-2 62.21 |||| 9.9e0 59.48 |||| 8.7e0 58.64 |||| 2.3e0
-FedProx 98.70 |||| 7.5-3 98.35 |||| 1.2e-3 88.33 |||| 3.5e-2 62.13 |||| 1.2e0 60.33 |||| 2.2e0 61.31 |||| 1.6e0
-FedDyn 98.40 |||| 8.0e-3 98.36 |||| 1.8e-3 88.26 |||| 3.5e-2 62.33 |||| 9.4e-1 61.35 |||| 2.6e0 62.52 |||| 1.5e0
-q-FFL 98.72 |||| 4.7e-3 98.61 |||| 1.1e-3 89.03 |||| 3.6e-2 63.33 |||| 9.4e-1 60.63 |||| 1.3e0 59.65 |||| 1.5e0
-FedFair 98.75 |||| 3.9e-3 98.72 |||| 1.2e-3 89.60 |||| 1.7e-2 62.38 |||| 8.5e-1 61.58 |||| 1.1e0 62.90 |||| 1.5e0
-Table 4: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
-Herterogeneity Scalability
-Datasets MNIST FashionMNIST CIFAR10
-Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
-FedAvg 93.40 |||| 1.1e11 94.67 |||| 4.1e9 75.60 |||| 3.7e9 52.61 |||| 6.1e9 51.58 |||| 2.0e9 55.90 |||| 1.5e9
-SCAFFLOD 93.95 |||| 7.0e10 94.06 |||| 1.9e9 72.56 |||| 3.6e9 53.85 |||| 6.2e9 54.34 |||| 2.9e8 53.98 |||| 2.2e9
-FedProx 90.50 |||| 5.3e10 94.23 |||| 3.1e9 71.38 |||| 5.6e9 51.34 |||| 4.4e9 53.67 |||| 4.5e8 54.14 |||| 2.2e9
-FedDyn 91.42 |||| 5.6e10 93.08 |||| 3.8e9 73.21 |||| 9.8e10 52.14 |||| 5.5e9 50.21 |||| 1.2e10 51.86 |||| 2.0e10
-ALI-DPFL 90.89 |||| 3.3e10 93.98 |||| 2.6e9 77.66 |||| 2.5e10 52.05 |||| 3.6e9 38.35 |||| 2.4e10 33.46 |||| 1.6e10
-q-FFL 93.74 |||| 7.8e10 93.86 |||| 2.5e8 82.20 |||| 7.4e8 48.46 |||| 4.7e9 51.12 |||| 2.7e8 49.83 |||| 1.1e7
-FedFDP 95.13 |||| 2.3e10 94.56 |||| 3.4e9 83.77 |||| 6.0e8 54.21 |||| 2.6e9 54.52 |||| 2.2e8 56.49 |||| 5.3e8
-Table 5: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
-A.2 Proof of Theorem 2
-As Algorithm 2 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$.
-$\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of local client, so they are both to perform differential privacy preservation, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively.
-We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
-A.2.1 Privacy Loss of $\mathbf{w}_{t+1}^{i}$
-Theorem 3.
-After $T$ rounds local updates, the RDP of the $\mathbf{w}_{T}^{i}$ in $i$-th client satisfies:
-$R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$ (22)
-where $\sigma$ is noise multiplier of the $w_{t+1}^{i}$, and $\alpha>1$ is the order.
-$proof.$ We will prove this in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
+A.2. Proof of Theorem 1
+$proof.$ We will prove Theorem 1 in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
 Definition 0.
-(RDP privacy budget of SGMMironov et al.
-[2019]).
+(RDP privacy budget of SGM(Mironov et al., 2019)).
 Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$.
 If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
-$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$ (23)
+(31) $R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$
 where
-$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$ (24)
+(32) $\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$
 with $\vartheta_{0}\triangleq\mathcal{N}\left(0,\sigma^{2}\right),\vartheta_{1}\triangleq\mathcal{N}\left(1,\sigma^{2}\right)\mbox{ and }\vartheta\triangleq(1-q)\vartheta_{0}+q\vartheta_{1}$
@@
 Thus, $SG_{q,\sigma}\text{ satisfies }\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
-Finally, the existing work Mironov et al. [2019] describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
-$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$ (25)
+Finally, the existing work (Mironov et al., 2019) describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
+(33) $A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$
 Definition 0.
-(Composition of RDPMironov [2017]).
+(Composition of RDP(Mironov, 2017)).
 For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
 From Definition 2 and Definition 3, the Theorem 1 is obtained.
-A.2.2 Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$
-Theorem 4.
-After $T$ rounds local updates, the RDP of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ in $i$-th client satisfies:
-$R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$ (26)
-where $\sigma_{l}$ is noise multiplier of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$, and $\alpha>1$ is the order.
-The proof is similar to that of the $w_{t+1}^{i}$, so we omit it.
-A.2.3 Proof Loss of FedFDP
-Since both $w_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 3, and then use Lemma 2 to convert it to $(\epsilon,\delta)$-DP.
-Therefore, the Theorem 2 is hold.
-Lemma 2.
-(Conversion from RDP to DP Balle et al.
-[2020]).
-if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
-A.3 Experiment Settings
-A.3.1 Parameters Settings
-For all datasets, we adopted a widely used heterogeneous settingsLin et al.
-[2020]Li et al.
-[2021a] to 10 clients which controlled by a Dirichlet distribution denoted as $Dir(\beta)$, where a smaller $\beta$ indicates a higher degree of heterogeneity.
-We set the default value of $\beta=0.1$ Lin et al.
-[2020]Wang et al.
-[2020].
-In all datasets, we use a 4-layer CNN architecture McMahan et al.
-[2017] which consists of two convolutional layers and two fully connected layers.
-For FedFair and its baseline algorithms, we set learning rate $\eta=0.1$.
-For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$.
-In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
-In all experiments, we use 25% of the data will form the test dataset, and the remaining 75% of the data will be used for training.
-All tasks will be repeated five times using different random seeds, and the mean and standard deviation will be reported.
-A.3.2 Baselines introduction
-The baseline methods included in the comparison are:
-  • FedAvgMcMahan et al. [2017]: FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFDP into the FedAvg algorithm.
-  • SCAFFLODKarimireddy et al.
-[2020]: SCAFFOLD uses control variates (variance reduction) to correct for the ’client-drift’ in its local updates.
-It requires significantly fewer communication rounds and is not affected by data heterogeneity or client sampling.
-  • FedProxLi et al. [2020a]: FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
-  • FedDynAcar et al.
-[2021]: FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
-  • ALI-DPFLLing et al.
-[2023]: ALI-DPFL dynamically obtains the optimal number of epochs for each round of communication by minimizing the convergence upper bound, thereby addressing the imbalance between local privacy budgets and communication resources.
-  • q-FFLLi et al.
-[2020b]: q-FFL amplifies the loss values of clients, equivalent to giving larger update weights to poorly performing clients, in order to cater to the underperforming clients, uniforming model performance across multiple devices.
-A.3.3 Herterogeneity and Scalability
-As shown in Table 4 and 5, we conducted additional experiments on the MNIST, FashionMNIST, and CIFAR10 datasets to compare the performance of FedFDP with that of six other algorithms in terms of Heterogeneity and Scalability.
-The Heterogeneity experiments were mainly conducted in three heterogeneous scenarios, extending from the default $Dir(0.1)$ to include $Dir(0.5)$ and $Dir(1)$.
-Scalability was expanded from the default 10 clients to include 20 and 50 clients.
-Heterogeneities.
-Tables 4 and 5 show that when confronted with diverse data distributions exhibiting varying levels of heterogeneity, both FedFair and FedFDP outperform baseline methods in accuracy while maintaining significant fairness.
-With the exception of the $Dir(0.5)$ scenario, where they fall slightly short of q-FFL, they respectively demonstrated fairness improvements of 17.0% and 50% for FedFair and 30.3% and 18.9% for FedFDP compared to the best baselines in $Dir(0.1)$ and $Dir(1.0)$.
-Scalability.
-Table 4 demonstrates that, with accuracy mostly superior to baselines, FedFair exhibited fairness improvements of 9.5%, 15.4%, and 0.01% across three client counts.
-Table 5 indicates that as the number of clients increases, FedFDP outperformed the best baselines by 0.36%, 0.18%, and 0.59% in accuracy, except in the 50 clients scenario, where it demonstrated fairness improvements of 27.8% and 24.1%.

### v3

@@
 FedFDP: Fairness-Aware Federated Learning with Differential Privacy
-Abstract.
+Abstract
 Federated learning (FL) is a new machine learning paradigm to overcome the challenge of data silos and has garnered significant attention.
-However, through our observations, a globally effective trained model may performance disparities in different clients.
-This implies that the jointly trained models by clients may lead to unfair outcomes.
-On the other hand, relevant studies indicate that the transmission of gradients or models in federated learning can also give rise to privacy leakage issues, such as membership inference attacks.
-To address the first issue mentioned above, we propose a fairness-aware federated learning algorithm, termed FedFair.
-Building upon FedFair, we introduce privacy protection to form the FedFDP algorithm to address the second issue mentioned above.
-In FedFDP, we devise a fairness-aware clipping strategy to achieve differential privacy while adjusting fairness.
-Additionally, for the extra uploaded loss values, we present an adaptive clipping approach to maximize utility.
-Furthermore, we theoretically prove that our algorithm converges and ensures differential privacy.
-Lastly, extensive experimental results demonstrate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness.
-Code and data is accessible at https://anonymous.4open.science/r/FedFDP-5607.
-Federated Learning, Differential Privacy, Fairness
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
-1. Introduction
-Federated Learning (FL) (McMahan et al., 2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without revealing their individual datasets.
-After training their local models, each client transmits only the model parameters, not the raw training data, to a central server.
-The server aggregates these parameters to update the global model, which is then sent back to clients for further training.
-Thanks to its ability to overcome data silos by jointly improving model performance without the need to upload raw data, FL has gained significant attention in recent years (Zhu et al., 2021)(Lee et al., 2022).
-In particular, many studies have made great breakthroughs in model performance (Wang et al., 2020; Li et al., 2020b, 2021b) and communication costs (Wang et al., 2019; Ling et al., 2023) in FL.
-Figure 1. The test accuracies of 50 clients utilizing a globally trained model in FL.
-[…]
-To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary (Li et al., 2021c).
-Current machine learning algorithms that protect fairness, such as  (Dwork et al., 2012; Cotter et al., 2019), have not adequately addressed the issue of performance disparities among clients in federated learning.
-As shown in Figure 1, the red dots represent the test accuracies of the global model on different clients after completing federated learning with the FedAvg algorithm (McMahan et al., 2017).
-The global model trained collaboratively by clients exhibits significant unbalanced performance across different datasets held by the various clients.
-An increasing number of scholars are beginning to pay attention to this performance disparities phenomenon and are starting to work on it  (Li et al., 2020c, 2021b; Zhao and Joshi, 2022).
-On the other hand, recent research indicates that trained parameters can leak data privacy, such as the recovery of original data (Zhu et al., 2019) or the membership inference attack (Song et al., 2017).
-Moreover, Melis et al.
-(Melis et al., 2022) demonstrated that clients’ training data can be leaked through shared models.
-In order to enhance data privacy in FL, numerous scholars have begun integrating techniques such as differential privacy (DP) (Dwork et al., 2014), homomorphic encryption (Paillier, 1999), and secure multi-party computation (Yao, 1982) into the federated learning framework.
-Among these, DP stands out as the premier approach to privacy preservation in federated learning, attributed to its modest computational overhead and robust mathematical underpinnings.
-This is particularly relevant in defending membership inference attacks, as several studies (Abadi et al., 2016; Carlini et al., 2019; Feldman, 2020) have shown that incorporating appropriate DP during the training phase can effectively prevent inadvertent disclosure of sensitive training information.
-Although numerous articles have addressed fairness and differential privacy individually within the context of federated learning, there is a scarcity of research that integrates these two aspects.
-Achieving a balance between model performance, fairness, and differential privacy protection within federated learning frameworks poses a significant challenge.
-This paper aims to investigate fairness and differential privacy protection mechanisms in federated learning.
-By analyzing the relationship between them, we propose a fairness-aware federated learning with differential privacy, effectively guaranteeing model performance when achieving fairness and privacy protection.
-Our main contributions are listed as follows:
-  1. (1)
-    Firstly, a fairness-aware federated algorithm is proposed, denoted as FedFair.
-It incorporates a novel loss function, designed to optimize the objectives of fairness and model performance simultaneously (Section 4).
-FedFair algorithm achieves the excellent fairness, as shown by the green dots in Figure1.
-  2. (2)
-    Building upon FedFair, an algorithm referred to as FedFDP is proposed to achieve differential privacy protection (Section 5).
-Particularly, we device a fairness-aware gradient clipping strategy to ensures both differential privacy and allows for the adjustment of fairness in federated learning (Section 5.1).
-Additionally, we propose an adaptive clipping approach method for the additional loss values uploaded by each client to achieve optimal utility (Section 5.2).
-  3. (3)
-    Furthermore, it is demonstrated that an optimal fairness parameter $\lambda$ is found, which, when determined, enables the fastest convergence of FedFDP in balancing model performance and fairness through the convergence analysis (Section 6).
-  4. (4)
-    And then, a privacy loss analysis is conducted on algorithm FedFDP using the notion of Rényi differential privacy (RDP) to verify its satisfaction of differential privacy requirements (Section 7).
-  5. (5)
-    Lastly, through a comprehensive empirical evaluation on three public datasets, it is confirmed that FedFair and FedFDP outperform the state-of-the-art solutions in terms of model performance and fairness (Section 8).
-2. Related Work
+However, federated learning faces challenges in fairness and data privacy.
+To address both of the above challenges simultaneously, we first propose a fairness-aware federated learning algorithm, termed FedFair.
+Then based on FedFair, we introduce differential privacy protection to form the FedFDP algorithm to address the trade-offs among fairness, privacy protection, and model performance.
+In FedFDP, we designed an fairness-aware gradient clipping technique to identify the relationship between fairness and differential privacy.
+Through convergence analysis, we determined the optimal fairness adjustment parameters to simultaneously achieve the best model performance and fairness.
+Additionally, for the extra uploaded loss values, we present an adaptive clipping method to minimize privacy budget consumption.
+Extensive experimental results demonstrate that FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness.
+Codes and datasets will be made public after acceptance.
+1 Introduction
+Federated Learning (FL) (McMahan et al.
+2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without exposing their respective datasets.
+After training their local models, each client only transmits model parameters to the central server, rather than the original training data.
+The server aggregates these parameters to update the global model, which is then sent back to the clients for further training.
+Due to its ability to improve model performance through collaboration without the need to upload original data, overcoming data silos, FL has received significant attention in recent years (Zhu, Hong, and Zhou 2021; Lee et al.
+2022; Chen et al.
+2024).
+Particularly, many studies have achieved significant breakthroughs in terms of model performance (Wang et al.
+2020; Li et al.
+2020a, 2021b) and communication costs (Wang et al.
+2019; Ling et al.
+2024) in FL.
+Figure 1: Client drift caused by Non-IID.
+To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary (Li et al.
+2021a).
+Current fairness-aware machine learning algorithms, such as (Dwork et al.
+2012; Cotter et al.
+2019), have not fully addressed the issue of client performance disparities in federated learning.
+As shown in Figure 1, due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (yellow squares) lead to significant performance imbalances of the collaboratively trained global model (red square) across different datasets held by various clients.
+An increasing number of scholars are beginning to focus on this phenomenon of unfair performance disparity (Li et al.
+2020b, 2021b; Zhao and Joshi 2022).
+On the other hand, recent studies have pointed out that training parameters may leak data privacy, such as recovering original data (Zhu, Liu, and Han 2019) or member inference attacks (Song, Ristenpart, and Shmatikov 2017).
+Even if only models or gradients are uploaded instead of original data, data leakage can still occur (Melis et al.
+2022).
+To enhance data privacy in FL, differential privacy (DP) (Dwork, Roth et al.
+2014) has become the preferred method for protecting privacy in federated learning due to its moderate computational overhead and solid mathematical foundation.
+Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information (Abadi et al.
+2016; Carlini et al.
+2019; Feldman 2020).
+Although some works have explored fairness and differential privacy in federated learning separately, no prior research has considered them in a unified framework.
+For the first time, we propose a fairness-aware federated learning with differential privacy.
+However, balancing differential privacy protection, fairness, and model performance presents a significant challenge.
+Specifically, there are two main challenges: First, the clipping and noise addition steps in differential privacy can impact fairness and model utility in federated learning.
+Second, fair federated learning requires uploading loss information to the central server, which incurs additional privacy budget consumption.
+In this paper, to address the first challenge, we analyze the relationship between the fairness loss function and differential privacy in federated learning.
+We design an fairness-aware gradient clipping strategy to balance fairness and differential privacy.
+This gradient clipping strategy not only meets the requirements for fairness but also satisfies the sensitivity control needed for differential privacy.
+Furthermore, through convergence analysis, we demonstrate that an optimal adjustment parameter can be found to achieve the best model performance and fairness.
+To address the second challenge, we design an adaptive clipping method for loss uploading, significantly reducing the privacy budget consumption in this part.
+Our main contributions are as follows:
+  • First, we propose a fairness-aware federated algorithm named FedFair. It combines a novel loss function aimed at optimizing the objectives of fairness and model performance simultaneously.
+  • Based on FedFair, we propose an algorithm called FedFDP to further equip the system with differential privacy.
+In particular, we designed a fairness-aware gradient clipping strategy that ensures differential privacy and allows for adjustment of fairness in federated learning.
+In addition, we propose an adaptive clipping method for the additional loss values uploaded by each client to achieve optimal utility.
+  • Further, we conducted a convergence analysis of FedFDP and found the optimal fairness parameter $\lambda$ by minimizing the upper bound of convergence.
+We analyzed the privacy loss of the FedFDP algorithm using the concept of Rényi Differential Privacy (RDP) to verify its compliance with differential privacy requirements.
+  • Finally, through comprehensive empirical evaluations on three public datasets, we confirmed that FedFair and FedFDP are superior to existing solutions in terms of model performance and fairness.
+What’s more, we investigated the impact of clipping bound and noise multiplier in differential privacy on fairness.
+2 Related Work
 Differentially private federated learning (DPFL) is widely used for privacy protection in federated learning.
-Many studies have focused on the trade-off between model performance, communication efficiency, and privacy budget in DPFL (Ullah et al., 2023; Zhou et al., 2023; Fu et al., 2022; Ling et al., 2023; Kerkouche et al., 2021; Fu et al., 2024).
+Many studies have focused on the trade-off between model performance, communication efficiency, and privacy budget in DPFL (Fu, Chen, and Han 2022; Zhou et al.
+2023; Ling et al.
+2024; Fu et al.
+2024).
 In addition that, various fairness concepts in machine learning have been proposed, each emphasizing different aspects and addressing the needs of different stakeholders.
-For example, concepts like Good-Intent Fairness (Mohri et al., 2019) and Selection Fairness (Zhou et al., 2021) have been introduced.
-Further discussions on these concepts can be found in additional literature (Cong et al., 2020; Lyu et al., 2020; Mehrabi et al., 2021; Shi et al., 2021).
+For example, concepts like Good-Intent Fairness (Mohri, Sivek, and Suresh 2019) and Selection Fairness (Zhou, Fang, and Hui 2021) have been introduced.
+Further discussions on these concepts can be found in additional literature (Cong et al.
+2020; Lyu et al.
+2020) .
 At present, the primary criterion for fairness in FL is group fairness, which aims to mitigate biases associated with sensitive attributes like gender and race.
-Relevant research has shown that in federated learning with differential privacy, adjusting parameters such as the clipping threshold and noise multiplier can help enhance the algorithm’s group fairness (Gu et al., 2022).
-There is a growing body of literature that addresses the concept of group fairness specifically within the framework of DPFL (Gálvez et al., 2021; Padala et al., 2021; Gu et al., 2022, 2020; Pentyala et al., 2022).
+Relevant research has shown that in federated learning with differential privacy, adjusting parameters such as the clipping threshold and noise multiplier can help enhance the algorithm’s group fairness (Gu et al.
+2022).
+There is a growing body of literature that addresses the concept of group fairness specifically within the framework of DPFL (Gálvez et al.
+2021; Padala, Damle, and Gujar 2021; Gu et al.
+2022, 2020; Pentyala et al.
+2022).
 The more predominant fairness criterion in FL is balanced performance fairness, as introduced by Li et al.
-(Li et al., 2020c).
+(Li et al.
+2020b).
 This approach ensures fairness by assessing the uniformity of model performance across participating clients.
@@
 Building on this, Zhao et al.
-(Zhao and Joshi, 2022) proposed an alternative to the loss amplification mechanism, advocating for a direct weight reallocation strategy.
+(Zhao and Joshi 2022) proposed an alternative to the loss amplification mechanism, advocating for a direct weight reallocation strategy.
 Concurrently, Li et al.
-(Li et al., 2021b) concentrated on enhancing the fairness of local models within personalized FL.
+(Li et al.
+2021b) concentrated on enhancing the fairness of local models within personalized FL.
 Our study pioneers the investigation of balanced performance fairness within the context of DPFL.
-In comparison to definitions such as group fairness, balanced performance fairness is more meaningful for data providers such as enterprises and institutions, moreover, it can effectively motivate clients participate in the federated system.
-However, achieving balanced performance fairness typically requires adjusting the direction of model updates.
-It is a challenge to refine the model for balanced performance fairness while integrating DP techniques, such as clipping and noise addition, into the model’s update process.
-To the best of our knowledge, we are the first to consider the work of balanced performance fairness in DPFL.
-3. Preliminaries
-In the beginning, Table 1 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
-Table 1. Summary of main notations
-$F(\mathbf{w}$) Global loss function
-$F_{i}(\mathbf{w}$) Local loss function for client $i$
-$T$ Communication rounds
-$t$ Communication round index
-$D_{i}$ Dataset of client $i$
-$D$ The union of $D_{i}$
-$\mathbf{w}_{t}^{i}$ Local model parameter at node $i$ at round $t$
-$\mathbf{w}^{*}$ Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
-$\eta$ Gradient descent step size
-$\lambda$ Fairness parameter
-$N$ Total number of clients
-$q$ batch sample ratio
-$|\mathcal{B}_{i}|$ Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
-$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
-$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
-$\sigma$ Noise multiplier for gradient
-$\sigma_{l}$ Noise multiplier for loss
-$C$ Original clipping bound for gradient
-$C_{l}$ Original Clipping bound for loss
-|||||$\cdot$||| | ⋅ | | $L_{2}$-norm
-3.1. Federated Learning and Fairness
+3 Preliminaries
+3.1 Federated Learning and Fairness
 We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model.
 The objective of FL is to minimize $F(\mathbf{\cdot})$, in other words, to seek:
-(1) $\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$
+$\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (1)
 Here, $F_{i}(\cdot)$ represents the local loss function of client $i$.
@@
 We have defined balanced performance fairness in FL as follows.
-Specifically, this type of fairness is defined based on the performance variance of the model across clients: the smaller the variance, the fairer the model.
-Definition 0.
-(Balanced performance fairness (Li et al., 2020c)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
-3.2. Differential Privacy
+Definition 1
+(Balanced performance fairness (Li et al.
+2020b)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
+In this study, we employ the loss of model denote the performance.
+Equation (2) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients.
+Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
+$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (2)
+3.2 Differential Privacy
 Differential privacy is a formal mathematical model that quantifies privacy leakage in data analysis algorithms.
 It stipulates that alterations to a single record within the training data should not induce significant statistical variations in the algorithm’s results.
-Definition 0.
-(Differential Privacy (Dwork et al., 2014)).
+Definition 2
+(Differential Privacy (Dwork, Roth et al.
+2014)).
 $(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
-(2) $\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$
+$\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$ (3)
 By adding random noise, we can achieve differential privacy for a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ according to Definition 2.
@@
 We use $l_{2}$ sensitivity as in the previous work.
-Definition 0.
-($\mathbf{l_{k}}$-Sensitivity(Dwork et al., 2006)) For a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, we define its $l_{k}$ norm sensitivity (denoted as $\Delta_{k}f$) over all neighboring datasets $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ differing in a single sample as
-(3) $\text{sup}_{D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}}||f(D_{i})-f(D_{i}^{\prime})||_{k}\leq\Delta_{k}f.$
-DPSGD. In gradient-based methods, DPSGD(Abadi et al., 2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
-  1. (1)
-    Firstly, computing the gradient of loss function:
-    (4) $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}).$
-  2. (2)
-    Secondly, clipping the gradient of per-sample to get the sensitivity:
-    (5) $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right),$
-    where $C$ represents the clipping bound.
-  3. (3)
-    Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier:
-    (6) $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$
+DPSGD. In gradient-based methods, DPSGD(Abadi et al. 2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
+  • Firstly, computing the gradient of loss function: $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
+  • Secondly, clipping the gradient of per-sample to get the sensitivity: $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right)$, where $C$ represents the clipping bound.
+  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier: $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right)$.
 The final step is to proceed with gradient descent as usual.
 When each client in FL utilizes DPSGD as the optimization algorithm to safeguard their private data, the framework transitions from FL to DPFL.
-Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, fairness hyperparameter $\lambda$, batch sample ratio $q$, local dataset $D_{i}$.
-Output: the final trained model $\mathbf{w}_{T}$
-1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
-2 SERVER
-3       for $t\in[0,1,\cdots,T-1]$  do
-4             for $i\in[1,2,\cdots,N]$ parallel  do
-5                   $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
-6
-7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
-8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t+1}^{i})$;
-9
-10      return $\mathbf{w}_{T}$;
-11
-12LOCAL
-13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
-14
-15      $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
-16       for batch $b\in\mathcal{B}_{i}$ do
-17             $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i},b)-F(\mathbf{w}_{t})$;
-18             $\eta_{i}=\eta\cdot\left(1+\lambda\cdot\Delta_{i}\right)$;
-19             $\mathbf{w}_{t}^{i}=\mathbf{w}_{t}^{i}-\eta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},b)$;
-20
-21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ , compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
-22       return $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})$;
-23
-Algorithm 1 FedFair
-4. Proposed Algorithm: FedFair
+4 FedFair: Equipping FL with Fairness
 We previously defined the performance objective in Equation (1) and balanced performance fairness in Definition 1.
 Here, we use these to develop a new objective function that informs our proposed fairness-aware federated learning algorithm.
-In this study, we employ the loss of model denote the performance.
-Equation (7) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients.
-Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
-(7) $\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$
 Upon incorporating Equation (2) into Equation (1), we introduce a comprehensive objective function that integrates considerations of fairness:
-(8) $\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$
+$\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (4)
 For client $i$ during round $t$, the refined objective is articulated as follows:
-(9) $\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$
+$\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$ (5)
 where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training.
-from round $t-1$, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning.
+from round $t-1$, as delineated in line 2 of Algorithm 2, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning.
 When $\lambda\rightarrow\infty$, the algorithm only focus on fairness.
 However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging.
-Fortunately, we theoretically found an optimal value for $\lambda$ in section 6.2.
+Fortunately, we theoretically found an optimal value for $\lambda$.
 In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$..
@@
 So, the gradient of Equation (5) as follows:
-(10) $\nabla H_{i}(\mathbf{w}_{t}^{i})=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$
+$\nabla H_{i}(\mathbf{w}_{t}^{i})=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (6)
 where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t}^{i})$.
 The stochastic gradient descent by using Equation (6) as show:
-(11) $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$
+$\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (7)
 where $\eta_{i}=\eta(1+\lambda\Delta_{i})$.
-Figure 2. FedFDP framework.
-[…]
+Figure 2: FedFDP framework.
 Therefore, our approach can be viewed as establishing an adaptive learning rate that allows the model update to achieve both performance and fairness objectives simultaneously.
-Based on these principles, we introduce the FedFair algorithm, as depicted in Algorithm 1.
-  • Server (lines 1-1): The server executes $T$ rounds of communication, each round sending the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
-  • Local (lines 1-1): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
-Furthermore, as the training process advances, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$.
-In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods (Zinkevich, 2003).
-5. Proposed Algorithm: FedFDP
-In this section, we propose the FedFDP by integrating differential privacy protection base on FedFair.
-The flow of FedFDP algorithm as shown in Figure 2.
-The FedFDP algorithm adds the following two additional processes to FedFair.
-Firstly, in Section 5.1, we present a fair-clipping strategy that not only achieves sensitivity but also allows adjustments to the level of fairness.
-Furthermore, in Section 5.2, we expand the use of differential privacy protection to the additional loss values that local clients need to upload, suggesting an adaptive clipping method to maximize utility.
-The overall FedFDP algorithm is described in Section 5.3.
-Input: loss function $F(\mathbf{w})$.
-Parameters: learning rate $\eta$, noise multiplier for gradient $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
-Output: the final trained model $\mathbf{w}_{T}$
+Based on these principles, we introduce the FedFair algorithm, as depicted in Appendix Algorithm 2.
+5 FedFDP: Further Equipping FedFair with Differential Privacy
+In this section, we present our algorithm, FedFDP, which further integrates FedFair with differential privacy.
+The flow of FedFDP algorithm is shown in Figure 2.
+It adds two additional processes to FedFair:
+  • A fair-clipping strategy is added to FedFair. This strategy not only achieves sensitivity but also allows adjustments to the level of fairness.
+  • An adaptive clipping method is employed when protecting the loss values that clients share with differential privacy, aiming to maximize utility.
+Next, we discuss the details of these two strategies.
 1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0}),C_{l}^{0}$;
@@
 27
-28      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})\leftarrow\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\cdot\mathcal{N}(0,\mathbf{I})\right)$;
+28      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\mathcal{N}(0,\mathbf{I})\right)$;
 29       return $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
@@
 Algorithm 1 FedFDP
-5.1. Fair-clipping Strategy for Gradient
-In the general DPSGD algorithm (Abadi et al., 2016; Fu et al., 2023; Bu et al., 2022; Yang et al., 2022), per-sample clipping is an essential process for obtaining sensitivity.
+5.1 Fair-clipping Strategy for Gradient
+In the general DPSGD algorithm (Abadi et al.
+2016; Fu et al.
+2023; Bu et al.
+2022; Yang et al.
+2022), per-sample clipping is an essential process for obtaining sensitivity.
 In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (5), into Equation (8):
-(12) $\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$
+$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$ (8)
 where $\xi_{j}\in\mathcal{B}_{t}^{i}$, $j=\{1,2,\cdots,|\mathcal{B}_{i}|\}$, and $|\mathcal{B}_{i}|$ is the batch size of client $i$.
 Next, we compute the gradient of Equation (8):
-$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\lambda\cdot\Delta_{i}^{j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$
-(13) $=\left(1+\lambda\cdot\Delta_{i}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$
+$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})=\left(1+\lambda\cdot\Delta_{i}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$ (9)
 where $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})$.
-To achieve differential privacy protection, we replace the gradient of loss function by Equation (5.1) based on Equations (4)-(6), and reorganize the gradient descent equation for client $i$:
-$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
-$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
-(14) $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$
+To ensure the algorithm is differentially private, we replace the gradient of loss function by Equation (9), and reformulate the computation of the gradient descent for client $i$ as follows:
+$\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (10)
 where:
-(15) $C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
-Hence, our strategy can be regarded as an fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model.
-When $\lambda=0$, Equation (15) corresponds to the traditional DP clipping method, and FedFDP no longer ensures fairness.
+$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$ (11)
+Hence, our strategy can be considered as a fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model.
+When $\lambda=0$, Equation (11) corresponds to the traditional DP clipping method without fairness.
 For large $C$, Equation (11) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness.
 Conversely, for small $C$, Equation (11) simplifies to the second term, resulting in the gradient being clipped to the bound $C$.
-Our conclusions can be verified from the experiments (Figure 4).
-5.2. Adaptive Clipping Method for Loss
+The impact of $C$ on the performance of FedFDP will be further discussed in the experiments.
+5.2 Adaptive Clipping Method for Loss
 Since the calculation of $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy.
-As shown in lines 2-2 of Algorithm 2, We need to clip $F_{i}(\mathbf{w}_{t}^{i})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy.
+As shown in lines 1-1 of Algorithm 1, We need to clip $F_{i}(\mathbf{w}_{t}^{i})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy.
 However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging.
@@
 The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1}^{i})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
-(16) $C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$
+$C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$ (12)
 where
-(17) $\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))).$
+$\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))).$ (13)
 To ensure the achievement of differential privacy during the clipping process, we use the noise-added loss as the clipping bound $C_{l}^{i,t}$.
 Each round’s clipping bound adheres to the post-processing nature of differential privacy when $t=\{1,\dots,T\}$.
-5.3. Overall Algorithm
+5.3 Overall Algorithm
 We utilize the equation form Equation (10) to enhance Algorithm 1. Compared with the Algorithm 2, FedFDP solely necessitates the application of differential privacy to gradient and loss.
-  • DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$.
-Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 2 and execute the gradient decent.
-  • DP for local loss (lines 2-2): As described in 5.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
-6. Optimal Fairness Parameter
-In this section, we examine the convergence of the FedFDP algorithm to identify an optimal fairness parameter $\lambda$ that minimizes the convergence bound.
-6.1. Convergence Analysis
-We analyze the convergence of Algorithm 2 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound.
-Prior to proving this, we need to establish several assumptions:
-Assumption 1.
-$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
-Assumption 2.
-$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
-Assumption 3.
-Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random.
-The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=[1,\cdots,N]$ , $t=[1,\cdots,T-1]$ and $j=[1,\cdots,|\mathcal{B}_{i}|]$
-Theorem 1.
-Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 2 as follows:
-(18) $\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$
-where:
-  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
-  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
-  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
-  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
-The meaning of $\Gamma$ is consistent with (Li et al., 2020a), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
-proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$.
-Finally, by utilizing Assumption 1, we derive (18).
-For the detailed proof, please refer to Appendix A.1333Due to the lengthy proof, please refer to https://anonymous.4open.science/r/FedFDP-5607/APPENDIX/FedFDP_appendix.pdf for the appendix..
-6.2. Optimal $\lambda$
-Theorem 2.
-There exists a computable optimal $\lambda$, denoted as $\lambda^{*}$, which ensures the fastest convergence of the Algorithm 2.
-Proof.
-When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},x_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{i}^{j}$.
-Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$.
-We rearrange the convergence upper bound in Equation (18) to obtain a function in terms of $\lambda$,
-(19) $P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$
-where $a_{1}=G^{2}Q_{1}^{3}$, $a_{2}=6G^{2}Q_{1}^{2}$, $a_{3}=9Q_{1}G^{2}+2L\Gamma Q_{1}+2Q_{0}\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{4}=4G^{2}+2L\Gamma+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{5}=2Q_{0}$.
-Let $\mathcal{F}(\lambda)=\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1}$, then $P(\lambda)=\frac{L}{2\mu T}\mathcal{F}(\lambda)$, where $\frac{L}{2\mu T}$ can be considered a constant term.
-Obviously, when $\mathcal{F}(\lambda)$ is minimized, $P(\lambda)$ is also minimized.
-The derivative of $\mathcal{F}(\lambda)$ is as follows:
-$\mathcal{F}^{\prime}(\lambda)$ $=\frac{(3a_{1}\lambda^{2}+2a_{2}\lambda+a_{3})(a_{5}\lambda+1)}{(a_{5}\lambda+1)^{2}}-\frac{a_{5}(a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4})}{(a_{5}\lambda+1)^{2}}$
-(20) $=\frac{2a_{1}a_{5}\lambda^{3}+(a_{2}a_{5}+3a_{1})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}}{(a_{5}\lambda+1)^{2}}.$
-Let the denominator of $\mathcal{F}^{\prime}(\lambda)$ be a new function:
-(21) $\mathcal{G}(\lambda)=2a_{1}a_{5}\lambda^{3}+(3a_{1}+a_{2}a_{5})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}.$
-Since $(a_{5}\lambda+1)^{2}>0$, the sign of $\mathcal{F}^{\prime}(\lambda)$ is consistent with that of $\mathcal{G}(\lambda)$.
-Using the discriminant of a cubic function $\Delta=b^{2}-3ac$ (a general cubic function is expressed as $a\lambda^{3}+b\lambda^{2}+c\lambda+d$),
-$\Delta$ $=(3a_{1}+a_{2}a_{5})^{2}-3(2a_{1}a_{5}\cdot 2a_{2})$
-$=9a_{1}^{2}+a_{2}^{2}a_{5}^{2}+6a_{1}a_{2}a_{5}-12a_{1}a_{2}a_{5}$
-(22) $=(3a_{1}-a_{2}a_{5})^{2}\geq 0.$
-The derivative of $\mathcal{G}(\lambda)$ is given by
-(23) $\mathcal{G}^{\prime}(\lambda)=6a_{1}a_{5}\lambda^{2}+2(a_{2}a_{5}+3a_{1})\lambda+2a_{2}.$
-It is evident that when $\lambda>0$, $\mathcal{G}^{\prime}(\lambda)>0$.
-If $\mathcal{F}(\lambda)$ has an extremum, then $\mathcal{F}^{\prime}(\lambda)$ has a root, meaning $\mathcal{G}(\lambda)$ has a root.
-Substituting $\lambda=0$, $\mathcal{G}(0)=a_{3}-a_{4}a_{5}$, evidently when $a_{3}<a_{4}a_{5}$, $\mathcal{G}(\lambda)$ has a root.
-Let $\lambda=x-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, replacing $\lambda$ with $x$ in $\mathcal{G}(\lambda)$ yields
-(24) $\mathcal{G}(x)=x^{3}+a_{6}x+a_{7},$
-where
-  • $a_{6}=\frac{-(3a_{1}-a_{2}a_{5})^{2}}{12a_{1}^{2}a_{5}^{2}}$,
-  • $a_{7}=\frac{108a_{1}^{2}a_{5}^{2}(a_{3}-a_{4}a_{5})-36(a_{1}a_{2}a_{5})(3a_{1}+a_{2}a_{5})+2(3a_{1}+a_{2}a_{5})^{3}}{216a_{1}^{3}a_{5}^{3}}$.
-Using the root-finding formula to calculate $\mathcal{G}(x)=0$ gives the solutions $x_{1},x_{2},x_{3}$.
-[…]
-[…]
-[…]
-where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{1,2,3}=x_{1,2,3}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$.
-Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root.
-If the equation has $k$ real roots, the optimal solution for $\lambda$ is $\lambda^{*}=\max_{k}(\lambda_{k})$.
-In conclusion, $\mathcal{F}^{\prime}(\lambda)\geq 0\iff G(\lambda)\geq 0\iff\lambda\geq\lambda^{*}$, therefore $\mathcal{F}(\lambda)$ has a minimum value $\mathcal{F}(\lambda^{*})$, and $P(\lambda)$ has a minimum value $P(\lambda^{*})$.
-7. Privacy Analysis
-As Algorithm 2 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$.
-$\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of local client, so they are both to perform differential privacy preservation, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively.
-We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
-7.1. Privacy Loss of $\mathbf{w}_{t+1}^{i}$
-Theorem 1.
-After $T$ rounds local updates, the RDP of the $\mathbf{w}_{T}^{i}$ in $i$-th client satisfies:
-(25) $R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$
-where $\sigma$ is noise multiplier of the $\mathbf{w}_{t+1}^{i}$, and $\alpha>1$ is the order.
-The detail of proof please refer Appendix A.2.
-7.2. Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$
-Theorem 2.
-After $T$ rounds local updates, the RDP of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ in $i$-th client satisfies:
-(26) $R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$
-where $\sigma_{l}$ is noise multiplier of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$, and $\alpha>1$ is the order.
-The proof is similar to that of the $\mathbf{w}_{t+1}^{i}$, so we omit it.
-7.3. Privacy Loss of FedFDP
-Since both $\mathbf{w}_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 3, and then use Lemma 4 to convert it to $(\epsilon,\delta)$-DP.
-Lastly, we can get the Privacy loss of FedFDP as follows.
-Theorem 3.
-(Privacy loss of FedFDP). The privacy loss in $i$-th client of FedFDP satisfies:
-[…]
-where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 1, and $R_{loss}^{i}(\alpha)$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 2.
-Lemma 0.
-(Conversion from RDP to DP (Balle et al., 2020)).
-if a randomized mechanism $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ satisfies $(\alpha,R)$-RDP , then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
-8. Experiments
-Table 2. FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
-Datasets MNIST FashionMNIST CIFAR10
-Accuracy Fairness Accuracy Fairness Accuracy Fairness
-FedAvg 98.67$\pm$0.09 7.3e-3$\pm$5.4e-3 87.05$\pm$1.02 3.6e-1$\pm$2.2e-1 62.03$\pm$1.00 1.1e0$\pm$8.9e-1
-SCAFFLOD 98.45$\pm$0.03 3.0e-2$\pm$2.6e-2 86.77$\pm$0.25 6.1e0$\pm$4.0e0 62.21$\pm$1.37 9.9e0$\pm$1.1e1
-FedProx 98.70$\pm$0.13 7.5e-3$\pm$5.1e-3 87.30$\pm$0.51 3.5e-1$\pm$2.1e-1 62.13$\pm$0.20 1.2e0$\pm$3.1e-1
-FedDyn 98.40$\pm$0.23 8.0e-3$\pm$2.1e-3 87.48$\pm$0.31 6.2e-1$\pm$3.5e-1 62.33$\pm$1.15 3.2e0$\pm$8.8e-1
-q-FFL 98.72$\pm$0.03 4.7e-3$\pm$1.0e-3 87.35$\pm$0.13 5.1e-1$\pm$5.2e-2 63.33$\pm$0.21 9.4e-1$\pm$3.3e-2
-FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
-Table 3. FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
+  • DPSGD (lines 1-1): In lines 1-1, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$.
+Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 1 and execute the gradient decent.
+  • DP for local loss (lines 1-1): As described in 5.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
+6 Finding Optimal Fairness Parameter
+In this section, we discuss how to identify an optimal fairness parameter $\lambda$ through the convergence analysis of the FedFDP algorithm.
 Datasets MNIST FashionMNIST CIFAR10
@@
 FedFDP 95.13$\pm$0.83 2.3e10$\pm$1.0e10 85.99$\pm$0.76 2.8e8$\pm$0.8e8 54.21$\pm$0.98 2.6e9$\pm$1.6e9
-Table 4. FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
+Table 1: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
+6.1 Convergence Analysis
+We analyze the convergence of Algorithm 1 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound.
+Prior to proving this, we need to establish several assumptions:
+Assumption 1
+$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
+Assumption 2
+$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
+Assumption 3
+Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random.
+The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=[1,\cdots,N]$ , $t=[1,\cdots,T-1]$ and $j=[1,\cdots,|\mathcal{B}_{i}|]$
+Theorem 1
+Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 1 as follows:
+$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (14)
+where:
+  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
+  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
+  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
+  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
+The meaning of $\Gamma$ is consistent with (Li et al. 2020c), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
+proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$.
+Finally, by utilizing Assumption 1, we derive Equation (14).
+For the detailed proof, please refer to Appendix.
 Herterogeneity Scalability
 Datasets MNIST FashionMNIST CIFAR10
-Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
-FedAvg 98.67 |||| 7.3e-3 98.37 |||| 1.2e-3 89.41 |||| 3.6e-2 62.03 |||| 1.1e0 60.89 |||| 1.5e0 59.78 |||| 1.5e0
-SCAFFLOD 98.45 |||| 5.0e-3 98.51 |||| 2.1e-3 86.39 |||| 3.4e-2 62.21 |||| 9.9e0 59.48 |||| 8.7e0 58.64 |||| 2.3e0
-FedProx 98.70 |||| 7.5-3 98.35 |||| 1.2e-3 88.33 |||| 3.5e-2 62.13 |||| 1.2e0 60.33 |||| 2.2e0 61.31 |||| 1.6e0
-FedDyn 98.40 |||| 8.0e-3 98.36 |||| 1.8e-3 88.26 |||| 3.5e-2 62.33 |||| 9.4e-1 61.35 |||| 2.6e0 62.52 |||| 1.5e0
-q-FFL 98.72 |||| 4.7e-3 98.61 |||| 1.1e-3 89.03 |||| 3.6e-2 63.33 |||| 9.4e-1 60.63 |||| 1.3e0 59.65 |||| 1.5e0
-FedFair 98.75 |||| 3.9e-3 98.72 |||| 1.2e-3 89.60 |||| 1.7e-2 62.38 |||| 8.5e-1 61.58 |||| 1.1e0 62.90 |||| 1.5e0
-Table 5. FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
-Herterogeneity Scalability
-Datasets MNIST FashionMNIST CIFAR10
-Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
+Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
 FedAvg 93.40 |||| 1.1e11 94.17 |||| 4.1e9 85.60 |||| 3.7e9 52.61 |||| 6.1e9 51.58 |||| 2.0e9 55.90 |||| 1.5e9
@@
 FedFDP 94.13 |||| 2.3e10 94.56 |||| 3.4e9 86.77 |||| 6.0e8 54.21 |||| 2.6e9 54.52 |||| 2.2e8 56.49 |||| 5.3e8
-[…]
+Table 2: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
 $\epsilon$ 1.0 2.0 3.0 4.0
@@
 FedFDP 63.36 1.0e6 83.15 2.1e7 85.59 9.5e7 86.97 3.8e8
-This section outlines our experimental framework, beginning with the presentation of baselines and default parameter configurations.
-We then detail the experimental assessment of FedFair and FedFDP across dimensions of performance, fairness, heterogeneity, and scalability.
-Ablation studies focusing on the hyperparameters $\lambda$, $C$ and $\epsilon$ of FedFDP are also included.
-All tasks will be repeated five times using different random seeds, and the mean and standard deviation will be reported.
-8.1. Default Settings
-We conducted experiments on three computer vision (CV) datasets: MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), and CIFAR10 (Krizhevsky, 2009).
-And we comparised our algorithms to six baselines:
-  1. (1)
-    FedAvg(McMahan et al., 2017): FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
-  2. (2)
-    SCAFFLOD(Karimireddy et al., 2020): SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
-  3. (3)
-    FedProx(Li et al., 2020b): FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
-  4. (4)
-    FedDyn(Acar et al., 2021): FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
-  5. (5)
-    ALI-DPFL(Ling et al., 2023): ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
-  6. (6)
-    q-FFL(Li et al., 2020c): q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
-For all datasets, we adopted a widely used heterogeneous settings (Lin et al., 2020; Li et al., 2021a) to 10 clients which controlled by a Dirichlet distribution denoted as $Dir(\beta)$, and the default value of $\beta=0.1$ (Lin et al., 2020; Wang et al., 2020).
-We use a 4-layer CNN architecture (McMahan et al., 2017) which consists of two convolutional layers and two fully connected layers as model architecture.
-In hyper-parameters, for FedFair and its baseline algorithms, we set learning rate $\eta=0.1$.
-For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$.
-In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
-8.2. Results
-In this subsection, we systematically evaluate the performance and fairness of the FedFair and FedFDP under varying heterogeneity levels, expanding from $Dir(0.1)$ to both $Dir(0.5)$ and $Dir(1)$.
-We also assess their scalability by scaling the number of clients from the standard 10 to encompass 20 and 50 clients.
-Furthermore, we examine the influence of three pivotal hyperparameters within the FedFDP algorithm, namely, the adjustments to $\lambda$, $C$ and $\epsilon$.
-8.2.1. FedFair
-Performance and Fairness.
-Table 2 presents the experimental results of the FedFair algorithm compared to baseline methods.
-FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
-Heterogeneity.
-Table 4 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity.
-Except for the $Dir(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $Dir(0.1)$ and $Dir(1.0)$, respectively.
-Scalability. Table 4 indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
-8.2.2. FedFDP
+[…]
+6.2 Optimal Fairness Parameter $\lambda$
+Based on the convergence analysis, we find the “best” fairness parameter $\lambda$ that can minimize the convergence bound. We have the following theorem to do so:
+Theorem 2
+There exists a computable optimal $\lambda$, denoted as $\lambda^{*}$, which ensures the fastest convergence of the Algorithm 1.
+proof sketch: By reorganizing the convergence upper bound of Equation (14) into a function of $\lambda$:
+$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$ (15)
+By using numerical analysis methods, the minimum value $P(\lambda)$ can be solved. For the meaning of $a_{1,2,3,4,5}$ and the proof process, please refer to the Appendix.
+7 Privacy Analysis
+Because each client has the same privacy loss in every round.
+As a result, we only assess each client’s privacy loss from the perspective of a single client.
+Following that, we examine the privacy loss of FedPDF for the $i$-th client as follows.
+Theorem 3
+(Privacy loss of FedFDP). The privacy loss in $i$-th client of FedFDP satisfies:
+[…]
+where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 4 in Appendix, and $R_{loss}^{i}(\alpha$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 5 in Appendix.
+The details of the proof please refer to Appendix.
+8 Experiments
+This section outlines our experimental framework.
+We detail the experimental assessment of FedFDP across dimensions of performance, fairness, heterogeneity, and scalability, including ablation experiments focusing on the fairness hyperparameter $\lambda$, privacy budget $\epsilon$, clipping bound $C$ and noise multiplier $\sigma$.
+In the appendix, we also provide an experimental assessment of FedFair across the same dimensions of performance, fairness, heterogeneity, and scalability.
+For an introduction to baselines, default values of hyperparameters, and the running environment, please refer to the Appendix.
+All tasks were repeated five times and report the mean and standard deviation.
+8.1 Results
 Performance and Fairness.
@@
 Table 2 shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity.
-Except for the $Dir(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $Dir(0.1)$ and $Dir(1.0)$, respectively.
+Except for the $\mathrm{Dir}(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
 Scalability.
@@
 Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
-Figure 3. Impact of different $\lambda$ in FedFDP algorithm
-[…]
-Figure 4. Impact of different $C$ in FedFDP algorithm
-[…]
+Figure 3: Impact of different $\lambda$ in FedFDP algorithm.
+Figure 4: Impact of different $C$ in FedFDP algorithm.
+Figure 5: Impact of different $\sigma$ in FedFDP algorithm.
 Impact of Different $\lambda$.
-We investigated the correlation between $\lambda$ and fairness $\Psi$ in the FedFDP algorithm.
-As depicted in Figure 3, our findings align with the theoretical analysis in Section 6.1: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
+We investigated the correlation between $\lambda$ and $\Psi$ in the FedFDP algorithm.
+As depicted in Figure 3, our findings align with the theoretical analysis: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
 Impact of Different $C$.
-In Section 5.1, we emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously.
+We emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously.
 As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10].
@@
 In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
-9. Conclusion
+Impact of Different $\sigma$.
+We tested the variations in accuracy and fairness on two datasets when $\sigma$ was set to $[1,1.5,2,2.5,3]$, under the condition of $\epsilon=2$ and $q=0.05$, which supports $T=[6,115,268,463,708]$ rounds of communication.
+As shown in Figure 5, an increase in $\sigma$ allows more training epochs to advance the model’s convergence process, but when $\sigma=3$, excessive noise hinders the convergence.
+When $\sigma$ is smaller, the model performance is relatively similar, with a lower $\Psi$ value.
+As $\sigma$ gradually increases, the performance differences between models become apparent, and the $\Psi$ value increases.
+9 Conclusion
 In this paper, a FedFair algorithm is initially proposed to effectively address fairness concerns by optimizing a novel local loss function.
@@
 Appendix A Appendix
-A.1. Proof of Theorem 1
+In the beginning, Table 4 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
+$F(\mathbf{w}$) Global loss function
+$F_{i}(\mathbf{w}$) Local loss function for client $i$
+$T$ Communication rounds
+$t$ Communication round index
+$D_{i}$ Dataset of client $i$
+$D$ The union of $D_{i}$
+$\mathbf{w}_{t}^{i}$ Local model parameter at node $i$ at round $t$
+$\mathbf{w}^{*}$ Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
+$\eta$ Gradient descent step size
+$\lambda$ Fairness parameter
+$N$ Total number of clients
+$q$ batch sample ratio
+$|\mathcal{B}_{i}|$ Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
+$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
+$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
+$\sigma$ Noise multiplier for gradient
+$\sigma_{l}$ Noise multiplier for loss
+$C$ Original clipping bound for gradient
+$C_{l}$ Original Clipping bound for loss
+|||||$\cdot$||| | ⋅ | | $L_{2}$-norm
+Table 4: Summary of main notations
+A.1 Extra Experiment
+We implement our experiment using PyTorch-1.8 and run all experiments on a server with one Intel i9 13900ks CPU (24 cores), 64GB of memory, and one NVIDIA 4090 GPU, running Windows 10.
+Datasets MNIST FashionMNIST CIFAR10
+Accuracy Fairness Accuracy Fairness Accuracy Fairness
+FedAvg 98.67$\pm$0.09 7.3e-3$\pm$5.4e-3 87.05$\pm$1.02 3.6e-1$\pm$2.2e-1 62.03$\pm$1.00 1.1e0$\pm$8.9e-1
+SCAFFLOD 98.45$\pm$0.03 3.0e-2$\pm$2.6e-2 86.77$\pm$0.25 6.1e0$\pm$4.0e0 62.21$\pm$1.37 9.9e0$\pm$1.1e1
+FedProx 98.70$\pm$0.13 7.5e-3$\pm$5.1e-3 87.30$\pm$0.51 3.5e-1$\pm$2.1e-1 62.13$\pm$0.20 1.2e0$\pm$3.1e-1
+FedDyn 98.40$\pm$0.23 8.0e-3$\pm$2.1e-3 87.48$\pm$0.31 6.2e-1$\pm$3.5e-1 62.33$\pm$1.15 3.2e0$\pm$8.8e-1
+q-FFL 98.72$\pm$0.03 4.7e-3$\pm$1.0e-3 87.35$\pm$0.13 5.1e-1$\pm$5.2e-2 63.33$\pm$0.21 9.4e-1$\pm$3.3e-2
+FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
+Table 5: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
+Herterogeneity Scalability
+Datasets MNIST FashionMNIST CIFAR10
+Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
+FedAvg 98.67 |||| 7.3e-3 98.37 |||| 1.2e-3 89.41 |||| 3.6e-2 62.03 |||| 1.1e0 60.89 |||| 1.5e0 59.78 |||| 1.5e0
+SCAFFLOD 98.45 |||| 5.0e-3 98.51 |||| 2.1e-3 86.39 |||| 3.4e-2 62.21 |||| 9.9e0 59.48 |||| 8.7e0 58.64 |||| 2.3e0
+FedProx 98.70 |||| 7.5-3 98.35 |||| 1.2e-3 88.33 |||| 3.5e-2 62.13 |||| 1.2e0 60.33 |||| 2.2e0 61.31 |||| 1.6e0
+FedDyn 98.40 |||| 8.0e-3 98.36 |||| 1.8e-3 88.26 |||| 3.5e-2 62.33 |||| 9.4e-1 61.35 |||| 2.6e0 62.52 |||| 1.5e0
+q-FFL 98.72 |||| 4.7e-3 98.61 |||| 1.1e-3 89.03 |||| 3.6e-2 63.33 |||| 9.4e-1 60.63 |||| 1.3e0 59.65 |||| 1.5e0
+FedFair 98.75 |||| 3.9e-3 98.72 |||| 1.2e-3 89.60 |||| 1.7e-2 62.38 |||| 8.5e-1 61.58 |||| 1.1e0 62.90 |||| 1.5e0
+Table 6: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
+Default Settings
+We conducted experiments on three computer vision (CV) datasets: MNIST (LeCun et al.
+1998), FashionMNIST (Xiao, Rasul, and Vollgraf 2017), and CIFAR10 (Krizhevsky 2009).
+And we comparised our algorithms to six baselines:
+  1. 1.
+    FedAvg(McMahan et al. 2017): FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
+  2. 2.
+    SCAFFLOD(Karimireddy et al. 2020): SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
+  3. 3.
+    FedProx(Li et al. 2020a): FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
+  4. 4.
+    FedDyn(Acar et al.
+2021): FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
+  5. 5.
+    ALI-DPFL(Ling et al.
+2024): ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
+  6. 6.
+    q-FFL(Li et al.
+2020b): q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
+For all datasets, we adopted a widely used heterogeneous settings (Lin et al.
+2020; Li, He, and Song 2021) to 10 clients which controlled by a Dirichlet distribution denoted as $\mathrm{Dir}(\beta)$, and the default value of $\beta=0.1$ (Lin et al.
+2020; Wang et al.
+2020).
+We use a 4-layer CNN architecture (McMahan et al.
+2017) which consists of two convolutional layers and two fully connected layers as model architecture.
+In hyper-parameters, for FedFair and its baseline algorithms, we set learning rate $\eta=0.1$.
+For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$.
+In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
+Extended description of FedFair
+Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, fairness hyperparameter $\lambda$, batch sample ratio $q$, local dataset $D_{i}$.
+Output: the final trained model $\mathbf{w}_{T}$
+1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
+2 SERVER
+3       for $t\in[0,1,\cdots,T-1]$  do
+4             for $i\in[1,2,\cdots,N]$ parallel  do
+5                   $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
+6
+7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
+8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t+1}^{i})$;
+9
+10      return $\mathbf{w}_{T}$;
+11
+12LOCAL
+13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
+14
+15      $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
+16       for batch $b\in\mathcal{B}_{i}$ do
+17             $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i},b)-F(\mathbf{w}_{t})$;
+18             $\eta_{i}=\eta\cdot\left(1+\lambda\cdot\Delta_{i}\right)$;
+19             $\mathbf{w}_{t}^{i}=\mathbf{w}_{t}^{i}-\eta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},b)$;
+20
+21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ , compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
+22       return $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})$;
+23
+Algorithm 2 FedFair
+  • Server (lines 2-2): The server executes $T$ rounds of communication, each round sending the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
+  • Local (lines 2-2): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 2-2. For line 2, we execute gradient descent by using $\eta_{i}$.
+Result of FedFair
+Furthermore, as the training process advances, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$.
+In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods (Zinkevich 2003).
+Performance and Fairness.
+Table 5 presents the experimental results of the FedFair algorithm compared to baseline methods.
+FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
+Heterogeneity.
+Table 6 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity.
+Except for the $\mathrm{Dir}(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
+Scalability. Table 6 indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
+A.2 Proof of Theorem 1
 For the purpose of validation, we introduce an additional variable $\mathbf{v}_{t}^{i}$ to represent the immediate result of a single-step DPSGD update from $\mathbf{w}_{t}^{i}$.
@@
 Consequently, the fair-clipping DPSGD in client $i$ at iteration $t$ transitions from Equation (10) to:
-(28) $\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$
+$\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (17)
 where:
 $C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
-In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by (Stich, 2018).
+In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by (Stich 2018).
 Therefore,
 $\mathbf{v}_{t+1}$ $=\mathbf{w}_{t}$
-(29) $\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
-A.1.1. Key Lemma
-Lemma 0.
+$\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (18)
+Key Lemma
+Lemma 1
 (Results of one iteration.)Assume Assumption 1-3 hold, we have:
@@
 $\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2}\Delta_{t}\leq\frac{L}{2t}v\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\Delta_{1}\right)$
-A.1.2. Proof of Lemma 1
+Proof of Lemma 1
 By the Equation (A.2), we get
@@
 where L-smooth inference as show:
-(30) $\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$
+$\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$ (19)
 Thus, we get
@@
 $A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
-A.2. Proof of Theorem 1
-$proof.$ We will prove Theorem 1 in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
-Definition 0.
-(RDP privacy budget of SGM(Mironov et al., 2019)).
+A.3 Proof of Theorem 2
+When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},x_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{i}^{j}$.
+Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$.
+We rearrange the convergence upper bound in Equation (14) to obtain a function in terms of $\lambda$,
+$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$ (20)
+where $a_{1}=G^{2}Q_{1}^{3}$, $a_{2}=6G^{2}Q_{1}^{2}$, $a_{3}=9Q_{1}G^{2}+2L\Gamma Q_{1}+2Q_{0}\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{4}=4G^{2}+2L\Gamma+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{5}=2Q_{0}$.
+Let $\mathcal{F}(\lambda)=\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1}$, then $P(\lambda)=\frac{L}{2\mu T}\mathcal{F}(\lambda)$, where $\frac{L}{2\mu T}$ can be considered a constant term.
+Obviously, when $\mathcal{F}(\lambda)$ is minimized, $P(\lambda)$ is also minimized.
+The derivative of $\mathcal{F}(\lambda)$ is as follows:
+$\mathcal{F}^{\prime}(\lambda)=\frac{2a_{1}a_{5}\lambda^{3}+(a_{2}a_{5}+3a_{1})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}}{(a_{5}\lambda+1)^{2}}.$ (21)
+Let the denominator of $\mathcal{F}^{\prime}(\lambda)$ be a new function:
+$\mathcal{G}(\lambda)=2a_{1}a_{5}\lambda^{3}+(3a_{1}+a_{2}a_{5})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}.$ (22)
+Since $(a_{5}\lambda+1)^{2}>0$, the sign of $\mathcal{F}^{\prime}(\lambda)$ is consistent with that of $\mathcal{G}(\lambda)$.
+Using the discriminant of a cubic function $\Delta=b^{2}-3ac$ (a general cubic function is expressed as $a\lambda^{3}+b\lambda^{2}+c\lambda+d$),
+$\Delta$ $=(3a_{1}+a_{2}a_{5})^{2}-3(2a_{1}a_{5}\cdot 2a_{2})$
+$=9a_{1}^{2}+a_{2}^{2}a_{5}^{2}+6a_{1}a_{2}a_{5}-12a_{1}a_{2}a_{5}$
+$=(3a_{1}-a_{2}a_{5})^{2}\geq 0.$ (23)
+The derivative of $\mathcal{G}(\lambda)$ is given by
+$\mathcal{G}^{\prime}(\lambda)=6a_{1}a_{5}\lambda^{2}+2(a_{2}a_{5}+3a_{1})\lambda+2a_{2}.$ (24)
+It is evident that when $\lambda>0$, $\mathcal{G}^{\prime}(\lambda)>0$.
+If $\mathcal{F}(\lambda)$ has an extremum, then $\mathcal{F}^{\prime}(\lambda)$ has a root, meaning $\mathcal{G}(\lambda)$ has a root.
+Substituting $\lambda=0$, $\mathcal{G}(0)=a_{3}-a_{4}a_{5}$, evidently when $a_{3}<a_{4}a_{5}$, $\mathcal{G}(\lambda)$ has a root.
+Let $\lambda=x-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, replacing $\lambda$ with $x$ in $\mathcal{G}(\lambda)$ yields
+$\mathcal{G}(x)=x^{3}+a_{6}x+a_{7},$ (25)
+where
+  • $a_{6}=\frac{-(3a_{1}-a_{2}a_{5})^{2}}{12a_{1}^{2}a_{5}^{2}}$,
+  • $a_{7}=\frac{108a_{1}^{2}a_{5}^{2}(a_{3}-a_{4}a_{5})-36(a_{1}a_{2}a_{5})(3a_{1}+a_{2}a_{5})+2(3a_{1}+a_{2}a_{5})^{3}}{216a_{1}^{3}a_{5}^{3}}$.
+Using the root-finding formula to calculate $\mathcal{G}(x)=0$ gives the solutions $x_{1},x_{2},x_{3}$.
+[…]
+[…]
+[…]
+where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{1,2,3}=x_{1,2,3}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$.
+Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root.
+If the equation has $k$ real roots, the optimal solution for $\lambda$ is $\lambda^{*}=\max_{k}(\lambda_{k})$.
+In conclusion, $\mathcal{F}^{\prime}(\lambda)\geq 0\iff G(\lambda)\geq 0\iff\lambda\geq\lambda^{*}$, therefore $\mathcal{F}(\lambda)$ has a minimum value $\mathcal{F}(\lambda^{*})$, and $P(\lambda)$ has a minimum value $P(\lambda^{*})$.
+A.4 Proof of Theorem 3
+As Algorithm 1 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$.
+$\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of local client, so they are both to perform differential privacy preservation, as in lines 1-1 and lines 1-1 of the Algorithm 1, respectively.
+We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
+Privacy Loss of $\mathbf{w}_{t+1}^{i}$.
+Theorem 4
+After $T$ rounds local updates, the RDP of the $\mathbf{w}_{T}^{i}$ in $i$-th client satisfies:
+$R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$ (26)
+where $\sigma$ is noise multiplier of the $w_{t+1}^{i}$, and $\alpha>1$ is the order.
+$proof.$ We will prove this in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
+Definition 3
+(RDP privacy budget of SGM(Mironov, Talwar, and Zhang 2019)).
 Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$.
 If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
-(31) $R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$
+$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$ (27)
 where
-(32) $\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$
+$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$ (28)
 with $\vartheta_{0}\triangleq\mathcal{N}\left(0,\sigma^{2}\right),\vartheta_{1}\triangleq\mathcal{N}\left(1,\sigma^{2}\right)\mbox{ and }\vartheta\triangleq(1-q)\vartheta_{0}+q\vartheta_{1}$
@@
 Thus, $SG_{q,\sigma}\text{ satisfies }\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
-Finally, the existing work (Mironov et al., 2019) describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
-(33) $A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$
-Definition 0.
-(Composition of RDP(Mironov, 2017)).
+Finally, the existing work (Mironov, Talwar, and Zhang 2019) describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
+$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$ (29)
+Definition 4
+(Composition of RDP(Mironov 2017)).
 For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
 From Definition 3 and Definition 4, the Theorem 4 is obtained.
+Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$.
+Theorem 5
+After $T$ rounds local updates, the RDP of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ in $i$-th client satisfies:
+$R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$ (30)
+where $\sigma_{l}$ is noise multiplier of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$, and $\alpha>1$ is the order.
+The proof is similar to that of the $w_{t+1}^{i}$, so we omit it.
+Proof Loss of FedFDP.
+Since both $w_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 4, and then use Lemma 2 to convert it to $(\epsilon,\delta)$-DP.
+Therefore, the Theorem 3 is hold.
+Lemma 2
+(Conversion from RDP to DP (Balle et al.
+2020)).
+if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.

### v4

@@
 Abstract
-Federated learning (FL) is a new machine learning paradigm to overcome the challenge of data silos and has garnered significant attention.
-However, federated learning faces challenges in fairness and data privacy.
-To address both of the above challenges simultaneously, we first propose a fairness-aware federated learning algorithm, termed FedFair.
-Then based on FedFair, we introduce differential privacy protection to form the FedFDP algorithm to address the trade-offs among fairness, privacy protection, and model performance.
-In FedFDP, we designed an fairness-aware gradient clipping technique to identify the relationship between fairness and differential privacy.
-Through convergence analysis, we determined the optimal fairness adjustment parameters to simultaneously achieve the best model performance and fairness.
-Additionally, for the extra uploaded loss values, we present an adaptive clipping method to minimize privacy budget consumption.
-Extensive experimental results demonstrate that FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness.
-Codes and datasets will be made public after acceptance.
-1 Introduction
-Federated Learning (FL) (McMahan et al.
-2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without exposing their respective datasets.
+Federated learning (FL) is an emerging machine learning paradigm designed to address the challenge of data silos, attracting considerable attention.
+However, FL encounters persistent issues related to fairness and data privacy.
+To tackle these challenges simultaneously, we propose a fairness-aware federated learning algorithm called FedFair.
+Building on FedFair, we introduce differential privacy to create the FedFDP algorithm, which addresses trade-offs among fairness, privacy protection, and model performance.
+In FedFDP, we developed a fairness-aware gradient clipping technique to explore the relationship between fairness and differential privacy.
+Through convergence analysis, we identified the optimal fairness adjustment parameters to achieve both maximum model performance and fairness.
+Additionally, we present an adaptive clipping method for uploaded loss values to reduce privacy budget consumption.
+Extensive experimental results show that FedFDP significantly surpasses state-of-the-art solutions in both model performance and fairness.
+Index Terms:
+Federated Learning, Differential Privacy, Fairness
+I Introduction
+Federated Learning (FL) [33] is a distributed machine learning framework that allows clients to collaboratively train a shared model without exposing their respective datasets.
 After training their local models, each client only transmits model parameters to the central server, rather than the original training data.
 The server aggregates these parameters to update the global model, which is then sent back to the clients for further training.
-Due to its ability to improve model performance through collaboration without the need to upload original data, overcoming data silos, FL has received significant attention in recent years (Zhu, Hong, and Zhou 2021; Lee et al.
-2022; Chen et al.
-2024).
-Particularly, many studies have achieved significant breakthroughs in terms of model performance (Wang et al.
-2020; Li et al.
-2020a, 2021b) and communication costs (Wang et al.
-2019; Ling et al.
-2024) in FL.
+Due to its ability to improve model performance through collaboration without the need to upload original data, overcoming data silos, FL has received significant attention in recent years [54, 23, 6].
+Particularly, many studies have achieved significant breakthroughs in terms of model performance [46, 27, 26] and communication costs [47, 31] in FL.
 Figure 1: Client drift caused by Non-IID.
-To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary (Li et al.
-2021a).
-Current fairness-aware machine learning algorithms, such as (Dwork et al.
-2012; Cotter et al.
-2019), have not fully addressed the issue of client performance disparities in federated learning.
-As shown in Figure 1, due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (yellow squares) lead to significant performance imbalances of the collaboratively trained global model (red square) across different datasets held by various clients.
-An increasing number of scholars are beginning to focus on this phenomenon of unfair performance disparity (Li et al.
-2020b, 2021b; Zhao and Joshi 2022).
-On the other hand, recent studies have pointed out that training parameters may leak data privacy, such as recovering original data (Zhu, Liu, and Han 2019) or member inference attacks (Song, Ristenpart, and Shmatikov 2017).
-Even if only models or gradients are uploaded instead of original data, data leakage can still occur (Melis et al.
-2022).
-To enhance data privacy in FL, differential privacy (DP) (Dwork, Roth et al.
-2014) has become the preferred method for protecting privacy in federated learning due to its moderate computational overhead and solid mathematical foundation.
-Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information (Abadi et al.
-2016; Carlini et al.
-2019; Feldman 2020).
-Although some works have explored fairness and differential privacy in federated learning separately, no prior research has considered them in a unified framework.
+To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary [25].
+Current fairness-aware machine learning algorithms, such as [10, 8], have not fully addressed the issue of client performance disparities in federated learning.
+As shown in Figure 1, due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (orange squares) lead to significant performance disparities of the collaboratively trained global model (red square) across different datasets held by various clients.
+An increasing number of scholars are beginning to focus on this phenomenon of unfair performance disparity [28, 26, 50].
+On the other hand, recent studies have pointed out that training parameters may leak data privacy, such as recovering original data [53] or member inference attacks [42].
+Even if only models or gradients are uploaded instead of original data, data leakage can still occur [35].
+To enhance data privacy in FL, differential privacy (DP) [11] has become the preferred method for protecting privacy in federated learning due to its moderate computational overhead and solid mathematical foundation.
+Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information [1, 5, 12].
+Although some works have explored balanced performance fairness and differential privacy in federated learning separately, no prior research has considered them in a unified framework.
 For the first time, we propose a fairness-aware federated learning with differential privacy.
 However, balancing differential privacy protection, fairness, and model performance presents a significant challenge.
-Specifically, there are two main challenges: First, the clipping and noise addition steps in differential privacy can impact fairness and model utility in federated learning.
-Second, fair federated learning requires uploading loss information to the central server, which incurs additional privacy budget consumption.
-In this paper, to address the first challenge, we analyze the relationship between the fairness loss function and differential privacy in federated learning.
+Specifically, there are two main challenges: First, the clipping and noise addition processes in differential privacy have the potential to affect both fairness and model utility in federated learning.
+Moreover, achieving fairness in federated learning requires uploading loss information to the server, which entails additional privacy budget consumption.
+In this paper, the relationship between the fairness loss function and differential privacy in federated learning is analyzed to address the first challenge.
 We design an fairness-aware gradient clipping strategy to balance fairness and differential privacy.
@@
 Our main contributions are as follows:
-  • First, we propose a fairness-aware federated algorithm named FedFair. It combines a novel loss function aimed at optimizing the objectives of fairness and model performance simultaneously.
+  1. 1.
+    First, a fairness-aware federated algorithm, FedFair, is proposed. It integrates a novel loss function specifically designed to simultaneously optimize fairness and model performance.
+  2. 2.
     Based on FedFair, we propose an algorithm called FedFDP to further equip the system with differential privacy.
@@
 In addition, we propose an adaptive clipping method for the additional loss values uploaded by each client to achieve optimal utility.
-  • Further, we conducted a convergence analysis of FedFDP and found the optimal fairness parameter $\lambda$ by minimizing the upper bound of convergence.
+  3. 3.
+    In addition, we conducted a convergence analysis of FedFDP and identified the fairness parameter $\lambda^{*}$ that results in the fastest convergence.
 We analyzed the privacy loss of the FedFDP algorithm using the concept of Rényi Differential Privacy (RDP) to verify its compliance with differential privacy requirements.
+  4. 4.
     Finally, through comprehensive empirical evaluations on three public datasets, we confirmed that FedFair and FedFDP are superior to existing solutions in terms of model performance and fairness.
 What’s more, we investigated the impact of clipping bound and noise multiplier in differential privacy on fairness.
-2 Related Work
-Differentially private federated learning (DPFL) is widely used for privacy protection in federated learning.
-Many studies have focused on the trade-off between model performance, communication efficiency, and privacy budget in DPFL (Fu, Chen, and Han 2022; Zhou et al.
-2023; Ling et al.
-2024; Fu et al.
-2024).
-In addition that, various fairness concepts in machine learning have been proposed, each emphasizing different aspects and addressing the needs of different stakeholders.
-For example, concepts like Good-Intent Fairness (Mohri, Sivek, and Suresh 2019) and Selection Fairness (Zhou, Fang, and Hui 2021) have been introduced.
-Further discussions on these concepts can be found in additional literature (Cong et al.
-2020; Lyu et al.
-2020) .
-At present, the primary criterion for fairness in FL is group fairness, which aims to mitigate biases associated with sensitive attributes like gender and race.
-Relevant research has shown that in federated learning with differential privacy, adjusting parameters such as the clipping threshold and noise multiplier can help enhance the algorithm’s group fairness (Gu et al.
-2022).
-There is a growing body of literature that addresses the concept of group fairness specifically within the framework of DPFL (Gálvez et al.
-2021; Padala, Damle, and Gujar 2021; Gu et al.
-2022, 2020; Pentyala et al.
-2022).
-The more predominant fairness criterion in FL is balanced performance fairness, as introduced by Li et al.
-(Li et al.
-2020b).
-This approach ensures fairness by assessing the uniformity of model performance across participating clients.
-The concept has gained broad acceptance.
-Building on this, Zhao et al.
-(Zhao and Joshi 2022) proposed an alternative to the loss amplification mechanism, advocating for a direct weight reallocation strategy.
-Concurrently, Li et al.
-(Li et al.
-2021b) concentrated on enhancing the fairness of local models within personalized FL.
-Our study pioneers the investigation of balanced performance fairness within the context of DPFL.
-3 Preliminaries
-3.1 Federated Learning and Fairness
-We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model.
-The objective of FL is to minimize $F(\mathbf{\cdot})$, in other words, to seek:
-$\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (1)
-Here, $F_{i}(\cdot)$ represents the local loss function of client $i$.
-The weights are defined as $p_{i}=\frac{|D_{i}|}{|D|}$, $\sum_{i=1}^{N}p_{i}=1$, $D_{i}$ represents the dataset of client $i$, and $D$ represents the union of all $D_{i}$.
-We have defined balanced performance fairness in FL as follows.
-Definition 1
-(Balanced performance fairness (Li et al.
-2020b)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
-In this study, we employ the loss of model denote the performance.
-Equation (2) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients.
-Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
-$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (2)
-3.2 Differential Privacy
-Differential privacy is a formal mathematical model that quantifies privacy leakage in data analysis algorithms.
-It stipulates that alterations to a single record within the training data should not induce significant statistical variations in the algorithm’s results.
-Definition 2
-(Differential Privacy (Dwork, Roth et al.
-2014)).
-$(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
-$\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$ (3)
-By adding random noise, we can achieve differential privacy for a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ according to Definition 2.
-The $l_{k}$-sensitivity determines how much noise is needed and is defined as follow.
-We use $l_{2}$ sensitivity as in the previous work.
-DPSGD. In gradient-based methods, DPSGD(Abadi et al. 2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
-  • Firstly, computing the gradient of loss function: $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
-  • Secondly, clipping the gradient of per-sample to get the sensitivity: $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right)$, where $C$ represents the clipping bound.
-  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier: $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right)$.
-The final step is to proceed with gradient descent as usual.
-When each client in FL utilizes DPSGD as the optimization algorithm to safeguard their private data, the framework transitions from FL to DPFL.
-4 FedFair: Equipping FL with Fairness
-We previously defined the performance objective in Equation (1) and balanced performance fairness in Definition 1.
-Here, we use these to develop a new objective function that informs our proposed fairness-aware federated learning algorithm.
-Upon incorporating Equation (2) into Equation (1), we introduce a comprehensive objective function that integrates considerations of fairness:
-$\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (4)
-For client $i$ during round $t$, the refined objective is articulated as follows:
-$\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$ (5)
-where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training.
-from round $t-1$, as delineated in line 2 of Algorithm 2, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning.
-When $\lambda\rightarrow\infty$, the algorithm only focus on fairness.
-However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging.
-Fortunately, we theoretically found an optimal value for $\lambda$.
-In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$..
-Consequently, within the context of Equation (5), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed.
-So, the gradient of Equation (5) as follows:
-$\nabla H_{i}(\mathbf{w}_{t}^{i})=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (6)
-where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t}^{i})$.
-The stochastic gradient descent by using Equation (6) as show:
-$\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (7)
-where $\eta_{i}=\eta(1+\lambda\Delta_{i})$.
-Figure 2: FedFDP framework.
-Therefore, our approach can be viewed as establishing an adaptive learning rate that allows the model update to achieve both performance and fairness objectives simultaneously.
-Based on these principles, we introduce the FedFair algorithm, as depicted in Appendix Algorithm 2.
-5 FedFDP: Further Equipping FedFair with Differential Privacy
-In this section, we present our algorithm, FedFDP, which further integrates FedFair with differential privacy.
-The flow of FedFDP algorithm is shown in Figure 2.
-It adds two additional processes to FedFair:
-  • A fair-clipping strategy is added to FedFair. This strategy not only achieves sensitivity but also allows adjustments to the level of fairness.
-  • An adaptive clipping method is employed when protecting the loss values that clients share with differential privacy, aiming to maximize utility.
-Next, we discuss the details of these two strategies.
-1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0}),C_{l}^{0}$;
-2 SERVER
-3       for $t\in[0,1,\cdots,T-1]$  do
-4             for $i\in[1,2,\cdots,N]$ parallel  do
-5                   $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
-6
-7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
-8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
-9
-10      return $\mathbf{w}_{T}$;
-11
-12LOCAL
-13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
-14       Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
-15       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
-16             $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
-17             Compute $C^{i,j}_{t}$ by Equation (11);
-18             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
-19
-20      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right)$;
-21
-22      Compute clipping bound of loss $C_{l}^{i,t}$ by Equation (12) and (13) ;
-23       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
-24             Compute $F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})$;
-25
-26            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
-27
-28      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\mathcal{N}(0,\mathbf{I})\right)$;
-29       return $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
-30
-Algorithm 1 FedFDP
-5.1 Fair-clipping Strategy for Gradient
-In the general DPSGD algorithm (Abadi et al.
-2016; Fu et al.
-2023; Bu et al.
-2022; Yang et al.
-2022), per-sample clipping is an essential process for obtaining sensitivity.
-In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (5), into Equation (8):
-$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$ (8)
-where $\xi_{j}\in\mathcal{B}_{t}^{i}$, $j=\{1,2,\cdots,|\mathcal{B}_{i}|\}$, and $|\mathcal{B}_{i}|$ is the batch size of client $i$.
-Next, we compute the gradient of Equation (8):
-$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})=\left(1+\lambda\cdot\Delta_{i}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$ (9)
-where $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})$.
-To ensure the algorithm is differentially private, we replace the gradient of loss function by Equation (9), and reformulate the computation of the gradient descent for client $i$ as follows:
-$\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (10)
-where:
-$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$ (11)
-Hence, our strategy can be considered as a fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model.
-When $\lambda=0$, Equation (11) corresponds to the traditional DP clipping method without fairness.
-For large $C$, Equation (11) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness.
-Conversely, for small $C$, Equation (11) simplifies to the second term, resulting in the gradient being clipped to the bound $C$.
-The impact of $C$ on the performance of FedFDP will be further discussed in the experiments.
-5.2 Adaptive Clipping Method for Loss
-Since the calculation of $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy.
-As shown in lines 1-1 of Algorithm 1, We need to clip $F_{i}(\mathbf{w}_{t}^{i})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy.
-However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging.
-A large $C_{l}^{i,t}$ avoids loss of fidelity to the original values, but it also leads to large injected noise.
-Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t}^{i})$.
-For the $i$-th client, the adaptive clipping method uses the $differentially$ $private$ $mean$ $of$ $the$ $previous$ $round$ as the clipping bound for the current round.
-The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1}^{i})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
-$C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$ (12)
-where
-$\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))).$ (13)
-To ensure the achievement of differential privacy during the clipping process, we use the noise-added loss as the clipping bound $C_{l}^{i,t}$.
-Each round’s clipping bound adheres to the post-processing nature of differential privacy when $t=\{1,\dots,T\}$.
-5.3 Overall Algorithm
-We utilize the equation form Equation (10) to enhance Algorithm 1. Compared with the Algorithm 2, FedFDP solely necessitates the application of differential privacy to gradient and loss.
-  • DPSGD (lines 1-1): In lines 1-1, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$.
-Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 1 and execute the gradient decent.
-  • DP for local loss (lines 1-1): As described in 5.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
-6 Finding Optimal Fairness Parameter
-In this section, we discuss how to identify an optimal fairness parameter $\lambda$ through the convergence analysis of the FedFDP algorithm.
-Datasets MNIST FashionMNIST CIFAR10
-Accuracy Fairness Accuracy Fairness Accuracy Fairness
-FedAvg 93.40$\pm$0.47 1.1e11$\pm$0.5e10 84.15$\pm$1.97 3.0e8$\pm$1.4e8 52.61$\pm$0.17 6.1e9$\pm$5.2e9
-SCAFFLOD 93.95$\pm$1.39 7.0e10$\pm$4.9e9 83.41$\pm$4.87 3.3e9$\pm$4.7e9 53.85$\pm$1.16 6.2e9$\pm$1.3e9
-FedProx 90.50$\pm$3.95 5.3e10$\pm$7.3e9 83.85$\pm$0.98 7.9e8$\pm$4.5e8 51.34$\pm$0.91 4.4e9$\pm$2.1e9
-FedDyn 91.42$\pm$2.25 5.6e10$\pm$9.8e9 84.04$\pm$1.21 8.8e8$\pm$3.2e8 52.14$\pm$1.39 5.5e9$\pm$1.2e9
-ALI-DPFL 90.89$\pm$1.65 3.3e10$\pm$8.3e9 83.65$\pm$1.68 6.6e8$\pm$3.5e8 52.05$\pm$1.26 3.6e9$\pm$1.6e9
-q-FFL 93.74$\pm$1.41 7.8e10$\pm$1.1e11 83.13$\pm$2.74 4.2e9$\pm$2.6e9 48.46$\pm$1.00 4.7e9$\pm$1.9e9
-FedFDP 95.13$\pm$0.83 2.3e10$\pm$1.0e10 85.99$\pm$0.76 2.8e8$\pm$0.8e8 54.21$\pm$0.98 2.6e9$\pm$1.6e9
-Table 1: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
-6.1 Convergence Analysis
-We analyze the convergence of Algorithm 1 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound.
-Prior to proving this, we need to establish several assumptions:
-Assumption 1
-$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
-Assumption 2
-$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
-Assumption 3
-Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random.
-The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=[1,\cdots,N]$ , $t=[1,\cdots,T-1]$ and $j=[1,\cdots,|\mathcal{B}_{i}|]$
-Theorem 1
-Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 1 as follows:
-$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (14)
-where:
-  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
-  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
-  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
-  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
-The meaning of $\Gamma$ is consistent with (Li et al. 2020c), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
-proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$.
-Finally, by utilizing Assumption 1, we derive Equation (14).
-For the detailed proof, please refer to Appendix.
-Herterogeneity Scalability
-Datasets MNIST FashionMNIST CIFAR10
-Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
-FedAvg 93.40 |||| 1.1e11 94.17 |||| 4.1e9 85.60 |||| 3.7e9 52.61 |||| 6.1e9 51.58 |||| 2.0e9 55.90 |||| 1.5e9
-SCAFFLOD 93.95 |||| 7.0e10 94.06 |||| 1.9e9 84.56 |||| 3.6e9 53.85 |||| 6.2e9 54.34 |||| 2.9e8 53.98 |||| 2.2e9
-FedProx 90.50 |||| 5.3e10 94.23 |||| 3.1e9 84.38 |||| 5.6e9 51.34 |||| 4.4e9 53.67 |||| 4.5e8 54.14 |||| 2.2e9
-FedDyn 91.42 |||| 5.6e10 93.08 |||| 3.8e9 85.21 |||| 9.8e10 52.14 |||| 5.5e9 50.21 |||| 1.2e10 51.86 |||| 2.0e10
-ALI-DPFL 90.89 |||| 3.3e10 93.98 |||| 2.6e9 83.66 |||| 2.5e10 52.05 |||| 3.6e9 38.35 |||| 2.4e10 33.46 |||| 1.6e10
-q-FFL 93.74 |||| 7.8e10 93.86 |||| 2.5e8 85.20 |||| 7.4e8 48.46 |||| 4.7e9 51.12 |||| 2.7e8 49.83 |||| 1.1e7
-FedFDP 94.13 |||| 2.3e10 94.56 |||| 3.4e9 86.77 |||| 6.0e8 54.21 |||| 2.6e9 54.52 |||| 2.2e8 56.49 |||| 5.3e8
-Table 2: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
-$\epsilon$ 1.0 2.0 3.0 4.0
-Accuracy Fairness Accuracy Fairness Accuracy Fairness Accuracy Fairness
-FedAvg 61.68 1.1e6 82.69 2.3e7 84.07 3.8e8 86.83 8.0e8
-SCAFFLOD 62.12 1.2e6 82.36 2.3e7 83.39 2.6e9 85.41 5.6e8
-FedProx 61.25 1.1e6 81.62 2.3e7 85.32 3.8e8 85.68 2.4e9
-FedDyn 62.38 1.5e6 82.43 3.1e7 83.29 2.9e8 84.51 5.2e9
-ALI-DPFL 61.78 1.1e6 83.01 2.3e7 84.12 3.9e8 85.58 4.5e9
-q-FFL 58.99 3.6e6 80.24 9.6e7 82.92 5.5e8 85.01 3.8e9
-FedFDP 63.36 1.0e6 83.15 2.1e7 85.59 9.5e7 86.97 3.8e8
-[…]
-6.2 Optimal Fairness Parameter $\lambda$
-Based on the convergence analysis, we find the “best” fairness parameter $\lambda$ that can minimize the convergence bound. We have the following theorem to do so:
-Theorem 2
-There exists a computable optimal $\lambda$, denoted as $\lambda^{*}$, which ensures the fastest convergence of the Algorithm 1.
-proof sketch: By reorganizing the convergence upper bound of Equation (14) into a function of $\lambda$:
-$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$ (15)
-By using numerical analysis methods, the minimum value $P(\lambda)$ can be solved. For the meaning of $a_{1,2,3,4,5}$ and the proof process, please refer to the Appendix.
-7 Privacy Analysis
-Because each client has the same privacy loss in every round.
-As a result, we only assess each client’s privacy loss from the perspective of a single client.
-Following that, we examine the privacy loss of FedPDF for the $i$-th client as follows.
-Theorem 3
-(Privacy loss of FedFDP). The privacy loss in $i$-th client of FedFDP satisfies:
-[…]
-where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 4 in Appendix, and $R_{loss}^{i}(\alpha$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 5 in Appendix.
-The details of the proof please refer to Appendix.
-8 Experiments
-This section outlines our experimental framework.
-We detail the experimental assessment of FedFDP across dimensions of performance, fairness, heterogeneity, and scalability, including ablation experiments focusing on the fairness hyperparameter $\lambda$, privacy budget $\epsilon$, clipping bound $C$ and noise multiplier $\sigma$.
-In the appendix, we also provide an experimental assessment of FedFair across the same dimensions of performance, fairness, heterogeneity, and scalability.
-For an introduction to baselines, default values of hyperparameters, and the running environment, please refer to the Appendix.
-All tasks were repeated five times and report the mean and standard deviation.
-8.1 Results
-Performance and Fairness.
-Table 1 displays the experimental results of the FedFDP algorithm compared to baseline methods.
-Under a consistent privacy budget of $\epsilon=3.52$, when accuracy matches the best baseline, our method significantly enhances fairness, with improvements of 30.3%, 6.7%, and 27.8% observed across three separate datasets.
-Heterogeneity.
-Table 2 shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity.
-Except for the $\mathrm{Dir}(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
-Scalability.
-Table 2 demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline.
-Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
-Figure 3: Impact of different $\lambda$ in FedFDP algorithm.
-Figure 4: Impact of different $C$ in FedFDP algorithm.
-Figure 5: Impact of different $\sigma$ in FedFDP algorithm.
-Impact of Different $\lambda$.
-We investigated the correlation between $\lambda$ and $\Psi$ in the FedFDP algorithm.
-As depicted in Figure 3, our findings align with the theoretical analysis: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
-Impact of Different $C$.
-We emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously.
-As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10].
-Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition.
-As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$.
-When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
-Different $\epsilon$.
-We conducted experiments with $\epsilon=[1,2,3,4]$ on the FashionMNIST dataset, as detailed in Table 3.
-For $\epsilon=[1,2,3]$, we allowed larger values of $\epsilon$ to support more iterations.
-At $\epsilon=4$, we supported a smaller noise multiplier $\sigma=1.65$ with a similar number of iterations as at $\epsilon=3$.
-The experimental results demonstrate that FedFDP consistently achieves performance on par with the best baseline in terms of accuracy across various privacy budgets.
-In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
-Impact of Different $\sigma$.
-We tested the variations in accuracy and fairness on two datasets when $\sigma$ was set to $[1,1.5,2,2.5,3]$, under the condition of $\epsilon=2$ and $q=0.05$, which supports $T=[6,115,268,463,708]$ rounds of communication.
-As shown in Figure 5, an increase in $\sigma$ allows more training epochs to advance the model’s convergence process, but when $\sigma=3$, excessive noise hinders the convergence.
-When $\sigma$ is smaller, the model performance is relatively similar, with a lower $\Psi$ value.
-As $\sigma$ gradually increases, the performance differences between models become apparent, and the $\Psi$ value increases.
-9 Conclusion
-In this paper, a FedFair algorithm is initially proposed to effectively address fairness concerns by optimizing a novel local loss function.
-The FedFDP, building upon FedFair, is introduced to incorporate a fairness-aware gradient clipping strategy and an adaptive clipping method for additional loss values, thereby achieving both fairness and differential privacy protection.
-Then, we find an optimal fairness parameter $\lambda^{*}$ through convergence analysis and numerical analysis methods, striking a balance between model performance and fairness.
-Subsequently, a comprehensive privacy analysis of the approach is conducted using RDP.
-Through extensive experiments, the results indicate that FedFair and FedFDP significantly outperform state-of-the-art solutions in terms of model performance and fairness.
-It is believed that our work contributes a valuable FL framework for addressing fairness and privacy challenges.
-Appendix A Appendix
-In the beginning, Table 4 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
+II Related Work
+II-A Differentially Private Federated Learning
+Differentially Private Federated Learning (DPFL) has become widely used to enhance privacy protection in federated learning.
+Enayat et al.
+[45] dynamically adjusted compression rates in response to training loss changes to minimize communication rounds.
+Zhou et al.
+[52] examined the impact of local iterations when using the Laplace mechanism for perturbation.
+Fu et al.
+[13] introduced adaptive noise addition, customizing it based on loss fluctuations across clients and communication rounds.
+Ling et al.
+[31] balanced communication resources and privacy budgets by identifying the locally optimal number of iterations through convergence analysis.
+Raouf et al.
+[20] improved model accuracy by updating only a fixed subset of model weights while keeping others unchanged.
+II-B Fairness in DPFL
+The above work in DPFL mainly focuses on model performance and privacy budget efficiency, without addressing fairness issues.
+In machine learning, a range of fairness concepts has been proposed, each emphasizing different aspects to meet diverse stakeholder needs.
+For instance, Good-Intent Fairness [38] and Selection Fairness [51] have been proposed.
+For a deeper understanding of these concepts, please refer to the relevant literature [7, 32, 34, 41].
+In contrast, some recent studies have begun exploring group fairness within DPFL, aiming to reduce biases related to sensitive attributes such as gender and race [15, 39, 16, 17, 40].
+Research suggests that adjusting parameters like clipping thresholds and noise multipliers can enhance group fairness in DPFL algorithms [16].
+Under the framework of DP, adjacent datasets are defined as datasets $D$ and $D^{\prime}$ that differ by a single element, ensuring that an adversary cannot determine whether the output originates from $D$ or $D^{\prime}$.
+This mechanism significantly enhances the protection of attribute privacy and effectively reduces potential discrimination risks.
+Jagielski et al.
+[18] pioneered research on fair learning under DP constraints, addressing the need for predictive models to avoid discrimination based on specific attributes.
+They designed two learning algorithms that ensure both DP and equal opportunity, addressing challenges related to the collection of sensitive attributes.
+Tran et al.
+[44] utilized DP and Lagrangian duality to design neural networks capable of addressing fairness constraints while protecting the privacy of sensitive attributes.
+This approach ensures compliance with legal and ethical neutrality requirements when handling data related to gender, race, or age.
+Similarly, Ding et al.
+[9] proposed a classification model that combines a functional mechanism with decision boundary fairness to achieve both fairness and DP.
+This approach addresses privacy concerns related to training data and mitigates unfair practices stemming from attributes such as gender and race.
+II-C Balanced Performance Fairness
+Unlike the studies mentioned above, our research shifts focus to a relatively unexplored area within DPFL—balanced performance fairness.
+While group fairness emphasizes reducing performance biases related to sensitive attributes across different data types and is primarily used in centralized machine learning, balanced performance fairness, introduced by Li et al.
+[28], is more suited to federated learning.
+It emphasizes balancing the performance of global model across all clients, which is especially meaningful for data providers like companies and institutions.
+Building on the work of [28], Zhao et al.
+[50] proposed a strategy that replaces loss amplification with direct weight reallocation.
+Meanwhile, Li et al.
+[26] concentrated on improving fairness in personalized federated learning for local models.
+Incorporating DP into model updates, with parameter adjustments such as clipping bounds and noise multipliers, can influence group fairness to some extent [16].
+However, achieving balanced performance fairness necessitates adjustments to the direction of model updates.
+Therefore, determining how to update the model under DP constraints to ensure balanced performance fairness, while accounting for the effects of clipping and noise addition, presents a significant challenge.
+To the best of our knowledge, this is the first study to address balanced performance fairness within the context of DPFL.
+III Preliminaries
+In the beginning, Table I presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
+TABLE I: Summary of main notations
 $F(\mathbf{w}$) Global loss function
@@
 |||||$\cdot$||| | ⋅ | | $L_{2}$-norm
-Table 4: Summary of main notations
-A.1 Extra Experiment
-We implement our experiment using PyTorch-1.8 and run all experiments on a server with one Intel i9 13900ks CPU (24 cores), 64GB of memory, and one NVIDIA 4090 GPU, running Windows 10.
-Datasets MNIST FashionMNIST CIFAR10
-Accuracy Fairness Accuracy Fairness Accuracy Fairness
-FedAvg 98.67$\pm$0.09 7.3e-3$\pm$5.4e-3 87.05$\pm$1.02 3.6e-1$\pm$2.2e-1 62.03$\pm$1.00 1.1e0$\pm$8.9e-1
-SCAFFLOD 98.45$\pm$0.03 3.0e-2$\pm$2.6e-2 86.77$\pm$0.25 6.1e0$\pm$4.0e0 62.21$\pm$1.37 9.9e0$\pm$1.1e1
-FedProx 98.70$\pm$0.13 7.5e-3$\pm$5.1e-3 87.30$\pm$0.51 3.5e-1$\pm$2.1e-1 62.13$\pm$0.20 1.2e0$\pm$3.1e-1
-FedDyn 98.40$\pm$0.23 8.0e-3$\pm$2.1e-3 87.48$\pm$0.31 6.2e-1$\pm$3.5e-1 62.33$\pm$1.15 3.2e0$\pm$8.8e-1
-q-FFL 98.72$\pm$0.03 4.7e-3$\pm$1.0e-3 87.35$\pm$0.13 5.1e-1$\pm$5.2e-2 63.33$\pm$0.21 9.4e-1$\pm$3.3e-2
-FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
-Table 5: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
-Herterogeneity Scalability
-Datasets MNIST FashionMNIST CIFAR10
-Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
-FedAvg 98.67 |||| 7.3e-3 98.37 |||| 1.2e-3 89.41 |||| 3.6e-2 62.03 |||| 1.1e0 60.89 |||| 1.5e0 59.78 |||| 1.5e0
-SCAFFLOD 98.45 |||| 5.0e-3 98.51 |||| 2.1e-3 86.39 |||| 3.4e-2 62.21 |||| 9.9e0 59.48 |||| 8.7e0 58.64 |||| 2.3e0
-FedProx 98.70 |||| 7.5-3 98.35 |||| 1.2e-3 88.33 |||| 3.5e-2 62.13 |||| 1.2e0 60.33 |||| 2.2e0 61.31 |||| 1.6e0
-FedDyn 98.40 |||| 8.0e-3 98.36 |||| 1.8e-3 88.26 |||| 3.5e-2 62.33 |||| 9.4e-1 61.35 |||| 2.6e0 62.52 |||| 1.5e0
-q-FFL 98.72 |||| 4.7e-3 98.61 |||| 1.1e-3 89.03 |||| 3.6e-2 63.33 |||| 9.4e-1 60.63 |||| 1.3e0 59.65 |||| 1.5e0
-FedFair 98.75 |||| 3.9e-3 98.72 |||| 1.2e-3 89.60 |||| 1.7e-2 62.38 |||| 8.5e-1 61.58 |||| 1.1e0 62.90 |||| 1.5e0
-Table 6: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
-Default Settings
-We conducted experiments on three computer vision (CV) datasets: MNIST (LeCun et al.
-1998), FashionMNIST (Xiao, Rasul, and Vollgraf 2017), and CIFAR10 (Krizhevsky 2009).
-And we comparised our algorithms to six baselines:
-  1. 1.
-    FedAvg(McMahan et al. 2017): FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
-  2. 2.
-    SCAFFLOD(Karimireddy et al. 2020): SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
-  3. 3.
-    FedProx(Li et al. 2020a): FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
-  4. 4.
-    FedDyn(Acar et al.
-2021): FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
-  5. 5.
-    ALI-DPFL(Ling et al.
-2024): ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
-  6. 6.
-    q-FFL(Li et al.
-2020b): q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
-For all datasets, we adopted a widely used heterogeneous settings (Lin et al.
-2020; Li, He, and Song 2021) to 10 clients which controlled by a Dirichlet distribution denoted as $\mathrm{Dir}(\beta)$, and the default value of $\beta=0.1$ (Lin et al.
-2020; Wang et al.
-2020).
-We use a 4-layer CNN architecture (McMahan et al.
-2017) which consists of two convolutional layers and two fully connected layers as model architecture.
-In hyper-parameters, for FedFair and its baseline algorithms, we set learning rate $\eta=0.1$.
-For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$.
-In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
-Extended description of FedFair
+III-A Federated Learning and Fairness
+We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model.
+The objective of FL is to minimize $F(\mathbf{\cdot})$, in other words, to seek:
+$\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (1)
+Here, $F_{i}(\cdot)$ represents the local loss function of client $i$.
+The weights are defined as $p_{i}=\frac{|D_{i}|}{|D|}$, $\sum_{i=1}^{N}p_{i}=1$, $D_{i}$ represents the dataset of client $i$, and $D$ represents the union of all $D_{i}$.
+We have defined balanced performance fairness in FL as follows.
+Definition 1
+(Balanced performance fairness [28]) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
+In this study, we employ the loss of model denote the performance.
+Equation (2) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients.
+Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
+$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (2)
+III-B Differential Privacy
+Differential privacy is a formal mathematical model that quantifies privacy leakage in data analysis algorithms.
+It stipulates that alterations to a single record within the training data should not induce significant statistical variations in the algorithm’s results.
+Definition 2
+(Differential Privacy [11]).
+$(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
+$\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$ (3)
+By adding random noise, we can achieve differential privacy for a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ according to Definition 2.
+The $l_{k}$-sensitivity determines how much noise is needed and is defined as follow.
+We use $l_{2}$ sensitivity as in the previous work.
+DPSGD. In gradient-based methods, DPSGD[1] is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
+  • Firstly, computing the gradient for data sample $\xi_{j}$: $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
+  • Secondly, clipping the gradient of per-sample to get the sensitivity:
+    $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right),$ (4)
+    where $C$ represents the clipping bound.
+  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier:
+    $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$ (5)
+The final step is to proceed with gradient descent as usual.
+When each client in FL utilizes DPSGD as the optimization algorithm to safeguard their private data, the framework transitions from FL to DPFL.
+IV FedFair: Equipping FL with Fairness
 Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, fairness hyperparameter $\lambda$, batch sample ratio $q$, local dataset $D_{i}$.
@@
 Algorithm 1 FedFair
-  • Server (lines 2-2): The server executes $T$ rounds of communication, each round sending the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
-  • Local (lines 2-2): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 2-2. For line 2, we execute gradient descent by using $\eta_{i}$.
-Result of FedFair
-Furthermore, as the training process advances, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$.
-In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods (Zinkevich 2003).
-Performance and Fairness.
-Table 5 presents the experimental results of the FedFair algorithm compared to baseline methods.
-FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
-Heterogeneity.
-Table 6 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity.
-Except for the $\mathrm{Dir}(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
-Scalability. Table 6 indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
-A.2 Proof of Theorem 1
-For the purpose of validation, we introduce an additional variable $\mathbf{v}_{t}^{i}$ to represent the immediate result of a single-step DPSGD update from $\mathbf{w}_{t}^{i}$.
-We interpret $\mathbf{w}_{t+1}^{i}$ as the parameter obtained after a single communication step.
-Consequently, the fair-clipping DPSGD in client $i$ at iteration $t$ transitions from Equation (10) to:
-$\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (17)
+We previously defined the performance objective in Equation (1) and balanced performance fairness in Definition 1.
+Here, we use these to develop a new objective function that informs our proposed fairness-aware federated learning algorithm.
+Upon incorporating Equation (2) into Equation (1), we introduce a comprehensive objective function that integrates considerations of fairness:
+$\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (6)
+For client $i$ during round $t$, the refined objective is articulated as follows:
+$\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$ (7)
+where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training.
+from round $t-1$, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning.
+When $\lambda\rightarrow\infty$, the algorithm only focus on fairness.
+However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging.
+Fortunately, we theoretically found an optimal value for $\lambda$ in Section VI-B.
+In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$..
+Consequently, within the context of Equation (7), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed.
+So, the gradient of Equation (7) as follows:
+$\nabla H_{i}(\mathbf{w}_{t}^{i})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})$
+$=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (8)
+Figure 2: FedFDP framework.
+where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t}^{i})$.
+The stochastic gradient descent by using Equation (IV) as show:
+$\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (9)
+where $\eta_{i}=\eta(1+\lambda\Delta_{i})$.
+Therefore, our approach can be recognized as establishing an dynamic learning rate that allows the model update to achieve both performance and fairness objectives simultaneously.
+Based on these principles, we introduce the FedFair algorithm, as describe in Algorithm 1:
+  • Server (lines 1-1): The server executes $T$ rounds of communication, each round broadcasting the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
+  • Local (lines 1-1): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
+V FedFDP: Further Equipping FedFair with Differential Privacy
+In this section, we present our algorithm, FedFDP, which further integrates FedFair with differential privacy.
+The flow of FedFDP algorithm is shown in Figure 2.
+It adds two additional processes to FedFair:
+  • A fair-clipping strategy is added to FedFair. This strategy not only achieves sensitivity but also allows adjustments to the level of fairness.
+  • An adaptive clipping method is employed when protecting the loss values that clients share with differential privacy, aiming to maximize utility.
+Next, we discuss the details of these two strategies.
+V-A Fair-clipping Strategy for Gradient
+In the general DPSGD algorithm [1, 14, 4, 49], per-sample clipping is an essential process for obtaining sensitivity.
+In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (7), into Equation (10):
+$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$ (10)
+where $\xi_{j}\in\mathcal{B}_{t}^{i}$, $j=\{1,2,\cdots,|\mathcal{B}_{i}|\}$, and $|\mathcal{B}_{i}|$ is the batch size of client $i$.
+Next, we compute the gradient of Equation (10):
+$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})=\left(1+\lambda\cdot\Delta_{i}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$ (11)
+where $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})$.
+To ensure the algorithm is safeguarded by differential privacy, we replace the gradient of the loss function with Equation (11) and reformulate the gradient descent computation for client $i$ as follows:
+$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\eta\tilde{g}^{i}_{t}$
+$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
+$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
+$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (12)
 where:
-$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
-In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by (Stich 2018).
-Therefore,
-$\mathbf{v}_{t+1}$ $=\mathbf{w}_{t}$
-$\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (18)
-Key Lemma
-Lemma 1
-(Results of one iteration.)Assume Assumption 1-3 hold, we have:
-$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq\left(1-\mu\eta C_{t}\right)\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
+$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$ (13)
+Hence, our strategy can be considered as a fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model.
+When $\lambda=0$, Equation (13) corresponds to the traditional DP clipping method without fairness.
+For large $C$, Equation (13) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness.
+Conversely, for small $C$, Equation (13) simplifies to the second term, resulting in the gradient being clipped to the bound $C$.
+The impact of $C$ on the performance of FedFDP will be further discussed in the experiments.
+V-B Adaptive Clipping Method for Loss
+Since calculating $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, differential privacy must be incorporated to ensure the algorithm preserves privacy.
+As shown in lines 2-2 of Algorithm 2, $F_{i}(\mathbf{w}_{t}^{i})$ must be clipped to the interval $[0,C_{l}^{i,t}]$, followed by the addition of Gaussian noise with a mean of 0 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy.
+However, determining an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging.
+A large $C_{l}^{i,t}$ prevents distortion of the original values but results in larger injected noise.
+Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t}^{i})$.
+For the $i$-th client, the adaptive clipping method uses the differentially private mean of the previous round as the clipping bound for the current round.
+The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1}^{i})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
+$C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$ (14)
+where
+$\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))).$ (15)
+Input: loss function $F(\mathbf{w})$, learning rate $\eta$, noise multiplier for gradient $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
+Output: the final trained model $\mathbf{w}_{T}$
+1 SERVER
+2       for $t\in[0,1,\cdots,T-1]$  do
+3             for $i\in[1,2,\cdots,N]$ parallel  do
+4                   $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
+5
+6            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
+7             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
+8
+9      return $\mathbf{w}_{T}$;
+10
+11LOCAL
+12       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
+13       Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
+14       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
+15             $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
+16             Compute $C^{i,j}_{t}$ by Equation (13);
+17             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
+18
+19      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right)$;
+20
+21      Compute clipping bound of loss $C_{l}^{i,t}$ by Equation (14) and (15) ;
+22       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
+23             Compute $F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})$;
+24
+25            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
+26
+27      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\mathcal{N}(0,\mathbf{I})\right)$;
+28       return $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
+29
+Algorithm 2 FedFDP
+To ensure differential privacy during the clipping process, we use the noise-added loss $C_{l}^{i,t}$ as the clipping bound.
+For each communication round, the clipping bound follows the post-processing property of differential privacy.
+V-C Overall Algorithm
+We utilize the equation form Equation (12) to enhance Algorithm 2. Compared with the Algorithm 1, FedFDP solely necessitates the application of differential privacy to gradient and loss.
+  • DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$.
+Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 2 and execute the gradient decent.
+  • DP for local loss (lines 2-2): As described in Section V-B, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
+VI Finding Optimal Fairness Parameter
+In this section, we discuss how to identify an optimal fairness parameter $\lambda$ through the convergence analysis of the FedFDP algorithm.
+VI-A Convergence Analysis
+We analyze the convergence of Algorithm 2 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound.
+Prior to proving this, we need to establish several assumptions:
+Assumption 1
+$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
+Assumption 2
+$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
+Assumption 3
+Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random.
+The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i\in[1,\cdots,N]$ , $t\in[0,\cdots,T-1]$ and $j\in[1,\cdots,|\mathcal{B}_{i}|]$
+Theorem 1
+Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 2 as follows:
+$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (16)
 where:
@@
   • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
-Let $\Delta_{t}\triangleq\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. It is evident that we always have $\mathbf{w}_{t+1}=\mathbf{v}_{t+1}$. According to Lemma 1, this implies:
-$\Delta_{t+1}\leq\left(1-\mu\eta C_{t}\right)\Delta_{t}+\eta^{2}A$
-We use mathematical induction to obtain $\Delta_{t}\leq\frac{v}{t}$ where $v=\max\{\frac{\beta^{2}A}{\mu\beta C_{t}-1},\Delta_{1}\}$, $\eta=\frac{\beta}{t}$ for some $\beta>\frac{1}{\mu}$.
-STEP 1. When $t=1$, the equation $\Delta_{1}\leq v$ holds obviously.
-STEP 2. We assume $\Delta_{t}\leq\frac{v}{t}$ holds.
-STEP 3.
-$\Delta_{t+1}$ $\leq\left(1-\mu\frac{\beta}{t}C_{t}\right)\frac{v}{t}+\frac{\beta^{2}A}{t^{2}}$
-$=\frac{t-1}{t^{2}}v+\left(\frac{\beta^{2}A}{t^{2}}-\frac{\mu\beta C_{t}-1}{t^{2}}v\right)$
-$\leq\frac{t-1}{t^{2}}v$
-$\leq\frac{v}{t+1}$
-Therefore, $\Delta_{t+1}\leq\frac{v}{t+1}$ holds, completing the proof by mathematical induction. Hence, $\Delta_{t}\leq\frac{v}{t}$ holds.
-Then by the $L$-smoothness of $F(\cdot)$, let $\beta=\frac{2}{\mu}$, we get
-$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2}\Delta_{t}\leq\frac{L}{2t}v\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\Delta_{1}\right)$
-Proof of Lemma 1
-By the Equation (A.2), we get
-$\left\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\right\|^{2}$
-$=\|\mathbf{w}_{t}-\mathbf{w}^{*}$
-$-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\|^{2}$
-$=\left\|\mathbf{w}_{t}-\mathbf{w}^{*}\right\|^{2}$
-$\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\mathcal{N}(0,\mathbf{I})]}_{\mathcal{A}_{1}}\rangle$
-$+\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}}_{\mathcal{A}_{2}}$
-Firstly, we process $\mathcal{A}_{2}$:
-$\mathcal{A}_{2}\leq\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2}}_{\mathcal{B}_{1}}$
-$\quad+\underbrace{\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\left\langle\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right\rangle}_{\mathcal{B}_{0}}$
-$\quad+\underbrace{\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\cdot\mathcal{N}(0,\mathbf{I})\|^{2}}_{\mathcal{B}_{2}}$
-Since $\mathbb{E}[\mathcal{B}_{0}]=0$, we focus on $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$:
-$\mathbb{E}[\mathcal{B}_{2}]\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
-where $\frac{1}{\hat{B}^{2}}\triangleq\max_{i}\frac{1}{|\mathcal{B}_{i}|}$.
-By the convexity of $\|\cdot\|^{2}$,
-$\mathcal{B}_{1}\leq\eta^{2}\sum_{i=1}^{N}p_{i}\left\|\frac{1}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2},$
-taking exception and according to assumption 3:
-$\mathbb{E}[\mathcal{B}_{1}]\leq\eta^{2}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\right]\leq\eta^{2}C_{t}^{2}G^{2},$
-where $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i}$, $C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
-Now, we obtain the bound for the expectation of $\mathcal{A}_{2}$:
-$\mathbb{E}[\mathcal{A}_{2}]$ $\leq\eta^{2}\left(\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}+G^{2}C_{t}^{2}\right)$
-The process of $\mathcal{A}_{1}$ show as below:
-$\mathcal{A}_{1}$ $=-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
-$\quad\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\mathcal{N}(0,\mathbf{I})\rangle}_{\mathcal{C}_{0}}$
-$=\mathcal{C}_{0}$
-$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{1}}$
-$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{2}}$
-It’s obvious that $\mathbb{E}[\mathcal{C}_{0}]=0$.
-By Cauchy-Schwarz inequality and AM-GM inequality, we have
-$\mathcal{C}_{1}$ $=-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
-$\leq\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|^{2}$
-$=\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\mathcal{B}_{1}$
-So we get
-$\mathbb{E}[\mathcal{C}_{1}]\leq\sum_{i=1}^{N}p_{i}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}$
-According to Assumption 2, we know that
-$-\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\nabla F_{i}(\mathbf{w}_{t}^{i})\rangle\leq-\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
-So we get
-$\mathcal{C}_{2}$ $\leq 2\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}\cdot$
-$\quad\left[-\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}\right]$
-$\leq-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
-$=-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}+F^{*}-F_{i}(\mathbf{w}^{*})\right)$
-$\quad-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
-$=-2\eta C_{t}\underbrace{\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}\right)}_{\mathcal{D}_{1}}-2\eta C_{t}\Gamma-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2},$
-where $\Gamma\triangleq\sum_{i=1}^{N}p_{i}(F^{*}-F_{i}^{*})=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$.
-Next, we proceed to handle $\mathcal{D}_{1}$.
-$\mathcal{D}_{1}$ $=\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}_{t})\right)+\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t})-F^{*}\right)$
-$\geq\sum_{i=1}^{N}p_{i}\langle\nabla F_{i}(\mathbf{w}_{t}),\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\rangle+(F(\mathbf{w}_{t})-F^{*})$
-(from the Assumption 2)
-$\geq-\frac{1}{2}\sum_{i=1}^{N}p_{i}\left[\eta\|\nabla F_{i}(\mathbf{w}_{t})\|^{2}+\frac{1}{\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
-$\quad+(F(\mathbf{w}_{t})-F^{*})$
-(from the AM-GM inequality)
-$\geq-\sum_{i=1}^{N}p_{i}\left[\eta L(F_{i}(\mathbf{w}_{t})-F_{i}^{*})+\frac{1}{2\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
-$\quad+(F(\mathbf{w}_{t})-F^{*})$
-(from the L-smooth inference)
-$\geq-(\eta L+1)\Gamma-\frac{1}{2\eta}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2},$
-where L-smooth inference as show:
-$\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$ (19)
-Thus, we get
-$\mathcal{C}_{2}\leq 2\eta^{2}C_{t}L\Gamma+C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
-To sum up,
-$\mathbb{E}[\mathcal{A}_{1}]$ $=\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}+2\eta^{2}C_{t}L\Gamma$
-$\quad+C_{t}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
-$=(1+C_{t})\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
-$\quad+\eta^{2}\left(G^{2}C_{t}^{2}+2L\Gamma C_{t}\right),$
-and
-$\quad\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]$
-$=\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|(\mathbf{w}_{t}-\mathbf{v}_{t+1}^{i})-(\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i})\|^{2}\right]$
-$\leq\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i}\|^{2}\right]$
-$\leq\mathbb{E}\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}$
-$=\mathbb{E}\left[\mathcal{A}_{2}\right]=\eta^{2}C_{t}^{2}G^{2}+\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
-So we get,
-$\mathbb{E}\left[\mathcal{A}_{1}\right]$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}$
-$\quad+\eta^{2}\left(2G^{2}C_{t}^{2}+2L\Gamma C_{t}+2\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}\right)$
-All in all, we get
-$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
-where
-$A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
-A.3 Proof of Theorem 2
-When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},x_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{i}^{j}$.
+The meaning of $\Gamma$ is consistent with [29], where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
+proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$.
+Finally, by utilizing Assumption 1, we derive Equation (16).
+For the detailed proof, please refer to Appendix A-A.
+VI-B Optimal Fairness Parameter $\lambda^{*}$
 Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$.
@@
 […]
-where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{1,2,3}=x_{1,2,3}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$.
+where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{l}=x_{l}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, $l=1,2,3$.
 Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root.
@@
 In conclusion, $\mathcal{F}^{\prime}(\lambda)\geq 0\iff G(\lambda)\geq 0\iff\lambda\geq\lambda^{*}$, therefore $\mathcal{F}(\lambda)$ has a minimum value $\mathcal{F}(\lambda^{*})$, and $P(\lambda)$ has a minimum value $P(\lambda^{*})$.
-A.4 Proof of Theorem 3
-As Algorithm 1 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$.
-$\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of local client, so they are both to perform differential privacy preservation, as in lines 1-1 and lines 1-1 of the Algorithm 1, respectively.
+VII Privacy Analysis
+Since each client incurs the same privacy loss in every round, we evaluate privacy loss from the perspective of an individual client.
+Accordingly, we analyze the privacy loss of FedPDF for the $i$-th client as follows.
+Theorem 2
+(Privacy loss of FedFDP). The privacy loss in $i$-th client of FedFDP satisfies:
+[…]
+where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 3, and $R_{loss}^{i}(\alpha)$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 4.
+VII-A Proof of Theorem 2
+Lemma 1
+(Conversion from RDP to DP [3]).
+if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
+As Algorithm 2 shown, the $i$-th client uploads model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$.
+Both $\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the private training data of the client, and thus both require differential privacy protection, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively.
 We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
-Privacy Loss of $\mathbf{w}_{t+1}^{i}$.
+VII-A1 Privacy Loss of $\mathbf{w}_{t+1}^{i}$
 Theorem 3
@@
 Definition 3
-(RDP privacy budget of SGM(Mironov, Talwar, and Zhang 2019)).
-Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$.
-If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
+(RDP privacy budget of SGM[37]). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
 $R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$ (25)
@@
 Further, it holds $\forall(q,\sigma)\in(0,1],\mathbb{R}^{+*},A_{\alpha}(q,\sigma)\geq B_{\alpha}(q,\sigma)$.
-Thus, $SG_{q,\sigma}\text{ satisfies }\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
-Finally, the existing work (Mironov, Talwar, and Zhang 2019) describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
+Thus, $SG_{q,\sigma}$ satisfies $\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
+Finally, the existing work [37] describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
 $A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$ (27)
 Definition 4
-(Composition of RDP(Mironov 2017)).
+(Composition of RDP[36]).
 For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
 From Definition 3 and Definition 4, the Theorem 3 is obtained.
-Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$.
+VII-A2 Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$.
 Theorem 4
@@
 The proof is similar to that of the $w_{t+1}^{i}$, so we omit it.
-Proof Loss of FedFDP.
+VII-A3 Proof Loss of FedFDP.
 Since both $w_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 4, and then use Lemma 1 to convert it to $(\epsilon,\delta)$-DP.
 Therefore, the Theorem 2 is hold.
+VIII Experiments
+TABLE II: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
+Datasets MNIST FashionMNIST CIFAR10
+Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
+FedAvg 98.67$\pm$0.09 7.3e-3$\pm$5.4e-3 87.05$\pm$1.02 3.6e-1$\pm$2.2e-1 62.03$\pm$1.00 1.1e0$\pm$8.9e-1
+SCAFFLOD 98.45$\pm$0.03 3.0e-2$\pm$2.6e-2 86.77$\pm$0.25 6.1e0$\pm$4.0e0 62.21$\pm$1.37 9.9e0$\pm$1.1e1
+FedProx 98.70$\pm$0.13 7.5e-3$\pm$5.1e-3 87.30$\pm$0.51 3.5e-1$\pm$2.1e-1 62.13$\pm$0.20 1.2e0$\pm$3.1e-1
+FedDyn 98.40$\pm$0.23 8.0e-3$\pm$2.1e-3 87.48$\pm$0.31 6.2e-1$\pm$3.5e-1 62.33$\pm$1.15 3.2e0$\pm$8.8e-1
+q-FFL 98.72$\pm$0.03 4.7e-3$\pm$1.0e-3 87.35$\pm$0.13 5.1e-1$\pm$5.2e-2 63.33$\pm$0.21 9.4e-1$\pm$3.3e-2
+FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
+TABLE III: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
+Heterogeneity Scalability
+Datasets MNIST FashionMNIST CIFAR10
+Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
+Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
+FedAvg 98.67 7.3e-3 98.37 1.2e-3 89.41 3.6e-2 62.03 1.1e0 60.89 1.5e0 59.78 1.5e0
+SCAFFLOD 98.45 5.0e-3 98.51 2.1e-3 86.39 3.4e-2 62.21 9.9e0 59.48 8.7e0 58.64 2.3e0
+FedProx 98.70 7.5e-3 98.35 1.2e-3 88.33 3.5e-2 62.13 1.2e0 60.33 2.2e0 61.31 1.6e0
+FedDyn 98.40 8.0e-3 98.36 1.8e-3 88.26 3.5e-2 62.33 9.4e-1 61.35 2.6e0 62.52 1.5e0
+q-FFL 98.72 4.7e-3 98.61 1.1e-3 89.03 3.6e-2 63.33 9.4e-1 60.63 1.3e0 59.65 1.5e0
+FedFair 98.75 3.9e-3 98.72 1.2e-3 89.60 1.7e-2 62.38 8.5e-1 61.58 1.1e0 62.90 1.5e0
+In this section, we present the comparative experimental results of FedFair and FedFDP with multiple baseline methods on three datasets: MNIST [22], FashionMNIST [48], and CIFAR10 [21].
+We provide a detailed description of the experimental evaluation of FedFair and FedFDP in terms of performance, fairness, heterogeneity, and scalability.
+In particular, we conduct ablation studies on the fairness hyperparameter $\lambda$, privacy budget $\epsilon$, clipping bound $C$, and noise multiplier $\sigma$.
+In all tables, $\uparrow$ indicates that larger values are better, while $\downarrow$ indicates that smaller values are better.
+We implement our experiment using PyTorch-1.8 and run all experiments on a server with one Intel i9 13900ks CPU (24 cores), 64GB of memory, and one NVIDIA 4090 GPU, running Windows 10.
+VIII-A Default Settings
+We compared our algorithms against six baseline methods. For FedFair, the baseline algorithms were left unmodified; for FedFDP, DP was added to each baseline algorithm to enable a fair comparison.
+  1. 1.
+    FedAvg[33]: FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
+  2. 2.
+    SCAFFLOD[19]: SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
+  3. 3.
+    FedProx[27]: FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
+  4. 4.
+    FedDyn[2]: FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
+  5. 5.
+    ALI-DPFL[31]: ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
+  6. 6.
+    q-FFL[28]: q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
+For all datasets, we adopted a widely used heterogeneous settings [30, 24] to 10 clients which controlled by a Dirichlet distribution denoted as $\mathrm{Dir}(\beta)$, and the default value of $\beta=0.1$ [30, 46].
+We use a 4-layer CNN architecture [33] which consists of two convolutional layers and two fully connected layers as model architecture.
+Hyperparameters.
+For FedFair and its baseline algorithms, we set learning rate $\eta=0.1$.
+For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$.
+In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
+TABLE IV: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
+Datasets MNIST FashionMNIST CIFAR10
+Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
+FedAvg 93.40$\pm$0.47 1.1e11$\pm$0.5e10 84.15$\pm$1.97 3.0e8$\pm$1.4e8 52.61$\pm$0.17 6.1e9$\pm$5.2e9
+SCAFFLOD 93.95$\pm$1.39 7.0e10$\pm$4.9e9 83.41$\pm$4.87 3.3e9$\pm$4.7e9 53.85$\pm$1.16 6.2e9$\pm$1.3e9
+FedProx 90.50$\pm$3.95 5.3e10$\pm$7.3e9 83.85$\pm$0.98 7.9e8$\pm$4.5e8 51.34$\pm$0.91 4.4e9$\pm$2.1e9
+FedDyn 91.42$\pm$2.25 5.6e10$\pm$9.8e9 84.04$\pm$1.21 8.8e8$\pm$3.2e8 52.14$\pm$1.39 5.5e9$\pm$1.2e9
+ALI-DPFL 90.89$\pm$1.65 3.3e10$\pm$8.3e9 83.65$\pm$1.68 6.6e8$\pm$3.5e8 52.05$\pm$1.26 3.6e9$\pm$1.6e9
+q-FFL 93.74$\pm$1.41 7.8e10$\pm$1.1e11 83.13$\pm$2.74 4.2e9$\pm$2.6e9 48.46$\pm$1.00 4.7e9$\pm$1.9e9
+FedFDP 95.13$\pm$0.83 2.3e10$\pm$1.0e10 85.99$\pm$0.76 2.8e8$\pm$0.8e8 54.21$\pm$0.98 2.6e9$\pm$1.6e9
+TABLE V: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
+Herterogeneity Scalability
+Datasets MNIST FashionMNIST CIFAR10
+Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
+Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
+FedAvg 93.40 1.1e11 94.17 4.1e9 85.60 3.7e9 52.61 6.1e9 51.58 2.0e9 55.90 1.5e9
+SCAFFLOD 93.95 7.0e10 94.06 1.9e9 84.56 3.6e9 53.85 6.2e9 54.34 2.9e8 53.98 2.2e9
+FedProx 90.50 5.3e10 94.23 3.1e9 84.38 5.6e9 51.34 4.4e9 53.67 4.5e8 54.14 2.2e9
+FedDyn 91.42 5.6e10 93.08 3.8e9 85.21 9.8e10 52.14 5.5e9 50.21 1.2e10 51.86 2.0e10
+ALI-DPFL 90.89 3.3e10 93.98 2.6e9 83.66 2.5e10 52.05 3.6e9 38.35 2.4e10 33.46 1.6e10
+q-FFL 93.74 7.8e10 93.86 2.5e8 85.20 7.4e8 48.46 4.7e9 51.12 2.7e8 49.83 1.1e7
+FedFDP 94.13 2.3e10 94.56 3.4e9 86.77 6.0e8 54.21 2.6e9 54.52 2.2e8 56.49 5.3e8
+[…]
+$\epsilon$ 1.0 2.0 3.0 4.0
+Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
+FedAvg 61.68 1.1e6 82.69 2.3e7 84.07 3.8e8 86.83 8.0e8
+SCAFFLOD 62.12 1.2e6 82.36 2.3e7 83.39 2.6e9 85.41 5.6e8
+FedProx 61.25 1.1e6 81.62 2.3e7 85.32 3.8e8 85.68 2.4e9
+FedDyn 62.38 1.5e6 82.43 3.1e7 83.29 2.9e8 84.51 5.2e9
+ALI-DPFL 61.78 1.1e6 83.01 2.3e7 84.12 3.9e8 85.58 4.5e9
+q-FFL 58.99 3.6e6 80.24 9.6e7 82.92 5.5e8 85.01 3.8e9
+FedFDP 63.36 1.0e6 83.15 2.1e7 85.59 9.5e7 86.97 3.8e8
+VIII-B Result of FedFair
+Performance and Fairness.
+Table II presents the experimental results of the FedFair algorithm compared to baseline methods.
+FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
+A plausible explanation is that, as the training process progresses, FedFair increasingly prioritizes fairness, resulting in a reduction in $\eta_{i}$.
+Compared to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods [55].
+Figure 3: Impact of different $\lambda$ in FedFDP algorithm.
+Figure 4: Impact of different $C$ in FedFDP algorithm.
+Figure 5: Impact of different $\sigma$ in FedFDP algorithm.
+Heterogeneity.
+Table III shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity.
+Except for the $\mathrm{Dir}(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
+Scalability. Table III indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
+VIII-C Results of FedFDP
+Performance and Fairness.
+Table IV displays the experimental results of the FedFDP algorithm compared to baseline methods.
+Under a consistent privacy budget of $\epsilon=3.52$, when accuracy matches the best baseline, our method significantly enhances fairness, with improvements of 30.3%, 6.7%, and 27.8% observed across three separate datasets.
+Heterogeneity.
+Table V shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity.
+Except for the $\mathrm{Dir}(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
+Scalability.
+Table V demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline.
+Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
+Impact of Different $\lambda$.
+We investigated the correlation between $\lambda$ and $\Psi$ in the FedFDP algorithm.
+As depicted in Figure 3, our findings align with the theoretical analysis: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
+Impact of Different $C$.
+We emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously.
+As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10].
+Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition.
+As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$.
+When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
+Different $\epsilon$.
+We conducted experiments with $\epsilon=[1,2,3,4]$ on the FashionMNIST dataset, as detailed in Table VI.
+For $\epsilon=[1,2,3]$, we allowed larger values of $\epsilon$ to support more iterations.
+At $\epsilon=4$, we supported a smaller noise multiplier $\sigma=1.65$ with a similar number of iterations as at $\epsilon=3$.
+The experimental results demonstrate that FedFDP consistently achieves performance on par with the best baseline in terms of accuracy across various privacy budgets.
+In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
+Impact of Different $\sigma$.
+We tested the variations in accuracy and fairness on two datasets when $\sigma$ was set to $[1,1.5,2,2.5,3]$, under the condition of $\epsilon=2$ and $q=0.05$, which supports $T=[6,115,268,463,708]$ rounds of communication.
+As shown in Figure 5, an increase in $\sigma$ allows more training epochs to advance the model’s convergence process, but when $\sigma=3$, excessive noise hinders the convergence.
+When $\sigma$ is smaller, the model performance is relatively similar, with a lower $\Psi$ value.
+As $\sigma$ gradually increases, the performance differences between models become apparent, and the $\Psi$ value increases.
+IX Conclusion
+In this paper, a FedFair algorithm is initially proposed to effectively address fairness concerns by optimizing a novel local loss function.
+The FedFDP, building upon FedFair, is introduced to incorporate a fairness-aware gradient clipping strategy and an adaptive clipping method for additional loss values, thereby achieving both fairness and differential privacy protection.
+Then, we find an optimal fairness parameter $\lambda^{*}$ through convergence analysis and numerical analysis methods, striking a balance between model performance and fairness.
+Subsequently, a comprehensive privacy analysis of the approach is conducted using RDP.
+Through extensive experiments, the results indicate that FedFair and FedFDP significantly outperform state-of-the-art solutions in terms of model performance and fairness.
+It is believed that our work contributes a valuable FL framework for addressing fairness and privacy challenges.
+Appendix A Appendix
+A-A Proof of Theorem 1
+For the purpose of validation, we introduce an additional variable $\mathbf{v}_{t}^{i}$ to represent the immediate result of a single-step DPSGD update from $\mathbf{w}_{t}^{i}$.
+We interpret $\mathbf{w}_{t+1}^{i}$ as the parameter obtained after a single communication step.
+Consequently, the fair-clipping DPSGD in client $i$ at iteration $t$ transitions from Equation (12) to:
+$\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (29)
+where:
+$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
+In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by [43].
+Therefore,
+$\mathbf{v}_{t+1}=\mathbf{w}_{t}\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (30)
+A-A1 Key Lemma
 Lemma 2
-(Conversion from RDP to DP (Balle et al.
-2020)).
-if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
+(Results of one iteration.)Assume Assumption 1-3 hold, we have:
+$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq\left(1-\mu\eta C_{t}\right)\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
+where:
+  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
+  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
+  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
+  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
+Let $\Delta_{t}\triangleq\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. It is evident that we always have $\mathbf{w}_{t+1}=\mathbf{v}_{t+1}$. According to Lemma 2, this implies:
+$\Delta_{t+1}\leq\left(1-\mu\eta C_{t}\right)\Delta_{t}+\eta^{2}A$
+We use mathematical induction to obtain $\Delta_{t}\leq\frac{v}{t}$ where $v=\max\{\frac{\beta^{2}A}{\mu\beta C_{t}-1},\Delta_{1}\}$, $\eta=\frac{\beta}{t}$ for some $\beta>\frac{1}{\mu}$.
+STEP 1. When $t=1$, the equation $\Delta_{1}\leq v$ holds obviously.
+STEP 2. We assume $\Delta_{t}\leq\frac{v}{t}$ holds.
+STEP 3.
+$\Delta_{t+1}\leq\left(1-\mu\frac{\beta}{t}C_{t}\right)\frac{v}{t}+\frac{\beta^{2}A}{t^{2}}=\frac{t-1}{t^{2}}v+\left(\frac{\beta^{2}A}{t^{2}}-\frac{\mu\beta C_{t}-1}{t^{2}}v\right)\leq\frac{t-1}{t^{2}}v\leq\frac{v}{t+1}$
+Therefore, $\Delta_{t+1}\leq\frac{v}{t+1}$ holds, completing the proof by mathematical induction. Hence, $\Delta_{t}\leq\frac{v}{t}$ holds.
+Then by the $L$-smoothness of $F(\cdot)$, let $\beta=\frac{2}{\mu}$, we get
+$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2}\Delta_{t}\leq\frac{L}{2t}v\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\Delta_{1}\right)$
+A-A2 Proof of Lemma 2
+By the Equation (30), we get
+$\left\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\right\|^{2}$
+$=\|\mathbf{w}_{t}-\mathbf{w}^{*}-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\|^{2}$
+$=\left\|\mathbf{w}_{t}-\mathbf{w}^{*}\right\|^{2}$
+$\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\mathcal{N}(0,\mathbf{I})]}_{\mathcal{A}_{1}}\rangle$
+$+\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}}_{\mathcal{A}_{2}}$
+Firstly, we process $\mathcal{A}_{2}$:
+$\mathcal{A}_{2}\leq\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2}}_{\mathcal{B}_{1}}$
+$\quad+\underbrace{\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\left\langle\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right\rangle}_{\mathcal{B}_{0}}$
+$\quad+\underbrace{\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\cdot\mathcal{N}(0,\mathbf{I})\|^{2}}_{\mathcal{B}_{2}}$
+Since $\mathbb{E}[\mathcal{B}_{0}]=0$, we focus on $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$:
+$\mathbb{E}[\mathcal{B}_{2}]\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
+where $\frac{1}{\hat{B}^{2}}\triangleq\max_{i}\frac{1}{|\mathcal{B}_{i}|}$.
+By the convexity of $\|\cdot\|^{2}$,
+$\mathcal{B}_{1}\leq\eta^{2}\sum_{i=1}^{N}p_{i}\left\|\frac{1}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2},$
+taking exception and according to assumption 3:
+$\mathbb{E}[\mathcal{B}_{1}]\leq\eta^{2}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\right]\leq\eta^{2}C_{t}^{2}G^{2},$
+where $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i}$, $C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
+Now, we obtain the bound for the expectation of $\mathcal{A}_{2}$:
+$\mathbb{E}[\mathcal{A}_{2}]$ $\leq\eta^{2}\left(\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}+G^{2}C_{t}^{2}\right)$
+The process of $\mathcal{A}_{1}$ show as below:
+$\mathcal{A}_{1}$ $=-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\mathcal{N}(0,\mathbf{I})\rangle}_{\mathcal{C}_{0}}$
+$=\mathcal{C}_{0}\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{1}}\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{2}}$
+It’s obvious that $\mathbb{E}[\mathcal{C}_{0}]=0$.
+By Cauchy-Schwarz inequality and AM-GM inequality, we have
+$\mathcal{C}_{1}$ $=-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
+$\leq\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|^{2}$
+$=\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\mathcal{B}_{1}$
+So we get
+$\mathbb{E}[\mathcal{C}_{1}]\leq\sum_{i=1}^{N}p_{i}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}$
+According to Assumption 2, we know that
+$-\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\nabla F_{i}(\mathbf{w}_{t}^{i})\rangle\leq-\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
+So we get
+$\mathcal{C}_{2}$ $\leq 2\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}\cdot\left[-\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}\right]$
+$\leq-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
+$=-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}+F^{*}-F_{i}(\mathbf{w}^{*})\right)-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
+$=-2\eta C_{t}\underbrace{\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}\right)}_{\mathcal{D}_{1}}-2\eta C_{t}\Gamma-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2},$
+where $\Gamma\triangleq\sum_{i=1}^{N}p_{i}(F^{*}-F_{i}^{*})=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$.
+Next, we proceed to handle $\mathcal{D}_{1}$.
+$\mathcal{D}_{1}$ $=\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}_{t})\right)+\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t})-F^{*}\right)$
+$\geq\sum_{i=1}^{N}p_{i}\langle\nabla F_{i}(\mathbf{w}_{t}),\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\rangle+(F(\mathbf{w}_{t})-F^{*})$
+(from the Assumption 2)
+$\geq-\frac{1}{2}\sum_{i=1}^{N}p_{i}\left[\eta\|\nabla F_{i}(\mathbf{w}_{t})\|^{2}+\frac{1}{\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]+(F(\mathbf{w}_{t})-F^{*})$
+(from the AM-GM inequality)
+$\geq-\sum_{i=1}^{N}p_{i}\left[\eta L(F_{i}(\mathbf{w}_{t})-F_{i}^{*})+\frac{1}{2\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]+(F(\mathbf{w}_{t})-F^{*})$
+(from the L-smooth inference)
+$\geq-(\eta L+1)\Gamma-\frac{1}{2\eta}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2},$
+where L-smooth inference as show:
+$\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$ (31)
+Thus, we get
+$\mathcal{C}_{2}\leq 2\eta^{2}C_{t}L\Gamma+C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
+To sum up,
+$\mathbb{E}[\mathcal{A}_{1}]$ $=\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}+2\eta^{2}C_{t}L\Gamma$
+$\quad+C_{t}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
+$=(1+C_{t})\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
+$\quad+\eta^{2}\left(G^{2}C_{t}^{2}+2L\Gamma C_{t}\right),$
+and
+$\quad\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]$
+$=\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|(\mathbf{w}_{t}-\mathbf{v}_{t+1}^{i})-(\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i})\|^{2}\right]$
+$\leq\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i}\|^{2}\right]$
+$\leq\mathbb{E}\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}$
+$=\mathbb{E}\left[\mathcal{A}_{2}\right]=\eta^{2}C_{t}^{2}G^{2}+\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
+So we get,
+$\mathbb{E}\left[\mathcal{A}_{1}\right]$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}+\eta^{2}\left(2G^{2}C_{t}^{2}+2L\Gamma C_{t}+2\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}\right)$
+All in all, we get
+$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
+where
+$A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}.$

### v6

@@
-FedFDP: Fairness-Aware Federated Learning with Differential Privacy
+
 Abstract
@@
 Extensive experimental results show that FedFDP significantly surpasses state-of-the-art solutions in both model performance and fairness.
-Index Terms:
-Federated Learning, Differential Privacy, Fairness
-I Introduction
+1 Introduction
 Federated Learning (FL) [30] is a distributed machine learning framework that allows clients to collaboratively train a shared model without exposing their respective datasets.
@@
 Particularly, many studies have achieved significant breakthroughs in terms of model performance [38, 24, 23] and communication costs [39, 28] in FL.
-Figure 1: Client drift caused by Non-IID.
 To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary [22].
 Current fairness-aware machine learning algorithms, such as [10, 8], have not fully addressed the issue of client performance disparities in federated learning.
-As shown in Figure 1, due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (orange squares) lead to significant performance disparities of the collaboratively trained global model (red square) across different datasets held by various clients.
+Due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (orange squares) lead to significant performance disparities of the collaboratively trained global model (red square) across different datasets held by various clients.
 An increasing number of scholars are beginning to focus on this phenomenon of unfair performance disparity [25, 23, 42].
@@
 To enhance data privacy in FL, differential privacy (DP) [11] has become the preferred method for protecting privacy in federated learning due to its moderate computational overhead and solid mathematical foundation.
-Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information [1, 5, 12].
+Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information [12].
 Although some works have explored balanced performance fairness and differential privacy in federated learning separately, no prior research has considered them in a unified framework.
@@
   3. 3.
-    In addition, we conducted a convergence analysis of FedFDP and identified the fairness parameter $\lambda^{*}$ that results in the fastest convergence.
+    Furthermore, we conducted a convergence analysis of FedFDP and identified the fairness parameter $\lambda^{*}$ that results in the fastest convergence.
 We analyzed the privacy loss of the FedFDP algorithm using the concept of Rényi Differential Privacy (RDP) to verify its compliance with differential privacy requirements.
@@
     Finally, through comprehensive empirical evaluations on three public datasets, we confirmed that FedFair and FedFDP are superior to existing solutions in terms of model performance and fairness.
-What’s more, we investigated the impact of clipping bound and noise multiplier in differential privacy on fairness.
-II Related Work
-II-A Differentially Private Federated Learning
-Differentially Private Federated Learning (DPFL) has become widely used to enhance privacy protection in federated learning.
-Enayat et al.
-[45] dynamically adjusted compression rates in response to training loss changes to minimize communication rounds.
-Zhou et al.
-[52] examined the impact of local iterations when using the Laplace mechanism for perturbation.
-Fu et al.
-[13] introduced adaptive noise addition, customizing it based on loss fluctuations across clients and communication rounds.
-Ling et al.
-[31] balanced communication resources and privacy budgets by identifying the locally optimal number of iterations through convergence analysis.
-Raouf et al.
-[20] improved model accuracy by updating only a fixed subset of model weights while keeping others unchanged.
-II-B Fairness in DPFL
-The above work in DPFL mainly focuses on model performance and privacy budget efficiency, without addressing fairness issues.
-In machine learning, a range of fairness concepts has been proposed, each emphasizing different aspects to meet diverse stakeholder needs.
-For instance, Good-Intent Fairness [38] and Selection Fairness [51] have been proposed.
-For a deeper understanding of these concepts, please refer to the relevant literature [7, 32, 34, 41].
-In contrast, some recent studies have begun exploring group fairness within DPFL, aiming to reduce biases related to sensitive attributes such as gender and race [15, 39, 16, 17, 40].
-Research suggests that adjusting parameters like clipping thresholds and noise multipliers can enhance group fairness in DPFL algorithms [16].
-Under the framework of DP, adjacent datasets are defined as datasets $D$ and $D^{\prime}$ that differ by a single element, ensuring that an adversary cannot determine whether the output originates from $D$ or $D^{\prime}$.
-This mechanism significantly enhances the protection of attribute privacy and effectively reduces potential discrimination risks.
-Jagielski et al.
-[18] pioneered research on fair learning under DP constraints, addressing the need for predictive models to avoid discrimination based on specific attributes.
-They designed two learning algorithms that ensure both DP and equal opportunity, addressing challenges related to the collection of sensitive attributes.
-Tran et al.
-[44] utilized DP and Lagrangian duality to design neural networks capable of addressing fairness constraints while protecting the privacy of sensitive attributes.
-This approach ensures compliance with legal and ethical neutrality requirements when handling data related to gender, race, or age.
-Similarly, Ding et al.
-[9] proposed a classification model that combines a functional mechanism with decision boundary fairness to achieve both fairness and DP.
-This approach addresses privacy concerns related to training data and mitigates unfair practices stemming from attributes such as gender and race.
-II-C Balanced Performance Fairness
-Unlike the studies mentioned above, our research shifts focus to a relatively unexplored area within DPFL—balanced performance fairness.
-While group fairness emphasizes reducing performance biases related to sensitive attributes across different data types and is primarily used in centralized machine learning, balanced performance fairness, introduced by Li et al.
-[28], is more suited to federated learning.
-It emphasizes balancing the performance of global model across all clients, which is especially meaningful for data providers like companies and institutions.
-Building on the work of [28], Zhao et al.
-[50] proposed a strategy that replaces loss amplification with direct weight reallocation.
-Meanwhile, Li et al.
-[26] concentrated on improving fairness in personalized federated learning for local models.
-Incorporating DP into model updates, with parameter adjustments such as clipping bounds and noise multipliers, can influence group fairness to some extent [16].
-However, achieving balanced performance fairness necessitates adjustments to the direction of model updates.
-Therefore, determining how to update the model under DP constraints to ensure balanced performance fairness, while accounting for the effects of clipping and noise addition, presents a significant challenge.
-To the best of our knowledge, this is the first study to address balanced performance fairness within the context of DPFL.
-III Preliminaries
-In the beginning, Table I presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
-TABLE I: Summary of main notations
-$F(\mathbf{w}$) Global loss function
-$F_{i}(\mathbf{w}$) Local loss function for client $i$
-$T$ Communication rounds
-$t$ Communication round index
-$D_{i}$ Dataset of client $i$
-$D$ The union of $D_{i}$
-$\mathbf{w}_{t}^{i}$ Local model parameter at node $i$ at round $t$
-$\mathbf{w}^{*}$ Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
-$\eta$ Gradient descent step size
-$\lambda$ Fairness parameter
-$N$ Total number of clients
-$q$ batch sample ratio
-$|\mathcal{B}_{i}|$ Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
-$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
-$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
-$\sigma$ Noise multiplier for gradient
-$\sigma_{l}$ Noise multiplier for loss
-$C$ Original clipping bound for gradient
-$C_{l}$ Original clipping bound for loss
-|||||$\cdot$||| | ⋅ | | $L_{2}$-norm
-III-A Federated Learning and Fairness
+What’s more, we investigated the impact of clipping bound and noise multiplier on fairness.
+In Section 2, we formalize our problem and introduce the three objectives of our method.
+Section 3 presents a fair federated learning approach—FedFair.
+In Section 4, we enhance FedFair by incorporating the DP guarantee, thereby arriving at the proposed FedFDP.
+A fairness parameter $\lambda$ is embedded within FedFDP, and in Section 5, we derive the analytical solution for the optimal $\lambda^{*}$ through convergence analysis and differential methods.
+The privacy analysis of FedFDP are examined in detail in Section 6.
+In Section 7, we conduct extensive experiments on three datasets, comparing six baseline algorithms to rigorously validate the effectiveness of both FedFair and FedFDP.
+Following the review of related work in Section 0.B, Section 9 provides a summary of our contributions.
+Due to space limitations, the detailed procedures of the convergence analysis are included in Appendix 0.A.
+2 Problem Formulation
+In this work, we propose FedFDP which aims to serve threefold goals as follows. Then, we will elaborate on and formalize how fairness and DP are defined in FL.
+  • Goal 1 (Fairness): Strive to achieve federated learning with the highest possible fairness, where lower values of the federated fairness metric (Equation 2) indicate better fairness.
+  • Goal 2 (Privacy): Implement federated learning under formal differential privacy guarantees (Equation 3).
+  • Goal 3 (Utility): Ensure that the federated learning model retains strong predictive performance, particularly in terms of classification accuracy.
+2.1 Fairness in FL
 We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model.
 The objective of FL is to minimize $F(\mathbf{\cdot})$, in other words, to seek:
-$\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (1)
-Here, $F_{i}(\cdot)$ represents the local loss function of client $i$.
-The weights are defined as $p_{i}=\frac{|D_{i}|}{|D|}$, $\sum_{i=1}^{N}p_{i}=1$, $D_{i}$ represents the dataset of client $i$, and $D$ represents the union of all $D_{i}$.
+$\mathbf{w}^{*}$ $=\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (1)
+Here, $F_{i}(\cdot)$ represents the local loss function of client $i$. The weights are defined as $p_{i}=\frac{|D_{i}|}{\sum_{j}|D_{j}|}$, $D_{i}$ represents the dataset of client $i$.
 We have defined balanced performance fairness in FL as follows.
 Definition 1
-(Balanced performance fairness [28]) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
-In this study, we employ the loss of model denote the performance.
+(Balanced performance fairness [25]) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the NN devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the NN devices.
+Differing from [28] that uses the variance of accuracy for "performance", we employ the variance of loss.
+This enables the metric to be incorporated directly into the objective function, facilitating more straightforward iterative optimization.
 Equation (2) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients.
 Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
-$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (2)
-III-B Differential Privacy
-Differential privacy is a formal mathematical model that quantifies privacy leakage in data analysis algorithms.
+$\Psi(\mathbf{w})=\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (2)
+2.2 DP guarantee in FL
+DP is a formal mathematical model that quantifies privacy leakage in data analysis algorithms.
 It stipulates that alterations to a single record within the training data should not induce significant statistical variations in the algorithm’s results.
 Definition 2
-(Differential Privacy [11]).
-$(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
+(DP [11]).
+$(\epsilon,\delta)$-DP is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single data sample, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
 $\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$ (3)
 By adding random noise, we can achieve differential privacy for a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ according to Definition 2.
-The $l_{k}$-sensitivity determines how much noise is needed and is defined as follow.
-We use $l_{2}$ sensitivity as in the previous work.
+The $l_{2}$-sensitivity determines how much noise is needed and is defined as follow.
 DPSGD. In gradient-based methods, DPSGD[1] is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
-  • Firstly, computing the gradient for data sample $\xi_{j}$: $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
+  • Firstly, computing the gradient for data sample $\xi_{j}$, donated by $\mathbf{g}^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
   • Secondly, clipping the gradient of per-sample to get the sensitivity:
-    $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right),$ (4)
+    $\mathbf{\hat{g}}^{i,j}_{t}=\mathbf{g}^{i,j}_{t}/\max\left(1,\frac{\|\mathbf{g}^{i,j}_{t}\|}{C}\right),$ (4)
     where $C$ represents the clipping bound.
-  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier:
-    $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$ (5)
+  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\cdot\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier, $\mathbf{I}$ is the identity matrix:
+    $\mathbf{\tilde{g}}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\mathbf{\hat{g}}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$ (5)
 The final step is to proceed with gradient descent as usual.
 When each client in FL utilizes DPSGD as the optimization algorithm to safeguard their private data, the framework transitions from FL to DPFL.
-IV FedFair: Equipping FL with Fairness
+3 FedFair: Equipping FL with Fairness
 Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, fairness hyperparameter $\lambda$, batch sample ratio $q$, local dataset $D_{i}$.
 Output: the final trained model $\mathbf{w}_{T}$
-1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
+1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$;
 2 SERVER
-3       for $t\in[0,1,\cdots,T-1]$  do
-4             for $i\in[1,2,\cdots,N]$ parallel  do
+3 for $t=0,1,\cdots,T-1$ do
+4    for $i=1,2,\cdots,N$ parallel do
 5       $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
@@
 12LOCAL
-13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
+13 $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
 14
-15      $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
-16       for batch $b\in\mathcal{B}_{i}$ do
+15 $\{\mathcal{B}_{i}\}\leftarrow$ Split $D_{i}$ to batches;
+16 for batch $b\in\{\mathcal{B}_{i}\}$ do
 17    $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i},b)-F(\mathbf{w}_{t})$;
@@
 $\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$ (7)
-where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training.
-from round $t-1$, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning.
-When $\lambda\rightarrow\infty$, the algorithm only focus on fairness.
+where $F(\mathbf{w}^{i}_{t})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{t})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training.
+from round t1t-1, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning.
+When $\lambda\rightarrow\infty$, the algorithm only focuses on fairness.
 However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging.
-Fortunately, we theoretically found an optimal value for $\lambda$ in Section VI-B.
+Fortunately, we theoretically found an optimal value for $\lambda$ in Section 5.1.
 In addition, as shown in Figure 2, we experimentally confirmed the existence of the optimal $\lambda$..
@@
 $=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (8)
-Figure 2: FedFDP framework.
 where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t}^{i})$.
-The stochastic gradient descent by using Equation (IV) as show:
+The stochastic gradient descent by using Equation (3) as show:
 $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (9)
@@
 Based on these principles, we introduce the FedFair algorithm, as describe in Algorithm 1:
+  1. 1.
     Server (lines 1-1): The server executes $T$ rounds of communication, each round broadcasting the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
-  • Local (lines 1-1): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
-V FedFDP: Further Equipping FedFair with Differential Privacy
+  2. 2.
+    Local (lines 1-1): For each batch $b$, we compute the dynamic learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
+4 FedFDP: Further Equipping FedFair with Differential Privacy
 In this section, we present our algorithm, FedFDP, which further integrates FedFair with differential privacy.
 The flow of FedFDP algorithm is shown in Figure 1.
-It adds two additional processes to FedFair:
-  • A fair-clipping strategy is added to FedFair. This strategy not only achieves sensitivity but also allows adjustments to the level of fairness.
-  • An adaptive clipping method is employed when protecting the loss values that clients share with differential privacy, aiming to maximize utility.
+In addition to adding Gaussian noise to the gradients, we introduce two additional processes to balance fairness and DP: 1) A fair-clipping strategy is added to FedFair.
+This strategy not only achieves sensitivity but also allows adjustments to the level of fairness.
+2) An adaptive clipping method is employed when protecting the loss values that clients share with differential privacy, aiming to maximize utility.
 Next, we discuss the details of these two strategies.
-V-A Fair-clipping Strategy for Gradient
+4.1 Fair-clipping Strategy for Gradient
 In the general DPSGD algorithm [1, 13, 5, 41], per-sample clipping is an essential process for obtaining sensitivity.
 In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (7), into Equation (10):
-$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$ (10)
+$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$ (10)
 where $\xi_{j}\in\mathcal{B}_{t}^{i}$, $j=\{1,2,\cdots,|\mathcal{B}_{i}|\}$, and $|\mathcal{B}_{i}|$ is the batch size of client $i$.
@@
 To ensure the algorithm is safeguarded by differential privacy, we replace the gradient of the loss function with Equation (11) and reformulate the gradient descent computation for client $i$ as follows:
-$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\eta\tilde{g}^{i}_{t}$
-$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
-$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
-$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (12)
+$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\eta\mathbf{\tilde{g}}^{i}_{t}$
+$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right]$
+$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right]\,$, (12)
 where:
@@
 The impact of $C$ on the performance of FedFDP will be further discussed in the experiments.
-V-B Adaptive Clipping Method for Loss
+4.2 Adaptive Clipping Method for Loss
 Since calculating $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, differential privacy must be incorporated to ensure the algorithm preserves privacy.
-As shown in lines 2-2 of Algorithm 2, $F_{i}(\mathbf{w}_{t}^{i})$ must be clipped to the interval $[0,C_{l}^{i,t}]$, followed by the addition of Gaussian noise with a mean of 0 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy.
+As shown in lines 2-2 of Algorithm 2, $F_{i}(\mathbf{w}_{t}^{i})$ must be clipped to the interval $0,C_{l}^{i,t}]$, followed by the addition of Gaussian noise with a mean of 0 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy.
 However, determining an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging.
-A large $C_{l}^{i,t}$ prevents distortion of the original values but results in larger injected noise.
+In federated learning, due to the heterogeneity of data held by different clients, $F_{i}(\mathbf{w}_{t}^{i})$ can vary significantly.
+A clipping norm that is too large introduces excessive noise, while one that is too small leads to aggressive clipping of gradient directions, impairing model performance.
+Therefore, an adaptive clipping strategy is necessary.
 Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t}^{i})$.
@@
 1 SERVER
-2       for $t\in[0,1,\cdots,T-1]$  do
-3             for $i\in[1,2,\cdots,N]$ parallel  do
+2 for $t=0,1,\cdots,T-1$ do
+3    for $i=1,2,\cdots,N$ parallel do
 4       $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
@@
 11LOCAL
-12       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
+12 $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
 13   Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
-14       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
+14 for $j=1,2,\cdots,|\mathcal{B}_{i}|$ do
 15    $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
 16      Compute $C^{i,j}_{t}$ by Equation (13);
-17             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
+17    $\mathbf{\tilde{g}}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
 18
-19      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right)$;
+19 $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\mathbf{\tilde{g}}_{t}^{i,j}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right)$;
 20
 21  Compute clipping bound of loss $C_{l}^{i,t}$ by Equation (14) and (15) ;
-22       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
+22 for $j=1,2,\cdots,|\mathcal{B}_{i}|$ do
 23      Compute $F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})$;
@@
 For each communication round, the clipping bound follows the post-processing property of differential privacy.
-V-C Overall Algorithm
+Figure 1: FedFDP framework.
+Besides the privately computed local model $\mathbf{w}_{t+1}^{i}$ from the “Fair-Clipping”, client $i$ is also required to upload the private loss $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$.
+4.3 Overall Algorithm
 We utilize the equation form Equation (12) to enhance Algorithm 2. Compared with the Algorithm 1, FedFDP solely necessitates the application of differential privacy to gradient and loss.
-  • DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$.
-Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 2 and execute the gradient decent.
-  • DP for local loss (lines 2-2): As described in Section V-B, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
-VI Finding Optimal Fairness Parameter
+  1. 1.
+    DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\mathbf{\tilde{g}}_{t}^{i,j}$.
+Then, we add mean 0 noise to the sum of per-sample processed gradient $\mathbf{\tilde{g}}_{t}^{i,j}$ in line 2 and execute the gradient decent.
+  2. 2.
+    DP for local loss (lines 2-2): As described in Section 4.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
+5 Finding Optimal Fairness Parameter
 In this section, we discuss how to identify an optimal fairness parameter $\lambda$ through the convergence analysis of the FedFDP algorithm.
-VI-A Convergence Analysis
-We analyze the convergence of Algorithm 2 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound.
-Prior to proving this, we need to establish several assumptions:
-Assumption 1
-$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
-Assumption 2
-$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
-Assumption 3
-Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random.
-The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i\in[1,\cdots,N]$ , $t\in[0,\cdots,T-1]$ and $j\in[1,\cdots,|\mathcal{B}_{i}|]$
-Theorem 1
-Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 2 as follows:
-$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (16)
-where:
-  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
-  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
-  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
-  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
-The meaning of $\Gamma$ is consistent with [29], where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
-proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$.
-Finally, by utilizing Assumption 1, we derive Equation (16).
-For the detailed proof, please refer to Appendix A-A.
-VI-B Optimal Fairness Parameter $\lambda^{*}$
-Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$.
-We rearrange the convergence upper bound in Equation (16) to obtain a function in terms of $\lambda$,
+All assumptions and detailed procedures for the convergence analysis are provided in Appendix 0.A.2, leading to the following result:
+Theorem 5.1(Convergence rate of FedFDP   (Simplified version of Theorem 0.A.1))
+$\mathbb{E}[F(\mathbf{w}_{T})]-F^{*}$ $\leq\mathcal{O}\left(\frac{LG^{2}\lambda^{2}}{\mu^{2}T}\right)+\mathcal{O}\left(\frac{L^{2}\Gamma\lambda}{\mu^{2}T}\right)$
+$\quad+\mathcal{O}\left(\frac{L\sigma^{2}C^{2}d}{\mu^{2}T\lambda}\right)+\mathcal{O}\left(\frac{L\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}}{T}\right)\,$, (16)
+where $L$, $\mu$, and $G$ are parameters defined in Assumptions 13, and $\Gamma$ is a measure of heterogeneity; see Appendix 0.A.1 for details.
+As demonstrated in (5.1), the fairness parameter $\lambda$ ($\lambda>0$) exhibits convex influence on the convergence rate.
+The analytical solution of this relationship will be obtained via extremum analysis in Section 5.1.
+5.1 Optimal Fairness Parameter $\lambda^{*}$
+Assuming $\Delta_{i}^{j}$ is bounded as $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$.
+We rearrange the convergence upper bound in Equation (5.1) to obtain a function in terms of $\lambda$:
 $P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$ (17)
@@
 If $\mathcal{F}(\lambda)$ has an extremum, then $\mathcal{F}^{\prime}(\lambda)$ has a root, meaning $\mathcal{G}(\lambda)$ has a root.
-Substituting $\lambda=0$, $\mathcal{G}(0)=a_{3}-a_{4}a_{5}$, evidently when $a_{3}<a_{4}a_{5}$, $\mathcal{G}(\lambda)$ has a root.
+Substituting $\lambda=0$, we have $\mathcal{G}(0)=a_{3}-a_{4}a_{5}$, evidently when $a_{3}<a_{4}a_{5}$, $\mathcal{G}(\lambda)$ has a root.
 Let $\lambda=x-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, replacing $\lambda$ with $x$ in $\mathcal{G}(\lambda)$ yields
@@
 Using the root-finding formula to calculate $\mathcal{G}(x)=0$ gives the solutions $x_{1},x_{2},x_{3}$.
+$x_{1}=\sqrt[3]{-\frac{a_{7}}{2}+\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}}+\sqrt[3]{-\frac{a_{7}}{2}-\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}},$
+$x_{2}=\omega\sqrt[3]{-\frac{a_{7}}{2}+\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}}+\omega^{2}\sqrt[3]{-\frac{a_{7}}{2}-\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}},$
+$x_{3}=\omega^{2}\sqrt[3]{-\frac{a_{7}}{2}+\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}}+\omega\sqrt[3]{-\frac{a_{7}}{2}-\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}}.$
+where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{l}=x_{l}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, $l\in\{1,2,3\}$.
+Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root.
+If the equation has $\mathrm{m}$ real roots, the optimal solution for $\lambda$ is $\lambda^{*}=\max_{\mathrm{m}}(\lambda_{\mathrm{m}})$.
+In conclusion, $\mathcal{F}^{\prime}(\lambda)\geq 0\iff G(\lambda)\geq 0\iff\lambda\geq\lambda^{*}$, therefore $\mathcal{F}(\lambda)$ has a minimum value $\mathcal{F}(\lambda^{*})$, and $P(\lambda)$ has a minimum value $P(\lambda^{*})$.
+6 Privacy Analysis
+At the beginning of privacy analysis, we clarify that: (1) FedFDP does not assume Secure Aggregation (SecAgg)[4], allowing the server to observe individual client uploads; (2) we guarantee sample-level DP via per-sample clipping to protect individual data records; and (3) SecAgg is orthogonal to our method—while it hides per-client updates, it cannot replace DP in preventing the aggregated model from memorizing sensitive data.
+As Algorithm 2 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$.
+$\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of client $i$, so they are both to perform differential privacy preservation, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively.
+We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
+6.1 Privacy Loss of $\mathbf{w}_{t+1}^{i}$
+Theorem 6.1
+After $T$ rounds local updates, the RDP of the $\mathbf{w}_{T}^{i}$ in $i$-th client satisfies:
+$R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$ (23)
+where $\sigma$ is noise multiplier of the $\mathbf{w}_{t+1}^{i}$, and $\alpha>1$ is the order.
+Proof.
+We will prove Theorem 6.1 in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
+Definition 3
+(RDP privacy budget of SGM[32]). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity $1$, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
+$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))\,$, (24)
+where
+$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)=\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\,,\\ B_{\alpha}(q,\sigma)=\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\,.\end{array}\right$.
+(25)
+with $\vartheta_{0}=\mathcal{N}\left(0,\sigma^{2}\right),\vartheta_{1}=\mathcal{N}\left(1,\sigma^{2}\right)\mbox{ and }\vartheta=(1-q)\vartheta_{0}+q\vartheta_{1}$ .
+Further, it holds for $\forall(q,\sigma)\in(0,1],\mathbb{R}^{+*},A_{\alpha}(q,\sigma)\geq B_{\alpha}(q,\sigma)$.
+Thus, $SG_{q,\sigma}$ satisfies $\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
+Finally, the existing work [32] describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$:
+$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)\,.$ (26)
+Definition 4
+(Composition of RDP[33]).
+For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as (X,Y)(X,Y)(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
+From Definition 3 and Definition 4, the Theorem 6.1 is obtained.
+6.2 Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$
+Theorem 6.2
+After $T$ rounds local updates, the RDP of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ in $i$-th client satisfies:
+$R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$ (27)
+where $\sigma_{l}$ is noise multiplier of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$, and $\alpha>1$ is the order.
+The proof is similar to that of the $\mathbf{w}_{t+1}^{i}$, so we omit it.
+6.3 Privacy Loss of FedFDP
+Since both $\mathbf{w}_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 4, and then use Lemma 1 to convert it to $(\epsilon,\delta)$-DP.
+Lastly, we can get the Privacy loss of FedFDP as follows.
+Theorem 6.3
+(Privacy loss of FedFDP). The privacy loss in $i$-th client of FedFDP satisfies:
+$\begin{split}\begin{aligned} (\epsilon^{i},\delta^{i})=&(R_{model}^{i}(\alpha)+R_{loss}^{i}(\alpha)+\ln((\alpha-1)/\alpha)\\ &-(\ln\delta+\ln\alpha)/(\alpha-1),\delta),\end{aligned}\end{split}$ (28)
+where $0<\delta<1$, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 6.1, and $R_{loss}^{i}(\alpha)$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 6.2.
+Lemma 1
+[…]
+7 Experiments
+In this section, we demonstrate the effectiveness of FedFair/FedFDP through answering the following three research questions:
+  • RQ1 How effective is FedFair / FedFDP under classification tasks with real-world datasets?
+  • RQ2 How robust is FedFair / FedFDP across different experimental settings?
+  • RQ3 What is the performance of FedFDP under differential hyper-parameters?
+To address the aforementioned three RQs, Section 7.1 introduces the default experimental settings used in this study, including the description of baselines and the configuration of the experimental code.
+Section 7.2 presents comparative experiments conducted on three datasets under default parameters involving FedFair, FedFDP, and six baseline methods to answer RQ1.
+Section 7.3 addresses RQ2 through experiments on heterogeneity and scalability.
+Finally, Section 7.4 performs ablation studies on $\lambda$, $C$, $\sigma$, $\epsilon$, and the target accuracy to answer RQ3.
+7.1 Default Settings
+Prior to addressing the three research questions, an outline of the default experimental configuration is provided.
+Baselines.
+We compared our algorithms against six baseline methods.
+For FedFair, the baseline algorithms were left unmodified; for FedFDP, DP was applied to each baseline algorithm to enable a fair comparison, which include FedAvg[30], SCAFFOLD[17], FedProx[24], FedDyn[2], ALI-DPFL[28] and q-FFL[25].
+Table 1: Comparison of Communication Overhead and Transmission Requirements. $d$ represent the dimension of model.
+Algorithm
+Comm. Cost
+(Per Round)
+Extra Transmitted
+Variables
+Description of Overhead
+FedAvg $\mathcal{O}(d)$ None Transmits only model parameters/gradients.
+FedProx $\mathcal{O}(d)$ None Adds proximal term locally.
+q-FFL $\mathcal{O}(d)$ Scalar Sends scalar loss for re-weighting.
+SCAFFOLD $\mathcal{O}(2d)$ Control Variates Transmits both model and control variates.
+FedDyn $\mathcal{O}(d)$ None Improve SCAFFOLD without control variables.
+ALI-DPFL $\mathcal{O}(d)$ Scalar Server broadcast scalar local steps to clients.
+FedFDP $\mathcal{O}(d)$ Scalar Transmits scalar loss for fairness/privacy.
+Communication Overhead Analysis.
+Table 1 presents a detailed comparison of the communication complexity per round for FedFDP and other state-of-the-art baselines.
+While methods like SCAFFOLD introduce control variates to mitigate client drift, they double the communication payload to $\mathcal{O}(2d)$, which can be prohibitive in bandwidth-constrained wireless networks.
+In contrast, FedFDP maintains a communication complexity of $\mathcal{O}(d)$, aligning with efficient baselines such as FedAvg and FedDyn .
+Although FedFDP incorporates a fairness-aware mechanism and differential privacy, the additional transmission overhead is restricted to a negligible scalar value (representing the fairness loss), similar to the approach in q-FFL and ALI-DPFL.
+This demonstrates that FedFDP achieves robust privacy and fairness without compromising communication efficiency.
+Tasks setting.
+By presenting the comparative experimental results of FedFair and FedFDP with multiple baseline methods on three datasets: MNIST [19], FashionMNIST [40], and CIFAR10 [18], we adopted a widely used heterogeneous settings [27, 21] to 10 clients which controlled by a Dirichlet distribution denoted as $\mathrm{Dir}(\beta)$, and the default value of $\beta=0.1$ [27, 38].
+We use a 4-layer CNN architecture [30] which consists of two convolutional layers and two fully connected layers as model architecture.
+Implementation environment.
+We implement our experiment using PyTorch-1.8 and run all experiments on a server with one Intel i9 13900ks CPU (24 cores), 64GB of memory, and one NVIDIA 4090 GPU, running on Windows 10.
+Hyperparameters.
+For FedFair and its baseline algorithms, we set learning rate $\eta=0.1$.
+For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1.0\times 10^{-5}$.
+In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
+7.2 Effective in Real-world Datasets
+To address RQ1, we conducted experiments on three datasets for both FedFair and FedFDP under the default parameter settings described in Section 7.1.
+The results, presented in Table 2 and Table 3, report the test accuracy ($\%$) and fairness measure ($\Psi$).
+Table 2: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
+Datasets MNIST FashionMNIST CIFAR10
+Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
 […]
@@
 […]
-where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{l}=x_{l}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, $l=1,2,3$.
-Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root.
-If the equation has $k$ real roots, the optimal solution for $\lambda$ is $\lambda^{*}=\max_{k}(\lambda_{k})$.
-In conclusion, $\mathcal{F}^{\prime}(\lambda)\geq 0\iff G(\lambda)\geq 0\iff\lambda\geq\lambda^{*}$, therefore $\mathcal{F}(\lambda)$ has a minimum value $\mathcal{F}(\lambda^{*})$, and $P(\lambda)$ has a minimum value $P(\lambda^{*})$.
-VII Privacy Analysis
-Since each client incurs the same privacy loss in every round, we evaluate privacy loss from the perspective of an individual client.
-Accordingly, we analyze the privacy loss of FedPDF for the $i$-th client as follows.
-Theorem 2
-(Privacy loss of FedFDP). The privacy loss in $i$-th client of FedFDP satisfies:
 […]
-where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 3, and $R_{loss}^{i}(\alpha)$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 4.
-VII-A Proof of Theorem 2
-Lemma 1
-(Conversion from RDP to DP [3]).
-if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
-As Algorithm 2 shown, the $i$-th client uploads model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$.
-Both $\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the private training data of the client, and thus both require differential privacy protection, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively.
-We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
-VII-A1 Privacy Loss of $\mathbf{w}_{t+1}^{i}$
-Theorem 3
-After $T$ rounds local updates, the RDP of the $\mathbf{w}_{T}^{i}$ in $i$-th client satisfies:
-$R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$ (24)
-where $\sigma$ is noise multiplier of the $w_{t+1}^{i}$, and $\alpha>1$ is the order.
-$proof.$ We will prove this in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
-Definition 3
-(RDP privacy budget of SGM[37]). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
-$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$ (25)
-where
-$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$ (26)
-with $\vartheta_{0}\triangleq\mathcal{N}\left(0,\sigma^{2}\right),\vartheta_{1}\triangleq\mathcal{N}\left(1,\sigma^{2}\right)\mbox{ and }\vartheta\triangleq(1-q)\vartheta_{0}+q\vartheta_{1}$
-Further, it holds $\forall(q,\sigma)\in(0,1],\mathbb{R}^{+*},A_{\alpha}(q,\sigma)\geq B_{\alpha}(q,\sigma)$.
-Thus, $SG_{q,\sigma}$ satisfies $\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
-Finally, the existing work [37] describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
-$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$ (27)
-Definition 4
-(Composition of RDP[36]).
-For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
-From Definition 3 and Definition 4, the Theorem 3 is obtained.
-VII-A2 Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$.
-Theorem 4
-After $T$ rounds local updates, the RDP of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ in $i$-th client satisfies:
-$R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$ (28)
-where $\sigma_{l}$ is noise multiplier of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$, and $\alpha>1$ is the order.
-The proof is similar to that of the $w_{t+1}^{i}$, so we omit it.
-VII-A3 Proof Loss of FedFDP.
-Since both $w_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 4, and then use Lemma 1 to convert it to $(\epsilon,\delta)$-DP.
-Therefore, the Theorem 2 is hold.
-VIII Experiments
-TABLE II: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
+[…]
+[…]
+Table 3: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
 Datasets MNIST FashionMNIST CIFAR10
 Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
-FedAvg 98.67$\pm$0.09 7.3e-3$\pm$5.4e-3 87.05$\pm$1.02 3.6e-1$\pm$2.2e-1 62.03$\pm$1.00 1.1e0$\pm$8.9e-1
-SCAFFLOD 98.45$\pm$0.03 3.0e-2$\pm$2.6e-2 86.77$\pm$0.25 6.1e0$\pm$4.0e0 62.21$\pm$1.37 9.9e0$\pm$1.1e1
-FedProx 98.70$\pm$0.13 7.5e-3$\pm$5.1e-3 87.30$\pm$0.51 3.5e-1$\pm$2.1e-1 62.13$\pm$0.20 1.2e0$\pm$3.1e-1
-FedDyn 98.40$\pm$0.23 8.0e-3$\pm$2.1e-3 87.48$\pm$0.31 6.2e-1$\pm$3.5e-1 62.33$\pm$1.15 3.2e0$\pm$8.8e-1
-q-FFL 98.72$\pm$0.03 4.7e-3$\pm$1.0e-3 87.35$\pm$0.13 5.1e-1$\pm$5.2e-2 63.33$\pm$0.21 9.4e-1$\pm$3.3e-2
-FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
-TABLE III: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 2 presents the experimental results of the FedFair algorithm compared to baseline methods.
+FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
+A plausible explanation is that, as the training process progresses, FedFair increasingly prioritizes fairness, resulting in a reduction in $\eta_{i}$.
+Compared to training methods with a fixed learning rate, learning rate decay accelerate the convergence of gradient-based methods [45].
+Table 3 presents the experimental results of the FedFDP algorithm compared to baseline methods.
+In comparison with Table 2, the introduction of DP leads to a slight decrease in accuracy, approximately in the range of $3\%$ - $10\%$.
+However, the fairness issue becomes more pronounced, with the corresponding metric increasing significantly in magnitude, indicating that DP exerts a substantial influence on fairness in FL.
+Under a fixed privacy budget of $\epsilon=3.52$, FedFDP significantly enhances fairness while matching the accuracy of the best-performing baseline, achieving improvements of $30.3\%$, $6.7\%$, and $27.8\%$ across three distinct datasets, respectively.
+7.3 Robust in Different Settings
+To answer RQ2, as shown in Table 4 and Table 5, we extended the heterogeneity setting from the default $\mathrm{Dir}(0.1)$ to $\mathrm{Dir}(0.5)$ and $\mathrm{Dir}(1)$, and scaled the number of clients from the default 10 to 20 and 50 to examine scalability.
+Table 4: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
 Heterogeneity Scalability
@@
 FedAvg 98.67 7.3e-3 98.37 1.2e-3 89.41 3.6e-2 62.03 1.1e0 60.89 1.5e0 59.78 1.5e0
-SCAFFLOD 98.45 5.0e-3 98.51 2.1e-3 86.39 3.4e-2 62.21 9.9e0 59.48 8.7e0 58.64 2.3e0
+SCAFFOLD 98.45 5.0e-3 98.51 2.1e-3 86.39 3.4e-2 62.21 9.9e0 59.48 8.7e0 58.64 2.3e0
 FedProx 98.70 7.5e-3 98.35 1.2e-3 88.33 3.5e-2 62.13 1.2e0 60.33 2.2e0 61.31 1.6e0
@@
 FedFair 98.75 3.9e-3 98.72 1.2e-3 89.60 1.7e-2 62.38 8.5e-1 61.58 1.1e0 62.90 1.5e0
-In this section, we present the comparative experimental results of FedFair and FedFDP with multiple baseline methods on three datasets: MNIST [22], FashionMNIST [48], and CIFAR10 [21].
-We provide a detailed description of the experimental evaluation of FedFair and FedFDP in terms of performance, fairness, heterogeneity, and scalability.
-In particular, we conduct ablation studies on the fairness hyperparameter $\lambda$, privacy budget $\epsilon$, clipping bound $C$, and noise multiplier $\sigma$.
-In all tables, $\uparrow$ indicates that larger values are better, while $\downarrow$ indicates that smaller values are better.
-We implement our experiment using PyTorch-1.8 and run all experiments on a server with one Intel i9 13900ks CPU (24 cores), 64GB of memory, and one NVIDIA 4090 GPU, running Windows 10.
-VIII-A Default Settings
-We compared our algorithms against six baseline methods. For FedFair, the baseline algorithms were left unmodified; for FedFDP, DP was added to each baseline algorithm to enable a fair comparison.
-  1. 1.
-    FedAvg[33]: FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
-  2. 2.
-    SCAFFLOD[19]: SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
-  3. 3.
-    FedProx[27]: FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
-  4. 4.
-    FedDyn[2]: FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
-  5. 5.
-    ALI-DPFL[31]: ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
-  6. 6.
-    q-FFL[28]: q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
-For all datasets, we adopted a widely used heterogeneous settings [30, 24] to 10 clients which controlled by a Dirichlet distribution denoted as $\mathrm{Dir}(\beta)$, and the default value of $\beta=0.1$ [30, 46].
-We use a 4-layer CNN architecture [33] which consists of two convolutional layers and two fully connected layers as model architecture.
-Hyperparameters.
-For FedFair and its baseline algorithms, we set learning rate $\eta=0.1$.
-For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$.
-In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
-TABLE IV: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
-Datasets MNIST FashionMNIST CIFAR10
-Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
-FedAvg 93.40$\pm$0.47 1.1e11$\pm$0.5e10 84.15$\pm$1.97 3.0e8$\pm$1.4e8 52.61$\pm$0.17 6.1e9$\pm$5.2e9
-SCAFFLOD 93.95$\pm$1.39 7.0e10$\pm$4.9e9 83.41$\pm$4.87 3.3e9$\pm$4.7e9 53.85$\pm$1.16 6.2e9$\pm$1.3e9
-FedProx 90.50$\pm$3.95 5.3e10$\pm$7.3e9 83.85$\pm$0.98 7.9e8$\pm$4.5e8 51.34$\pm$0.91 4.4e9$\pm$2.1e9
-FedDyn 91.42$\pm$2.25 5.6e10$\pm$9.8e9 84.04$\pm$1.21 8.8e8$\pm$3.2e8 52.14$\pm$1.39 5.5e9$\pm$1.2e9
-ALI-DPFL 90.89$\pm$1.65 3.3e10$\pm$8.3e9 83.65$\pm$1.68 6.6e8$\pm$3.5e8 52.05$\pm$1.26 3.6e9$\pm$1.6e9
-q-FFL 93.74$\pm$1.41 7.8e10$\pm$1.1e11 83.13$\pm$2.74 4.2e9$\pm$2.6e9 48.46$\pm$1.00 4.7e9$\pm$1.9e9
-FedFDP 95.13$\pm$0.83 2.3e10$\pm$1.0e10 85.99$\pm$0.76 2.8e8$\pm$0.8e8 54.21$\pm$0.98 2.6e9$\pm$1.6e9
-TABLE V: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
-Herterogeneity Scalability
+Table 5: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
+Heterogeneity Scalability
 Datasets MNIST FashionMNIST CIFAR10
@@
 FedAvg 93.40 1.1e11 94.17 4.1e9 85.60 3.7e9 52.61 6.1e9 51.58 2.0e9 55.90 1.5e9
-SCAFFLOD 93.95 7.0e10 94.06 1.9e9 84.56 3.6e9 53.85 6.2e9 54.34 2.9e8 53.98 2.2e9
+SCAFFOLD 93.95 7.0e10 94.06 1.9e9 84.56 3.6e9 53.85 6.2e9 54.34 2.9e8 53.98 2.2e9
 FedProx 90.50 5.3e10 94.23 3.1e9 84.38 5.6e9 51.34 4.4e9 53.67 4.5e8 54.14 2.2e9
@@
 FedFDP 94.13 2.3e10 94.56 3.4e9 86.77 6.0e8 54.21 2.6e9 54.52 2.2e8 56.49 5.3e8
-[…]
+Heterogeneity.
+Table 4 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity.
+Except for the $\mathrm{Dir}(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
+Table 5 shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity.
+Except for the $\mathrm{Dir}(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
+Scalability.
+Table 4 indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
+Table 5 demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline.
+Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
+Figure 2: Impact of different $\lambda$ in FedFDP algorithm.
+Figure 3: Impact of different $C$ in FedFDP algorithm.
+Figure 4: Impact of different $\sigma$ in FedFDP algorithm.
+7.4 Performance under Different Hyper-parameters
+In response to RQ3, we performed extensive hyper-parameter experiments.
+In Table 6, we observed results under fixed privacy budgets $\epsilon=\{1.0,2.0,3.0,4.0\}$.
+Table 7 presents experiments aiming at fixed target accuracies $\{80\%,85\%,90\%,95\%\}$.
+Fig.
+2 explores the existence of an optimal $\lambda$, with results consistent with those in Section 5.
+The influence of the clipping norm $C$ is investigated in Fig.
+3, and the effect of the noise multiplier $\sigma$ is examined in Fig.
+4.
+Impact of Different $\lambda$.
+In Section 5.1, we derive a closed-form solution for $\lambda^{*}$, which, however, relies on constants or upper bounds (e.g., $L$, $\mu$, $G$, $\Gamma$) whose specific values change with different models, datasets, and loss functions.
+Therefore, we conduct the experiment in Fig.
+2 to investigate the practical range of $\lambda^{*}$.
+We investigated the correlation between $\lambda$ and $\Psi$ in the FedFDP algorithm.
+As depicted in Fig.
+2, our findings align with the theoretical analysis: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
+Impact of Different $C$.
+We emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously.
+As illustrated in Fig.
+3, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range {0.01, 0.1, 1, 5, 10}.
+Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition.
+As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$.
+When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
+Impact of Different $\sigma$.
+We tested the variations in accuracy and fairness on two datasets when $\sigma$ was set to $\{1,1.5,2,2.5,3\}$, under the condition of $\epsilon=2$ and $q=0.05$, which supports $T\in\{6,115,268,463,708\}$ rounds of communication.
+As shown in Fig.
+4, an increase in $\sigma$ allows more training epochs to advance the model’s convergence process, but when $\sigma=3$, excessive noise hinders the convergence.
+When $\sigma$ is smaller, the model performance is relatively similar, with a lower $\Psi$ value.
+As $\sigma$ gradually increases, the performance differences between models become apparent, and the $\Psi$ value increases.
+Table 6: The average accuracy ($\%$) and fairness ($\Psi$) on different privacy budget ($\epsilon$) setting at FashionMNIST dataset.
 $\epsilon$ 1.0 2.0 3.0 4.0
@@
 FedAvg 61.68 1.1e6 82.69 2.3e7 84.07 3.8e8 86.83 8.0e8
-SCAFFLOD 62.12 1.2e6 82.36 2.3e7 83.39 2.6e9 85.41 5.6e8
+SCAFFOLD 62.12 1.2e6 82.36 2.3e7 83.39 2.6e9 85.41 5.6e8
 FedProx 61.25 1.1e6 81.62 2.3e7 85.32 3.8e8 85.68 2.4e9
@@
 FedFDP 63.36 1.0e6 83.15 2.1e7 85.59 9.5e7 86.97 3.8e8
-VIII-B Result of FedFair
-Performance and Fairness.
-Table II presents the experimental results of the FedFair algorithm compared to baseline methods.
-FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
-A plausible explanation is that, as the training process progresses, FedFair increasingly prioritizes fairness, resulting in a reduction in $\eta_{i}$.
-Compared to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods [55].
-Figure 3: Impact of different $\lambda$ in FedFDP algorithm.
-Figure 4: Impact of different $C$ in FedFDP algorithm.
-Figure 5: Impact of different $\sigma$ in FedFDP algorithm.
-Heterogeneity.
-Table III shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity.
-Except for the $\mathrm{Dir}(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
-Scalability. Table III indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
-VIII-C Results of FedFDP
-Performance and Fairness.
-Table IV displays the experimental results of the FedFDP algorithm compared to baseline methods.
-Under a consistent privacy budget of $\epsilon=3.52$, when accuracy matches the best baseline, our method significantly enhances fairness, with improvements of 30.3%, 6.7%, and 27.8% observed across three separate datasets.
-Heterogeneity.
-Table V shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity.
-Except for the $\mathrm{Dir}(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
-Scalability.
-Table V demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline.
-Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
-Impact of Different $\lambda$.
-We investigated the correlation between $\lambda$ and $\Psi$ in the FedFDP algorithm.
-As depicted in Figure 3, our findings align with the theoretical analysis: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
-Impact of Different $C$.
-We emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously.
-As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10].
-Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition.
-As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$.
-When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
-Different $\epsilon$.
-We conducted experiments with $\epsilon=[1,2,3,4]$ on the FashionMNIST dataset, as detailed in Table VI.
-For $\epsilon=[1,2,3]$, we allowed larger values of $\epsilon$ to support more iterations.
+Impact of Different $\epsilon$.
+We conducted experiments with $\epsilon\in\{1,2,3,4\}$ on the FashionMNIST dataset, as detailed in Table 6.
+For $\epsilon\in\{1,2,3\}$, we allowed larger values of $\epsilon$ to support more iterations.
 At $\epsilon=4$, we supported a smaller noise multiplier $\sigma=1.65$ with a similar number of iterations as at $\epsilon=3$.
@@
 In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
-Impact of Different $\sigma$.
-We tested the variations in accuracy and fairness on two datasets when $\sigma$ was set to $[1,1.5,2,2.5,3]$, under the condition of $\epsilon=2$ and $q=0.05$, which supports $T=[6,115,268,463,708]$ rounds of communication.
-As shown in Figure 5, an increase in $\sigma$ allows more training epochs to advance the model’s convergence process, but when $\sigma=3$, excessive noise hinders the convergence.
-When $\sigma$ is smaller, the model performance is relatively similar, with a lower $\Psi$ value.
-As $\sigma$ gradually increases, the performance differences between models become apparent, and the $\Psi$ value increases.
-IX Conclusion
+Table 7: The privacy budget ($\epsilon$) and fairness ($\Psi$) on different target accuracy setting at MNIST $\mathrm{Dir}(0.05)$.
+Acc. 80% 85% 90% 95%
+$\epsilon(\downarrow)$ $\Psi(\downarrow)$ $\epsilon(\downarrow)$ $\Psi(\downarrow)$ $\epsilon(\downarrow)$ $\Psi(\downarrow)$ $\epsilon(\downarrow)$ $\Psi(\downarrow)$
+FedAvg 2.43 2.10e9 3.02 8.90e9 3.45 1.30e10 4.89 1.30e11
+SCAFFOLD 2.55 3.30e9 3.15 1.10e10 3.67 3.70e10 5.32 1.10e11
+FedProx 3.02 1.30e9 2.98 1.30e10 4.31 5.80e10 4.89 2.50e11
+FedDyn 2.68 2.20e9 3.55 7.80e9 3.98 1.10e10 5.15 3.30e11
+ALI-DPFL 2.85 3.10e9 2.98 1.30e10 3.67 2.30e10 5.32 9.80e10
+q-FFL 2.68 9.80e8 3.15 9.70e9 4.12 8.80e10 4.04 1.02e11
+FedFDP 2.25 6.60e8 2.70 1.20e9 3.05 5.80e9 3.77 2.80e10
+$\epsilon$ at Different Target Accuracy.
+TABLE 7 presents experiments on MNIST under the $\mathbf{Dir}(0.05)$ partition, reporting the privacy budget $\epsilon$ and fairness measure $\Psi$ at the point of achieving accuracy targets $\{80\%,85\%,90\%,95\%\}$.
+As the target accuracy increases (from $80\%$ to $95\%$), the privacy budget ($\epsilon$) for most methods rises slightly (e.g., FedAvg increases from 2.43 to 3.02), while fairness ($\Psi$) deteriorates significantly (e.g., FedAvg rises from 2.10×109 to 1.3×1011), highlighting the intensified trade-off between privacy and fairness under higher accuracy demands.
+FedFDP demonstrates exceptional performance in balancing these objectives: At the high $95\%$ accuracy target, its $\epsilon$ (2.70) is lower than all other methods, and its $\Psi$ (1.20×109) is significantly lower than competitors by an order of magnitude.
+At the $80\%$ accuracy target, FedFDP achieves the lowest $\Psi$ (6.60×108) and the lowest $\epsilon$ (2.25) in the entire table, validating its comprehensive superiority.
+Figure 5: Tradeoff between privacy and fairness on $60\%$ accuracy.
+Tradeoff between Privacy and Fairness.
+With the target accuracy fixed at $60\%$ on both MNIST and FashionMNIST, we investigate the trade-off between fairness and privacy.
+As shown in Fig.
+5, as the privacy budget $\epsilon$ increases, the fairness metric $\Psi$ decreases significantly on both datasets, indicating a clear trade-off between privacy and fairness under the given accuracy constraint.
+The noise introduced by the privacy mechanism, while protecting data, amplifies performance disparities across groups, whereas relaxing privacy constraints creates room for fairness optimization.
+This monotonically decreasing relationship reveals the tripartite trade-off characteristics among privacy, fairness, and accuracy.
+8 Limitations and Future Work
+Despite the promising results achieved by FedFDP in balancing fairness, privacy, and utility, there are several limitations in the current study that open avenues for future research.
+First, regarding heterogeneity, our current evaluation primarily focuses on statistical heterogeneity (Non-IID data) simulated via Dirichlet partitioning.
+We have not yet deeply explored model heterogeneity, where clients may possess different network architectures due to varying storage or memory capacities.
+Furthermore, the challenges associated with cross-domain data—where feature spaces or distributions differ fundamentally across clients—remain unaddressed.
+Future work could investigate adapting the fairness-aware gradient clipping strategy to support heterogeneous model architectures and cross-domain federated learning scenarios.
+Second, the current framework assumes that all selected clients participate successfully in the training process (full participation).
+However, in practical real-world deployments, device heterogeneity (e.g., varying computational power) and unstable network conditions (e.g., bandwidth limits and high latency) often lead to the problem of stragglers.
+These stragglers can significantly delay the synchronization process in the global aggregation phase.
+In this work, we have not specifically optimized FedFDP for such system-level challenges.
+Future iterations of FedFDP could explore asynchronous update mechanisms or robust client selection protocols to mitigate the impact of stragglers and communication dropouts.
+9 Conclusion
 In this paper, a FedFair algorithm is initially proposed to effectively address fairness concerns by optimizing a novel local loss function.
@@
 It is believed that our work contributes a valuable FL framework for addressing fairness and privacy challenges.
-Appendix A Appendix
-A-A Proof of Theorem 1
+Acknowledge
+This work is supported by National Natural Science Foundation of China Key Program (Grant No.
+62132005), Natural Science Foundation of Shanghai (Grant No.
+22ZR1419100), and CAAI-Huawei MindSpore Open Fund (Grant No.
+CAAIXSJLJJ-2022-005A).
+Appendix 0.A Appendix
+In the beginning, Table 8 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
+Table 8: Summary of main notations
+$F(\mathbf{w}$) Global loss function
+$F_{i}(\mathbf{w}$) Local loss function for client $i$
+$T$ Communication rounds
+$t$ Communication round index
+$D_{i}$ Dataset of client $i$
+$D$ The union of $D_{i}$
+$\mathbf{w}_{t}^{i}$ Local model at client $i$ at round $t$
+$\mathbf{w}^{*}$ Optimal model that minimizes $F(\mathbf{\mathbf{w}})$
+$\eta$ Gradient descent step size
+$\lambda$ Fairness parameter
+NN Total number of clients
+$q$ batch sample ratio
+$|\mathcal{B}_{i}|$ Batch size of client $i$, with the mathematical expectation equals to $|D_{i}|\cdot q$
+$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
+$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
+$\sigma$ Noise multiplier for gradient
+$\sigma_{l}$ Noise multiplier for loss
+$C$ Original clipping bound for gradient
+$C_{l}$ Original clipping bound for loss
+$||\cdot||$ $L_{2}$-norm
+0.A.1 Convergence Analysis
+We analyze the convergence of Algorithm 2 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound.
+Prior to proving this, we need to establish several assumptions:
+Assumption 1
+$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{\top}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
+Assumption 2
+$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{\top}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
+Assumption 3
+Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random.
+The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i\in[1,\cdots,N]$ , $t\in[0,\cdots,T-1]$ and $j\in[1,\cdots,|\mathcal{B}_{i}|]$
+Theorem 0.A.1
+Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 2 as follows:
+$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (29)
+where:
+  • $A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
+  • $\hat{B}=\min_{i}|\mathcal{B}_{i}|$,
+  • $\Gamma=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
+  • $C_{t}=\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}=\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
+The meaning of $\Gamma$ is consistent with [26], where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
+proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$.
+Finally, by utilizing Assumption 1, we derive Equation (29).
+For the detailed proof, please refer to Appendix 0.A.2.
+0.A.2 Proof of Theorem 0.A.1
 For the purpose of validation, we introduce an additional variable $\mathbf{v}_{t}^{i}$ to represent the immediate result of a single-step DPSGD update from $\mathbf{w}_{t}^{i}$.
@@
 Therefore,
-$\mathbf{v}_{t+1}=\mathbf{w}_{t}\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (30)
-A-A1 Key Lemma
+$\mathbf{v}_{t+1}=\mathbf{w}_{t}-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (31)
+0.A.2.1 Key Lemma
 Lemma 2
@@
 where:
-  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
-  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
-  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
-  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
-Let $\Delta_{t}\triangleq\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. It is evident that we always have $\mathbf{w}_{t+1}=\mathbf{v}_{t+1}$. According to Lemma 2, this implies:
+  • $A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
+  • $\hat{B}=\min_{i}|\mathcal{B}_{i}|$,
+  • $\Gamma=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
+  • $C_{t}=\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}=\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
+Let $\Delta_{t}=\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. It is evident that we always have $\mathbf{w}_{t+1}=\mathbf{v}_{t+1}$. According to Lemma 2, this implies:
 $\Delta_{t+1}\leq\left(1-\mu\eta C_{t}\right)\Delta_{t}+\eta^{2}A$
@@
 $\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2}\Delta_{t}\leq\frac{L}{2t}v\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\Delta_{1}\right)$
-A-A2 Proof of Lemma 2
+0.A.2.2 Proof of Lemma 2
 By the Equation (31), we get
@@
 $\mathbb{E}[\mathcal{B}_{2}]\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
-where $\frac{1}{\hat{B}^{2}}\triangleq\max_{i}\frac{1}{|\mathcal{B}_{i}|}$.
+where $\frac{1}{\hat{B}^{2}}=\max_{i}\frac{1}{|\mathcal{B}_{i}|}$.
 By the convexity of $\|\cdot\|^{2}$,
@@
 $\mathbb{E}[\mathcal{B}_{1}]\leq\eta^{2}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\right]\leq\eta^{2}C_{t}^{2}G^{2},$
-where $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i}$, $C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
+where $C_{t}=\sum_{i=1}^{N}p_{i}C_{t}^{i}$, $C_{t}^{i}=\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
 Now, we obtain the bound for the expectation of $\mathcal{A}_{2}$:
@@
 $=-2\eta C_{t}\underbrace{\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}\right)}_{\mathcal{D}_{1}}-2\eta C_{t}\Gamma-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2},$
-where $\Gamma\triangleq\sum_{i=1}^{N}p_{i}(F^{*}-F_{i}^{*})=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$.
+where $\Gamma=\sum_{i=1}^{N}p_{i}(F^{*}-F_{i}^{*})=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$.
 Next, we proceed to handle $\mathcal{D}_{1}$.
@@
 $A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
+Appendix 0.B Related Work
+While various fairness concepts exist in machine learning [7, 29], recent studies in DPFL have predominantly focused on group fairness to mitigate biases related to sensitive attributes like gender and race [14, 34, 15].
+Notable works have integrated DP with fairness constraints (e.g., equal opportunity or decision boundary fairness) to protect attribute privacy while reducing discrimination [16, 37, 9].
+In contrast, our research shifts focus to balanced performance fairness, a relatively unexplored area within DPFL.
+Unlike group fairness which targets demographic parity, balanced performance fairness—introduced by Li et al.
+[25] and further explored in [42, 23]—prioritizes uniform model performance across clients, making it particularly suitable for federated settings involving diverse institutions.
+While parameter tuning (e.g., clipping thresholds) can improve group fairness [15], achieving balanced performance fairness necessitates fundamentally adjusting the direction of model updates.
+Balancing these directional adjustments with the noise and clipping inherent in DP presents a unique challenge.
+To the best of our knowledge, this is the first study to address balanced performance fairness within the context of DPFL.